A Complete Bibliography of Proceedings of the ACM on Human-Computer Interaction (PACMHCI)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

02 March 2022 Version 1.01

Title word cross-reference

#BlackLivesMatter [ASS18]. #pray4victims [BL18].

2 [OSM^{+}20]. 3 [CPB18, KIK20, OSM^{+}20, SEC^{+}20].

19 [HSVR21, YGHR21].

2.0 [KE18]. 2020 [CM20, KLS20, KLS21]. 2021 [BKL^{+}21, BLS21, FJIH21, GMM21]. 2022 [BMdS^{+}21].

360Anywhere [SCY^{+}18]. 360theater [SLN21].

A/B [VZV19]. Aadhaar [SJ21]. AB4Web [VZV19]. Abilities [HBJR21]. Absolute [DWMV21]. Abstracting [AV19]. Abstraction [KE18]. Abuse [RYD21]. Academia [KG21]. Academic [ZD19]. Acceptance [SAE^{+}21]. Access [LZAD21, PKC^{+}21]. Accessibility [BP17b, BMPP20, SCP^{+}21, UAAR20]. Account [HLV21]. Accountability [SSH^{+}21]. Accounting [JA21]. Accuracy [BKG^{+}21, CYGW21, YU20]. Accurate [VMD^{+}20]. Acoustic [ZWZ^{+}21]. Acquisition [UYM21]. Across [BBNT21, FD20, LHT^{+}20]. Acting [ASS18]. Action [HWS21, LO17, NHH20]. Actionable [GYD17]. Active [GLZ^{+}21]. Activists [BWDP20, GYD17]. Activities [BF17, FCE^{+}18, HFBL19, MTA^{+}17].

Activity [BZC^{+}17, BRM21, GRG20, KBKH21, MMM^{+}21]. Actuated [SLT^{+}21]. Ad [SCY^{+}18]. Ad-hoc [SCY^{+}18]. Adaptation [TRO^{+}21]. Adaptations [YHR^{+}19]. Adapting [BS21]. Adaptive [BCCV17]. Address [AL18]. Addressing [RYD21]. Adherence [RAZ^{+}20]. Adolescents [BUSA^{+}21]. Adoption [SAE^{+}21]. ADQDA [LE21]. Adult [ZRL21]. Adults [KSX21]. Advantages [KDH^{+}21]. Advertise [AOT^{+}18]. Advice [HKM^{+}20]. Affect [MFM^{+}20]. Affected [IOES20]. Affects [DBS^{+}21]. Affinity [LE21, SPT^{+}20]. Afforded [Ann20]. afraid [CCWH21]. After [LHPH20, SSH^{+}21, CH20]. Agency [BSS18, HMB^{+}21, RYD21]. Agent [HLV21]. Agents [ZMR21]. Aging [BP17b, SAE^{+}21]. Agriculture [NPT19]. AI [ALD^{+}20, GLZ^{+}21, GHHC21, GHHS21, KG20, LBM21, MBT^{+}20, MHN^{+}21, RYCC21, RGR^{+}21]. AI-Enabled [GHHC21]. AI-Led [KG20]. AI-Mediated [MHN^{+}21]. Air [HGG20, YMG^{+}20, BLC^{+}21]. Airtime [STM^{+}19]. Algorithm [GC19]. Algorithm-in-the-Loop [GC19]. Algorithmic [KDEA21, NM21, SB20b]. Algorithms [BKG^{+}21]. Alignment [CS18a, CT18]. Allocation [HBJR21]. Aloud [MCI20]. Alternatives [VZV19]. Altruism [Vya19]. Ambiguous [CKP^{+}18]. Among [BKKD21, MAR^{+}19, SKV^{+}21]. Amplification [LHPH20]. Analogies [CCH^{+}18]. Analysing [MPB^{+}19]. Analysis [HRRS20, KML^{+}21, KMR20, LE21, MRRS19, SCP^{+}21, SB17, TAS^{+}19, TD19, WGGS20, YHR^{+}19]. Analytics [SB17]. Analyzing [GH21, HSG19, ZMZL21]. Android [MP19]. Annotation [MSY20]. Annotations [CSS^{+}17]. Anomaly [PCMP18]. Anonymity [MCYZ19]. Answer [MWM^{+}19]. Answering [DMC^{+}20]. Antecedents [XZL^{+}20]. Anxiety [DBS^{+}21]. Any [SCY^{+}18]. App [CEM17]. Application [BB20, Sal17, SB17, SLD21, SHY^{+}18, TD19]. Applications [CMN^{+}21, HMB^{+}21, LAR21, VMD^{+}20]. Approach [BCCV17, CFG^{+}17, dACZ19, CXL17, DDFK19, FSC21, GGF^{+}20, KBKH21, LDP18, NM21, SWT20, TVS17]. Apps [GMM19]. April [BKL^{+}21]. Arabian [ADAY21]. Architecture [SVW19]. Archive [CLKEF21]. Art [BVM21]. Arts [LKS21]. Ask [MWM^{+}19]. AskHistorians [Gil20]. Asking [MCYZ19]. Aspects [AS17, SB20b]. Ass [PKC^{+}21]. Assembling [HFL19]. Assembly [HBJR21, KFSC18, MAN^{+}20, YOFC19]. Assertions [BR17]. Assessment [BahRKM18, BZSM20]. Assignment [EHH^{+}18, HvBKG20]. Assist [CXL17, DDFK19]. Assistance [KFSC18]. Assistant [CLL^{+}21]. Assistants [CR20]. Assisted [JCD19]. Assisting [BBW21]. Assurance [CCS^{+}18]. Asymmetric [OG21, WACdA^{+}21]. Asymmetries [NSJ21]. Asynchronous [CCNY19, CMY^{+}17]. Attach [BR20]. Attention [KCC20, LAHW19]. Attitudes [RA21b]. Attribute [CLP^{+}21]. Attribution [BB18]. Audio [CSS^{+}17, TRO^{+}21]. Audit [HJM20]. Auditory [MRGN20]. Augmentation [BRM21]. Augmented [ALG^{+}21, LZY^{+}21, MAN^{+}20, SCY^{+}18, SHY^{+}18, SLN21, TRO^{+}21, TG18, WO20]. Augmenting [WDP^{+}20]. Author [EHH^{+}18]. Authority [HFL19]. Autogrip [BR20]. Automated [SWT20, TD19, ZMR21]. Automatic [BMPP20, GB18, RA21b]. Automatically [HSG19]. Autonomy [HFL19]. Avatar [DBS^{+}21, KRMM21]. Awareness [CEM17, KTF^{+}21, MRGN20, STM^{+}19, SPT^{+}20, WDP^{+}20]. Away [HFBL19]. Axis [SLT^{+}21, SLT^{+}21].

B [VZV19]. Back [FC20]. Back-of-Device [FC20]. Backgrounds [HWYW21]. Bad [CLL^{+}21]. Badge [BAAP20]. Balancing [KTF^{+}21]. Ban [CPS^{+}17]. Bangladesh [AHCD17, SAR21]. Barrier [UAAR20]. Based [ALG^{+}21, ABQ17, BJ21, BahRKM18, BAAP20, CMN^{+}21, CLwH^{+}20, DHH18, GB18, KMM21, LNM21, RLGG21, See20, SWT20, SSGB21, YHR^{+}19, AV19, BKM20, BCCV17, dACZ19, CHC^{+}17, DBS^{+}21, GH21, HBJR21, HvBKG20, LO17, MTA^{+}17, SDO21, TBR20, VN19, Vya19, YGHR21, ZWZ^{+}21, KKE^{+}21]. Batches [CLP^{+}21]. BCI [HRC^{+}21]. Be [DFJ^{+}21]. begets [MIS^{+}20]. Beginner [WDP^{+}20]. Behavior [CSA^{+}21, HSG19, KYB21, MMdSP21, MBT^{+}20, PCMP18, SWT20, VMP21]. Behavior-Based [SWT20]. Behavioral [CGS^{+}21, MYS^{+}21, RCM21, TD19]. Beidian [CCX^{+}20]. Being [DFJ^{+}21, CCWH21]. Better [TDC17, WB21]. Between [BLH^{+}17, CR20, DWMV21, FCE^{+}18, MSY20, XHH17, BZC^{+}17, CCH^{+}18, LZAD21, PBC20, SWT20, SSNC20, WACdA^{+}21]. Beyond [CYW18]. Bézier [GVNK20]. Biases [AOT^{+}18]. Bikers [BBW21]. Binary [SS21]. Bindings [KKS19]. Biological [MDM^{+}19]. BioNets [MDM^{+}19]. Biophilia [RVBC^{+}21]. Bird [SBAK21]. Birds [VMP21]. Bit [BVM21]. Bitcoin [KPW19]. Black [ERT21, RI20]. Blind [MRGN20]. Blogging [BP17b]. ’Blue [RAZ^{+}20]. Board [SBAK21, TRO^{+}21]. Bodily [GBN21]. Body [GB18, HOSK21, IOES20]. Bolster [STM^{+}19]. Boost [AOT^{+}18]. Borders [AWM21]. Bots [WWSG21, BKJ^{+}20]. bought [IRA^{+}19]. Boundary [CS18b]. Brainstorming [HSG19, LW19]. Break [ZRL21]. Breakdowns [GMM19]. Breaking [UAAR20]. Breathe [ERT21]. breeds [BTM^{+}21]. Brevity [GAW19]. Bridging [KG21]. Bringing [LAR21].

Browser [GGF^{+}20]. Build [HSVR21]. Building [CSN19, GB18, WWT^{+}21]. Built [FKK19, ZRL21]. Built-In [ZRL21]. Businesses [KBFK21]. Button [BLC^{+}21]. Bystander [AFKL20, KTF^{+}21].

Call [HDPL21]. called [BTM^{+}21]. Cameras [Wil20]. Campaign [DDFK19, Kee19]. Campaigns [STM^{+}19]. Can [AOT^{+}18, MNS^{+}21, PKC^{+}21, TCH20, ETE^{+}19]. Cancer [SWF^{+}20]. Candy [BBS19]. Can’t [CPS^{+}17, ERT21]. Canvases [KDH^{+}21]. Capacitive [KIK20]. CAPath [KIK20]. Capture [LAHW19]. Capturing [DRAQ18, JKW^{+}21]. Card [LKG21]. Cards [LKG21]. Care [COKL20, FCR^{+}17, KWdCL^{+}21, RCM21, SWF^{+}20, Vya19]. Career [BKKD21]. Caret [LZY^{+}21]. Caring [MDZ19]. Cars [YMG^{+}20]. Case [ADD21, AKIS21, BB18, CFG^{+}17, CEM17, CHC^{+}17, FB19, Gar21, Gil20, HFL19, KFSC18, KG20, LLDH21, XZL^{+}20, vBGH^{+}19]. CASS [WWT^{+}21]. Caste [VMP21]. Catalogs [AWM21]. Catalysts [SKV^{+}21]. Categorization [CT18]. Causal [BB18, GAW19]. causing [RKNES20]. CAVE [GB18]. CCBL [TDC17]. Centered [KRK^{+}21, TDC17]. Centric [AFKL20, KKE^{+}21, LAHW19]. cesspool [Gil20]. Chains [MHG^{+}21]. Chairs [KFMH17]. Challenge [RAZ^{+}20]. Challenges [AHCD17, CCWH21, CCNY19, DRA^{+}17, FSC21, GSC21, KOOL19, KBFK21, KWdCL^{+}21, LKS21, MMdSP21, SGF20, SHY^{+}18, SWF^{+}20, WWSG21]. Challenging [GRG20]. Change [FKK19, GH21, GYD17, KJH19]. Changes [HGG20, ZBL19]. Changing [PBC20, PCL^{+}21]. Channel [ABQ17, MNP^{+}20, NSJ21, RLGG21]. Characteristics [DMC^{+}20]. Characterization [DDZ18]. Characterizing [MMH21, PM21, SKV^{+}21]. Charity [SS19]. Chat [BKJ^{+}20].

Chat-bots [BKJ^{+}20]. Chatbot [BKM20, LYH20, WWT^{+}21]. Chatbot-based [BKM20]. Chats [HWYW21]. CHI [APW^{+}21, GMM21]. Child [MFM21]. Children [RCM21, YGHR21]. China [LX20, LAW19]. Choice [AOT^{+}18, SSGB21]. Choices [Gar21, HWYW21]. Choke [BKM21]. Chronic [BLH^{+}17]. Citizen [HFL19, JØC^{+}20, MM19, MHG^{+}21]. Citizens [SJ21]. City [SBAK21]. Civil [DRA^{+}17]. Class [VN19]. Classes [BPO21]. Classification [BDSL17, CGM^{+}18]. Cleaner [BVM21]. Clickbaits [CSMG17]. Client [KG21]. Climate [GH21]. Close [FB19]. Close-Knit [FB19]. Closer [LAR21]. Cloud [MV19]. Clouds [AV19]. Clues [AMF18]. Clustering [SPT^{+}20]. Clutch [BKM21]. Co [BUSA^{+}21, BPO21, CSN19, KDEA21, LBM21, MFM21, OG21, SSNC20]. Co-Designing [LBM21, BUSA^{+}21]. Co-Editing [CSN19]. Co-Editors [SSNC20]. Co-learner [BPO21]. Co-located [OG21]. Co-play [MFM21]. Co-Produce [KDEA21]. Coach [CKS18]. Coaching [CKS18]. Cocode [BPO21]. Code [BKJ^{+}20, LPFD21]. Code-mix [BKJ^{+}20]. Coerciveness [BBV19]. Cognitive [CGS^{+}21, HvBKG20, KFSC18, WGGS20]. Cohabitation [XHH17]. Coin [KPW19]. Collaborate [ZMW20]. Collaboration [ALG^{+}21, ALD^{+}20, CCNY19, OR19, SBPP20, SCY^{+}18, WMBO19]. Collaborations [KG21, MWM^{+}19]. Collaborative [BP17a, BLL^{+}21, CXL17, DDZ18, FCR^{+}17, KDH^{+}21, KTBL^{+}20, MMdSP21, MAR^{+}19, MRGN20, MDZ19, SWF^{+}20, TBDB21]. Collaboratively [BHNA18, CNH^{+}17]. Collaborators [ZD19]. Collective [SS21, MNS^{+}21]. College [GSC21]. Collocated [NSJ21]. Coloring [DRAQ18].

Combating [Wil20]. Combining [Hig20]. Command [SB20a]. Commands [PKB^{+}19]. Comment [BGRK17]. Commentator [LUFW20]. Comments [LHPH20]. commerce [CCX^{+}20]. Communal [BHM19]. Communicate [ADD21]. Communication [BB20, BLH^{+}17, CYW18, FLC17, GMM19, HWS21, LOSD20, MM19, MHN^{+}21, NSJ21, OM20, SAR21]. Communities [CILM18, CWB^{+}18, DPA17, FB19, FD20, PSM21, SDF^{+}21, SK20]. Community [CXL17, CCS^{+}18, DAQ17, DMC^{+}20, HMB^{+}21, JVP^{+}19, JKW^{+}21, KJH19, KG20, NPT19, See20, WWT^{+}21, ZD19]. Community-AI [KG20]. Community-Based [See20]. CommunityClick [JKW^{+}21]. Companies [SSH^{+}21]. Company [ADD21]. Comparative [RA21a]. Comparing [VZV19]. Competitive [BKM21, MMdSP21]. Complex [CKP^{+}18, CMY^{+}17, DWZ^{+}20, HRRS20, KDH^{+}21, MRRS19]. Complexity [CSS^{+}17, KMR20, SSNC20]. Component [RH19]. Composite [CYGW21, OG21]. Compromise [RJZ^{+}21]. Computational [WMBO19]. Computer [AM21, Hao21, KTBL^{+}20, LW19, MSY20, SHD21, TD19]. Computer-Mediated [LW19]. Computers [Ame18]. Computing [AMJ^{+}19, BWV20, DWW^{+}17, FLP^{+}21, JVP^{+}19]. Concept [BSS18, DFJ^{+}21]. Conceptual [MCC^{+}18, RS19]. Conceptualizations [RA21b]. Conceptualize [MHG^{+}21]. Concerns [LKS21]. Conditions [BLH^{+}17]. Conduct [LPFD21]. Conducting [BUSA^{+}21, BV21]. Conductive [KIK20]. Configurable [ZMR21]. Configuration [TG18]. Conflicts [CT18]. Conformity [FTP20]. confrontation [RKNES20]. Congress [MM19]. Connect [ZD19]. connection [HFBL19]. Consequences [BDSL17, XZL^{+}20]. Considerations [CCNY19, Fie19, VMD^{+}20]. Considering [WGG20]. Consistencies [BL18]. Consistency [SWT20, SSNC20]. Conspiracy [PSM21]. Constrained [NMBS21]. Constraints [RIK21]. Construct [BS21]. Constructed [BHNA18]. Constructionist [Ame18]. Consumption [CSMG17]. Contagion [LNM21]. Content [CWB^{+}18, FTP20, MAN^{+}20, RJZ^{+}21, RS19, See20, VXHK21]. Contestability [VXHK21]. Context [MH19, MYS^{+}21, SAR21, TDC17, TCH20, TVS17, YHR^{+}19]. Context- [YHR^{+}19]. Contexts [DHH18, LM20, PBC20]. Continued [KLS21]. Continuum [DFJ^{+}21]. Contribute [DMC^{+}20, GPH21]. Contributor [BZC^{+}17]. Control [HRC^{+}21, HGG20, LZAD21]. Controlled [COKL20, YMG^{+}20]. Controllers [KMM21]. Controls [GSK^{+}17]. Conversation [CMP^{+}21, LKS21]. Conversational [CR20, QGB20, SPT^{+}20]. Conversations [BGRK17, LPFD21, SKV^{+}21]. Cool [BJ21]. Cooperation [Ame18]. Cooperative [ALD^{+}20, BF17, Pri19]. Coordinates [RRF20]. Coordination [Pri19, RCM21]. Coping [dQPMdHRdAN21, TSHK20]. COPSE [MTA^{+}17]. Copy [LZY^{+}21]. Copy-and-Paste [LZY^{+}21]. Copyright [FB19]. Cord [BHNA18]. Coronavirus [Wil20]. Corporate [LRSS20]. Correctness [SSNC20]. Cost [MCYZ19]. Costs [HWS21, LRSS20]. Couch [SS19]. Counseling [KG21]. cousin [IRA^{+}19]. COVID [HSVR21, YGHR21]. COVID-19 [HSVR21, YGHR21]. Craft [Vya19]. Craft-based [Vya19]. CRDT [SSNC20]. Created [CNH^{+}17]. Creating [AL21, KDH^{+}21, KKS19]. Creative [KBFK21]. Creativity [CILM18, FB19, LW19]. Credibility [BZSM20]. Crisis [CH20]. Criteria [BZSM20]. Critical [Ame18, BBS19, RI20]. cro [WGN^{+}21]. Cross [LE21, PCL^{+}21, RDS18, SHY^{+}18, VN19]. Cross-Device [RDS18, SHY^{+}18, VN19, LE21]. Cross-disciplinary [PCL^{+}21]. Crossing [FLC20]. Crowd [BB18, CILM18, CKS18, CMP^{+}21, GPH21, MBT^{+}20, OTH20, TSS21]. CrowdCog [HvBKG20]. CrowdFolio [FKDG21]. Crowds [FM21]. Crowdsourced [AS17, BZSM20, CSS^{+}17, FGG17]. Crowdsourcing [HvBKG20, MMdSP21, QGB20, WYY^{+}21, vBGH^{+}19]. CrowdUI [OTH20]. Crowdwork [WMM^{+}19]. Crowdworker [SGF20]. Crowdworkers [DHH18, WMM^{+}19]. CSCW [BL17, CPB18, ERT21, WGG20]. CSCW1 [BKL^{+}21]. CSCW2 [KLS20, BLS21]. CSCW3 [KLS21]. Cues [AMR21, BLC^{+}21]. Culture [DRA^{+}17, IRA^{+}19, VKI21, Vya19]. Cultures [LHT^{+}20]. Curious [CEM17]. Curveball [MMM^{+}21]. Curves [GVNK20]. Customizable [MPB^{+}19]. Customization [dACZ19, DBS^{+}21]. Customizations [GMM19]. Customize [dQPMdHRdAN21]. cutting [Hig20].

D [CPB18, KIK20, OSM^{+}20, SEC^{+}20]. D-Printed [KIK20]. DANTE [SPT^{+}20]. Data [AL18, CFD^{+}21, CCS^{+}18, EJB^{+}20, FSN^{+}20, Fie19, FCE^{+}18, FSC21, GYD17, GSK^{+}17, HOZ^{+}21, HRRS20, KKE^{+}21, KG21, LLDH21, LE21, MWM^{+}19, MFM^{+}20, MDZ19, MSY20, MNS^{+}21, MRRS19, RA21b, SB17, SLT^{+}21, TBDB21, WMBO19, YHR^{+}19, ZMW20, FSC21]. Data-Centric [KKE^{+}21]. Data-Driven [CFD^{+}21, CCS^{+}18, YHR^{+}19]. Dataset [SHD21]. Datasets [SHD21]. Daughters [RYD21]. Decals [NMBS21]. December [KLS21]. Decision [GSC21, GPH19, GC19, KOOL19]. Decision-Making [KOOL19]. Decisions [AMR21]. Deck [LKG21]. Declarative [DS20]. Decomposed [FKDG21]. Decreasing [HRC^{+}21]. Deep [AH18, CXL17, LYH20, SPT^{+}20]. Deeper [KML^{+}21, VH21]. Defined [CLwH^{+}20, MLSH21]. Deformable [NMBS21]. Delayed [HKF19]. Delays [HKF19]. Dementia [KRK^{+}21]. Democracy [MDZ19]. Demography [SS21]. Demonstrating [MPB^{+}19]. Departing [BSS18]. Depression [BRN^{+}19, SWF^{+}20, ZBL19]. Depth [Wil20]. Description [MNP^{+}20, SGF20]. Design [BKM20, BAAP20, BWDP20, BDSL17, BDW21, BS21, BP17b, BV21, CLKEF21, CMP^{+}21, CCNY19, DMC^{+}20, EJB^{+}20, FGG17, GN20, Hao21, JCD19, KRK^{+}21, KWdCL^{+}21, LDP18, LKG21, MRGN20, NPT19, OTH20, RI20, VZV19]. Designers [SLN21]. Designing [BCCV17, CRN21, FKK19, FLP^{+}21, GHHC21, LYH20, LBM21, MRRS19, NKV21, VN19, YMG^{+}20, ZD19, BUSA^{+}21]. Designs [AFKL20, BLC^{+}21, CYGW21]. Desk [LAHW19]. Desk-Centric [LAHW19]. Detecting [CSA^{+}21]. Detection [PCMP18, TRO^{+}21, ZBL19]. Developer [LLDH21]. Developers [LLDH21, MHG^{+}21]. Developing [HBJR21]. Development [SHD21, SK20, SHY^{+}18]. Device [FC20, HDPL21, HOSK21, LAHW19, RDS18, SLT^{+}21, SHY^{+}18, VN19, LE21]. Devices [BR20, SCMS18]. DG3 [DS20]. Diagnosis [GHHC21, MT21]. Diagram [VN19]. Diagramming [LE21]. Diary [ABC21]. Did [BZC^{+}17]. Differences [BZSM20, DMC^{+}20, SSNC20]. Differentiating [CYC^{+}21]. Difficulty [Hig20, RGR^{+}21]. Digital [AHCD17, AMR21, AFT^{+}18, CR20, DBS^{+}21, FKK19, Hao21, HFBL19, HGG20, LO17, MFM21, SS21, TAS^{+}19, ZSOJ20]. Digitization [MAR^{+}19]. Dimension [SLT^{+}21]. Dimensional [YRIV20, UYM21].

Dis [HFBL19]. Disabilities

[HBJR21, Tan21]. Disaster [BL18]. Disciplinary [SHD21, PCL^{+}21]. Disclosure [DRA^{+}17, LYH20]. Disclosures [AMF18, ERB^{+}17]. Disconnects [TBDB21]. Discourse [ASS18, GKS21, Hao21, TAS^{+}19]. Discovery [BP17a]. Discussion [CMY^{+}17]. Discussions [XZL^{+}20]. Disorders [FTP20, TBDB21]. Display [WACdA^{+}21, ZMZL21]. Displays [RRF20]. Distributed [BKM20, MV19, SBPP20]. Distribution [GSK^{+}17, YU20]. Disturb [WWSG21]. Disturbance [TCH20]. Dive [KML^{+}21]. Diverse [KTBL^{+}20]. Diversifies [CYW18]. Diversity [dQPMdHRdAN21]. Djnn [MCC^{+}18]. Djnn/Smala [MCC^{+}18]. Do [BKJ^{+}20, GPH19, RJZ^{+}21, SHD21, MHG^{+}21, ZMW20]. Does [FKK19]. Dog [HDPL21]. Dog-to-Human [HDPL21]. Domain [DWZ^{+}20, MWM^{+}19, Rib19]. Domain-independent [DWZ^{+}20]. Domestic [RYD21]. Donations [SS19]. Don’t [AMR21, CEM17, TSHK20, WWSG21]. Down [IOES20]. Drive [KPW19]. Driven [CFD^{+}21, CCS^{+}18, MMM^{+}21, YHR^{+}19]. Driver [DC17, KMR20]. Drivers [MH19]. Drone [HMB^{+}21]. Dropout [MMH21]. Drowsiness [KMR20]. Dual [NSJ21, YU20]. Dual-channel [NSJ21]. During [GSC21, Hig20, Kee19, LHPH20, YGHR21, HSVR21]. Dynamic [BZC^{+}17, GBN21, HBJR21, NMBS21]. Dynamics [Kee19, MSY20].

E-commerce [CCX^{+}20]. Ear [CLwH^{+}20]. Ear-Based [CLwH^{+}20]. Early [KYB21, SLD21]. easier [RKNES20]. Eating [FTP20, PBC20, TBDB21]. Ecological [BahRKM18, KBKH21]. Economy [DWW^{+}17, NKV21]. Ecosystem [DWB18]. Ecosystems [GMM19]. Edit [MD20]. Edit-a-thons [MD20]. Editing [CXL17, CSN19, DDZ18]. Editor [KFMH17]. Editorial [BKL^{+}21, BLS21, BMdS^{+}21, CNC^{+}17, CM20, FJIH21, GMM21, KLS20, KLS21]. Editors [CT18, SSNC20, EVC^{+}19, KMHLF18, LGS19]. Education [AMJ^{+}19, CCWH21, LKS21]. Educational [KRMM21]. Educators [AL21, PM21]. Effect [AMR21, ERLW17, GRG20, MRGN20, OM20]. Effects [ALG^{+}21, ASP20, AOT^{+}18, BBNT21, CSS^{+}17, GAW19, HWYW21, KRMM21, KYB21, MCYZ19, MHN^{+}21]. Efficacy [CPS^{+}17]. Efficient [BB18, SBPP20]. Effort [GBN21]. EICS [JVP^{+}19]. Elderly [PCMP18]. Electroencephalography [KFSC18]. Elements [GBN21, NMBS21]. Elicitation [MV19, WO20]. Eliciting [SB20a]. elite [GN20]. Embody [SLT^{+}21, KKE^{+}21]. Emerging [HMB^{+}21]. EMG [KKE^{+}21]. EMG-Based [KKE^{+}21]. Emotion [BKM20, GKS21, RA21b]. Emotional [GBN21, LHPH20]. Emotions [FCE^{+}18, GBN21]. Empathy [WB21, ZAS^{+}21]. Empirical [ALG^{+}21, CSJ^{+}18, JCD19, KML^{+}21]. Employees [LRSS20]. Empower [MNS^{+}21]. Emulation [CLKEF21]. Enable [HKM^{+}20]. Enabled [RA21b, SWF^{+}20, GHHC21]. Enablers [RYCC21]. Enabling [BR20, FLC20, WACdA^{+}21]. End [BR20, BahRKM18, dQPMdHRdAN21]. End-User [BR20, BahRKM18]. End-users [dQPMdHRdAN21]. Ended [BP17a]. Endoscopic [MON^{+}21]. Enforcement [FB19]. Engage [KDEA21]. Engagement [AKIS21, DAQ17, DC17, GH21, GBN21, JØC^{+}20, MMH21, PSM21, RGR^{+}21]. Engaging [FCE^{+}18]. Engine [ERLW17, VH21]. Engineering [BWV20, CMN^{+}21, FLP^{+}21, JVP^{+}19, SVV21]. English [LHT^{+}20]. Enhancing [LAR21, ZSOJ20]. enjoy [CCWH21]. Enough [CNH^{+}17]. Ensuring [SWT20]. Entering [GSK^{+}17]. Enterprise [CWB^{+}18]. Entertainment [GH21]. Entity [RH19]. Entity-Component-System [RH19]. Entrepreneurs [DDFK19]. Entry [ASP20]. Environment [SCY^{+}18, YOFC19]. Environments [GB18, LAHW19, SVW19, SBPP20, TD19]. Ephemeral [EJB^{+}20]. Ephemerality [MCYZ19]. Epilepsy [MT21]. Episodic [KSX21]. Epistemic [AS17]. Equity [EHH^{+}18]. Era [CSMG17]. Ergonomics [AFT^{+}18]. Errors [ASP20, CLL^{+}21]. Escalator [AKIS21]. Esports [LUFW20]. Estimating [QGB20]. Estimation [KCC20]. Ethical [BUSA^{+}21, BS21, Fie19]. EU [dACZ19]. Evaluate [LHI20]. Evaluating [DHH18, KMM21, RAZ^{+}20, WGGS20, YOFC19]. Evaluation [AFT^{+}18, BLC^{+}21, Hig20, RA21a]. Evaluations [DRK18]. Event [CH20, CHC^{+}17]. Event-based [CHC^{+}17]. Events [LHPH20, PKB^{+}19]. Ever [MNS^{+}21]. every [BLH^{+}17]. Everyday [BB20, XHH17]. Evidence [LHPH20, RJZ^{+}21]. evil [BTM^{+}21]. Evolution [CNH^{+}17]. Evolving [MT21]. Examined [CPS^{+}17]. Examining [ABC21, ASS18]. Example [KMR20]. Exchange [RDS18, WB21]. Execution [WYY^{+}21]. Exercising [HFL19]. Exergame [BRM21, KRK^{+}21]. Exhibits [LBM21]. Existing [TRO^{+}21]. Expanded [CSMV21]. Expanding [YRIV20]. Expectations [SSH^{+}21]. Experience [ALG^{+}21, BVM21, GBN21, LAR21, POG^{+}21, Tan21]. Experiences [AMF18, CSN19, HBJR21, JSP^{+}21, YGHR21]. Experiencing [LAW19]. Experiment [JCD19, OR19]. Experimentation [KKE^{+}21]. Experiments [AOT^{+}18, BV21]. Expert [BZSM20, FGG17, NHH20]. Experts [MWM^{+}19]. Explainable [GLZ^{+}21]. Explanation [KG20]. Explanations [GLZ^{+}21]. Exploiting [DS20]. Exploration [BB18, RVBC^{+}21, RDS18, WGM21, ZSOJ20]. Exploratory [CCX^{+}20]. Explore [GGF^{+}20]. Exploring [BKM21, BLC^{+}21, BAAP20, CLKEF21, CLwH^{+}20, EJB^{+}20, JZB^{+}21, JSP^{+}21, KJH19, LX20, MDM^{+}19, NSJ21, WGG20, XZL^{+}20]. Exposure [DBS^{+}21]. Expression [GMM19]. Expressions [GBN21, SB20a]. Extend [KIK20]. Extended [BF17]. Extending [dACZ19]. Extract [NM21]. Extraction [HMB^{+}21]. Extremist [PM21]. Eye [KCC20, WGN^{+}21, WGN^{+}21]. Eye-gaze [WGN^{+}21]. Eyes [AH18].

Fabric [VMD^{+}20]. Fabricated [HRC^{+}21]. Fabrication [Ann20]. Facebook [BHM19, DWB18, SSGB21]. Facial [GBN21]. Facilitated [CYW18]. Facilitating [CMY^{+}17]. Facilitation [YGHR21]. Factors [AFT^{+}18, PSM21]. Faculty [LKS21]. Fair [vBGH^{+}19]. Fairness [BKG^{+}21]. Fakey [MAMP21]. Familie [Gar21]. Family [IRA^{+}19, JA21, MFM21, PBC20]. Fandom [FD20]. Farm [HMB^{+}21]. Fatherhood [JAB19]. Fatigue [KMR20]. Fear [DBS^{+}21]. Feature [BDL21]. Features [ASP20, BF17, MDM^{+}19]. Feedback [BR20, BBW21, CMP^{+}21, DC17, FLC17, FGG17, FKDG21, HOSK21, JKW^{+}21, KBFK21, KYB21, RLGG21, TRO^{+}21, WB21, YHR^{+}19, YOFC19]. Feedforward [CLV^{+}19]. Feminist [RI20]. Fertility [FCR^{+}17, FCE^{+}18, FSC21]. Few [AH18]. Fidelity [SLN21]. Financial [PKC^{+}21]. Find [BKJ^{+}20, MWM^{+}19]. Finding [CCH^{+}18]. Findings [WGGS20]. Fine [DWZ^{+}20, NB20]. Fine-grained [DWZ^{+}20, NB20]. Finger [RLGG21, YRIV20]. Fingers [BR20]. First [BMdS^{+}21]. Fixed [JCD19]. Fixed-Gaze [JCD19]. Fixes [GGF^{+}20]. Flamers [PM21]. Flecto [KVDG20]. Flex [LHI20]. Flex-ER [LHI20]. Flexible [BMPP20, SCMS18]. Flicks [YRIV20]. Flight [HGG20]. Flood [SLD21]. Fluid [LE21]. Foldable [KVDG20]. Folk [KDEA21]. Fomites [Wil20]. Food [PBC20]. Force [BR20, RA21a]. Forces [HOSK21]. Formal [CMN^{+}21]. Formation [FB19]. Forming [HDPL21]. Forms [UAAR20, JØC^{+}20]. Fortunettes [CLV^{+}19]. Forum [LLDH21]. Forums [BTM^{+}21]. Foster [GYD17]. Four [RA21a]. Fragmentation [WMM^{+}19]. Frames [KJH19]. Framework [CGM^{+}18, HWS21, HRRS20, KE18, MCC^{+}18, OG21, RI20, RDS18, RS19, SCMS18, TRO^{+}21]. Framing [AOT^{+}18, DRK18, KG21, WB21]. Franca [CYW18]. Freeform [NMBS21]. Freelancers [BKKD21, LRSS20]. Friction [SB20b]. Friends [CEM17, SK20]. Frustration [HRC^{+}21]. Full [GB18]. Full-Body [GB18]. Fun [BZC^{+}17]. Function [LW19]. Future [BWV20, BDW21, CLV^{+}19, GSC21, BL17].

G [CS18a, HOSK21]. G-Forces [HOSK21]. G-Gene [CS18a]. Game [AL21, ALD^{+}20, BP17a, CGS^{+}21, DBS^{+}21, FcLDM19, GH21, GBN21, Hao21, KML^{+}21, LOSD20, MAMP21, MFM21, RI20, RGR^{+}21, SBAK21]. Game-based [DBS^{+}21, GH21]. Gameful [LNM21]. Gamemasters [AL21]. Gameplay [CSA^{+}21]. Gamers [TRO^{+}21, ZRL21]. Games [BKM21, DFJ^{+}21, GH21, GHHS21, Hao21, HRC^{+}21, KRMM21, KMM21, KBKH21, KML^{+}21, MHG^{+}21, RVBC^{+}21, TRO^{+}21, VKI21]. Gamification [LNM21]. Gaming [BJ21, MFM21, ZRL21]. Gaps [BDW21, DMC^{+}20]. Gastroenterological [MON^{+}21]. Gather [SS19]. Gaussian [YU20]. Gay [DWB18]. Gaze [JCD19, KCC20, MON^{+}21, WGN^{+}21].

Gelicit [MV19]. Gender [BKG^{+}21, DMC^{+}20, EHH^{+}18, MWH19, MGG^{+}21, SS21]. Gender-Imbalanced [MWH19]. Gene [CS18a, CS18a]. Generating [BR17, SSGB20]. Generation [CFG^{+}17, DS20, RS19, SDO21]. Generic [CPB18]. Genetic [Sal17, SDO21]. Genre [AWM21]. Geo [WYY^{+}21]. Geo-Social [WYY^{+}21]. Gestatten [KCC20]. Gesto [PKB^{+}19]. Gesture [AFT^{+}18, DS20, DWZ^{+}20, GVNK20, HG19, KCC20, MV19, WO20, YMG^{+}20]. GestureCards [HG19]. Gestures [CS18a, CLwH^{+}20, FLC20, LW19, MLSH21, OSM^{+}20, PKB^{+}19]. Get [TSHK20]. Girls [KBKH21]. Github [LPFD21]. Glimpse [BWV20]. Go [CYC^{+}21, GPH19, IRA^{+}19]. Goal [KYB21]. Going [LAW19]. Golfers [WDP^{+}20]. good [BLH^{+}17]. Got [BVM21]. Governance [SDF^{+}21]. Gradual [ZRL21]. grained [DWZ^{+}20, NB20]. Graphic [DMC^{+}20]. Graphical [KVDG20, NMBS21, SDO21, SVV21]. Graphics [ZSOJ20]. Grassroots [AL18, NPT19]. Gratifications [LX20]. Grid [KIK20, SDO21]. Grid-based [SDO21]. Grief [BHM19]. Ground [KMR20]. Grounding [CR20, HWS21]. Group [DPA17, DRK18, KWdCL^{+}21, LW19, VKI21, ZBL19, BMdS^{+}21]. Groups [HSVR21, MWH19, MAR^{+}19, YGHR21]. Groupware [MRRS19]. Growth [CKS18]. GUI [BV21, CLV^{+}19]. Guide [JVP^{+}19]. Guided [HRRS20, WGM21]. Guidelines [RAZ^{+}20]. Guiding [WGM21].

Hackathons [NHH20]. Hackers [Ame18]. Halls [JKW^{+}21]. Hand [CSMV21, FC20, LO17, LW19, ZSOJ20]. Hand-based [LO17]. Handed [FC20]. Handling [GVNK20]. Handwritten [ZWZ^{+}21]. Haptic [HOSK21, YMG^{+}20]. Harassment [BDSL17, SSH^{+}21]. Harmful [MBT^{+}20]. Hate [CPS^{+}17, MIS^{+}20]. HCI [CM20, ERT21, NKV21, PCL^{+}21, WGG20]. Head [HOSK21, JCD19]. Head-Worn [HOSK21]. Health [BLL^{+}21, FCE^{+}18, LYH20, MYS^{+}21, RCM21, WWT^{+}21]. Healthcare [BLH^{+}17, FSC21]. HeartMob [BDSL17]. Heavy [ZRL21]. Helmet [HOSK21]. Help [ETE^{+}19]. Helpful [SGF20]. Here [CPS^{+}17]. Heterogeneous [CSN19, HvBKG20]. Hexa [BLC^{+}21]. Hexa-metric [BLC^{+}21]. Hidden [CSJ^{+}18]. Hide [HWYW21]. Hierarchies [VMP21]. High [CYC^{+}21]. Highly [FCR^{+}17]. Hiring [LRSS20]. Historical [BDW21, Gil20]. History [Ame18]. HistoryViewer [SB17]. HMD [WACdA^{+}21]. Hoaxes [AH18]. hoc [SCY^{+}18]. Holistic [FKDG21, LE21]. Home [TDC17]. Homepreneurship [Ann20]. Homes [SAE^{+}21]. Honor [CSA^{+}21]. Hope [SGF20]. Hospital [HKM^{+}20]. Households [WGG20]. Human [AFT^{+}18, ALD^{+}20, BBV19, CGM^{+}18, CLP^{+}21, GYD17, HDPL21, HOSK21, KRK^{+}21, KTBL^{+}20, MMM^{+}21, POG^{+}21, TD19, GHHS21]. Human-AI [ALD^{+}20, GHHS21]. Human-Centered [KRK^{+}21]. Human-Computer [KTBL^{+}20, TD19]. Human-Machine [CGM^{+}18]. Human-side [CLP^{+}21]. Humanitarian [DC17]. Humanities [LKS21]. Hurry [HKF19]. Hybrid [HG19, NSJ21]. HybridPointing [DWMV21]. Hyper [DRA^{+}17]. Hyper-Selective [DRA^{+}17]. Hypotheses [AV17].

ICT [ETE^{+}19, TAS^{+}19]. ICTs [DRA^{+}17]. IDE [SB20a]. Ideation [KTBL^{+}20]. Identification [SJ21]. Identifying [DPA17, KFSC18]. Identities [SS21]. Identity [DWB18, KDEA21]. Ideological [KPW19]. ideologies [GN20]. Image [BKG^{+}21, CLP^{+}21, MGG^{+}21, SGF20]. ImageSense [KTBL^{+}20]. Imaginaries [HMB^{+}21]. Imbalanced [MWH19].

Immersion [KTF^{+}21]. Immersive [HOZ^{+}21, KRK^{+}21, LHI20, SEC^{+}20]. Immigrants [LM20]. Impact [CLP^{+}21, FC20, MH19, MCI20, MGG^{+}21, MP19]. Impacts [DRK18]. Impairments [JSP^{+}21, TRO^{+}21]. Imperfect [GYD17]. Implications [GN20, Hao21, JCD19, NPT19]. Importance [VH21]. Imposition [MSY20]. Improve [JKW^{+}21, MAMP21, WB21]. Improvement [Hig20]. Improves [CLL^{+}21]. Improving [MWH19, SPT^{+}20]. In-Situ [KFSC18]. Inaccurate [ABC21]. Incels [RJZ^{+}21]. Incentives [STM^{+}19]. Inclusion [IRA^{+}19]. Inclusive [HBJR21, RI20, TRO^{+}21]. Inclusivity [JKW^{+}21]. Income [MAR^{+}19]. Increase [WMM^{+}19]. Independence [BHNA18]. independent [DWZ^{+}20]. Index [YRIV20]. Indian [VMP21]. Indicating [ERB^{+}17]. Indicators [GPH21]. Indirect [AMF18]. Individual [HBJR21]. Individuals [BRN^{+}19]. Industrial [HMB^{+}21, TG18]. Industry [KG21]. Infer [TG18]. INFEX [RDS18]. Influence [BJ21, BDL21, FKDG21]. Influencers [LNM21]. Informal [LM20, LBM21]. Informatics [EJB^{+}20]. Information [ASS18, Kee19, KOOL19, MM19, NM21, NPT19, RDS18, VN19]. Informational [SB20b]. Infrastructure [BBV19, BBS19, SJ21]. Infrastructures [JAB19]. Ingroup [WB21]. Initiative [CCH^{+}18]. Injury [BHNA18]. Inmates [TAS^{+}19]. Innovation [KJH19]. Input [BAAP20, GSK^{+}17, HRC^{+}21, SCMS18, YRIV20, ZMZL21]. Inputs [KIK20]. Insights [BDSL17]. Inspection [MCI20]. Instant [CKP^{+}18, COKL20, HKF19, YHR^{+}19]. Instantiating [MTA^{+}17]. Institutions [FKK19]. Instructions [KMM21]. Instrumenting [SB17]. Integer [LZS^{+}21]. Integrating [SCMS18]. Integration [OR19]. Intelligent [CFG^{+}17, CLL^{+}21, KTBL^{+}20]. Intentions [SB20a]. Inter [DPA17]. Inter-Group [DPA17]. Interactants [SPT^{+}20]. Interacting [WWSG21]. Interaction [AV17, AM21, BLC^{+}21, CYC^{+}21, CSMV21, CLL^{+}21, GGF^{+}20, GB18, KG20, LAR21, LHI20, MCC^{+}18, MAN^{+}20, RRF20, SVW19, TD19, TBR20, UAAR20, WACdA^{+}21]. Interaction-Oriented [MCC^{+}18]. Interactionism [PBC20]. Interactions [CLwH^{+}20, MDM^{+}19, RH19, WO20, ZBL19, GHHS21]. Interactive [AS17, BWV20, BDW21, CMN^{+}21, DDFK19, FBG17, FLP^{+}21, Gar21, JVP^{+}19, KE18, KKS19, LKG21, SB17, TBR20, WGGS20, WGM21, ZMR21]. Intercultural [AMJ^{+}19]. Interface [CMY^{+}17, KKE^{+}21, MNP^{+}20, Sal17, VZV19, YHR^{+}19, ZSOJ20]. Interfaces [AFT^{+}18, CLKEF21, CPB18, GLZ^{+}21, KIK20, KVDG20, MTA^{+}17, MNP^{+}20, NMBS21, dQPMdHRdAN21, PCL^{+}21, RH19, SVV21, SLN21, VN19, XHH17, YMG^{+}20]. Intergenerational [LOSD20, PBC20]. Intermediaries [CCX^{+}20]. Intermediate [CRN21]. International [ADD21]. Internet [CSJ^{+}18, HDPL21]. Interpersonal [MHN^{+}21, TSHK20]. Interplay [AS17, BZC^{+}17, FCE^{+}18]. Interpretations [HKF19]. Interruption [SBPP20]. Interruptions [WGN^{+}21]. Intervention [MAMP21, ZRL21]. Interventions [JZB^{+}21, RA21b]. Intimate [BTM^{+}21]. Intrinsic [LW19]. Introduction [BMdS^{+}21]. Investigating [AMR21, BZSM20, CILM18, CSS^{+}17, IOES20, MWH19, POG^{+}21, RVBC^{+}21]. Investigation [BWDP20, VH21]. Invisible [TSS21]. Involving [Fie19]. IRL [SK20]. Islamic [IRA^{+}19]. ism [GHHS21]. ISS [CM20, FJIH21]. Issues [CSN19]. IVR [BP17b].

Jam [FcLDM19]. Jams [FcLDM19]. JavaFX [FBG17]. Join [PSM21]. Joint [HWS21]. Journey [SAE^{+}21]. Journeys [SWF^{+}20]. Just [TSHK20]. Justice [SSH^{+}21]. Justifications [BTM^{+}21]. JustSpeak [ZMR21].

KAVE [GB18]. Kenyan [WGG20]. Key [BLC^{+}21, BBV19]. Keyboard [BB20, CSMV21]. Keyboards [ASP20]. Kick [PKC^{+}21]. Kinect [GB18]. Knit [FB19]. Knitted [VMD^{+}20]. know [BLH^{+}17]. Knowledge [ABC21, KDEA21, OR19]. Korero [CMY^{+}17].

Labeling [CLP^{+}21]. Labor [TSS21]. Landing [Sal17]. Lands [FD20]. Language [BB20, CYW18, HSG19, MCC^{+}18, MHN^{+}21, MNP^{+}20, TDC17, ZAS^{+}21]. LanguageLogger [BB20]. Laptop [CSMV21]. Large [BB18, DWMV21, FLC20, RRF20, WACdA^{+}21]. largest [Gil20]. Last [BZC^{+}17]. Layer [SDF^{+}21]. layered [TVS17]. Layout [KIK20, SDO21]. Layouts [LZS^{+}21, SEC^{+}20]. Lead [AOT^{+}18]. Leakage [CHC^{+}17]. Learn2Earn [STM^{+}19]. Learned [GHHC21, KRK^{+}21, Rib19]. learner [BPO21]. Learners [DHH18]. Learning [Ame18, AMJ^{+}19, CXL17, DHH18, FcLDM19, GLZ^{+}21, JØC^{+}20, JA21, LM20, LBM21, LX20, MON^{+}21, RIK21, YGHR21, vBGH^{+}19]. Led [KG20]. Lens [SB20b, NKV21]. Lenses [GKS21]. Lessons [AMJ^{+}19, GHHC21, KRK^{+}21]. Let [BKJ^{+}20]. Letter [SN18]. Levels [Hig20]. Leveraging [LM20]. LGBTQ [DWB18]. LibraryThing [AWM21]. Life [AV19, GSC21, LOSD20, SBAK21, WMM^{+}19]. Life-Tags [AV19]. Light [HFBL19]. Lightweight [DAQ17, JZB^{+}21]. Like [SJ21]. Limits [GC19]. Line [CYGW21, HBJR21, VZV19]. Lingua [CYW18]. Linguistic [ERB^{+}17]. Links [DPA17, DRAQ18, VH21]. Listen [FLC17]. Literacy [LBM21, MAMP21]. Literature [APW^{+}21, DWW^{+}17, DFJ^{+}21]. Little [AM21, BVM21]. Live [CCWH21, LUFW20, LHPH20, SS19]. Live-Streaming [LHPH20]. Livestreaming [LAW19, SK20]. Living [CWB^{+}18, CS18b]. Loads [NB20]. Local [ZD19]. located [OG21]. Location [BJ21, LNM21]. Location-Based [BJ21, LNM21]. Logo [Ame18]. Long [CWB^{+}18, HFBL19]. Long-Term [CWB^{+}18, HFBL19]. Longitudinal [BKKD21]. looking [Pri19]. Loop [GC19, KG21]. Low [CYC^{+}21, LAR21, MAR^{+}19, SJ21]. Low-Income [MAR^{+}19]. Low-resolution [SJ21]. Low-Vision [LAR21]. Lying [IOES20].

M [WGN^{+}21]. Machine [CGM^{+}18, GLZ^{+}21, vBGH^{+}19]. Macro [CSJ^{+}18]. MADE [SLT^{+}21]. MADE-Axis [SLT^{+}21]. Magic [KG21]. Maintain [RKNES20]. Maintenance [SSNC20]. Maker [MWH19, Vya19]. Makers [MWH19]. Makerspaces [LZAD21]. Makes [PSM21]. Making [BKKD21, DDFK19, GYD17, GPH19, GC19, KOOL19]. Mall [AKIS21]. Management [BKM20, BL17, CRN21, GB18, NSJ21, SBPP20]. Managing [BRN^{+}19]. Mandarin [LHT^{+}20]. Mani [LO17]. Mani-Pull-Action [LO17]. Manifest [VMP21]. Manipulating [HRC^{+}21]. Manipulation [CFG^{+}17, ERLW17, MMM^{+}21, WO20]. Manual [KFSC18]. Many [CT18]. Mapping [DC17, PKB^{+}19]. Mappings [ZMZL21]. Maps [NM21, SEC^{+}20]. Marital [ADAY21]. Markers [ERB^{+}17]. Matching [ZD19]. Matchmaking [ADAY21]. Materialities [MM19]. Materials [CMY^{+}17]. Mature [LM20].

Maximization [WYY^{+}21]. MDE [LDP18]. Me [SEC^{+}20, WWSG21, IRA^{+}19]. Means [LLDH21]. Measurement [APW^{+}21]. Measurements [CYGW21]. Measures [GBN21]. Measuring [ALD^{+}20, HJM20]. MechanicalHeart [CGM^{+}18]. Mechanics [KML^{+}21]. Media [AMF18, BSS18, BWDP20, CSMG17, CH20, DWB18, EJB^{+}20, ERB^{+}17, GAW19, GSC21, GPH19, GPH21, GN20, JZB^{+}21, KBFK21, MAMP21, RKNES20, RA21b, RAZ^{+}20, SSH^{+}21]. Mediated [LW19, MHN^{+}21]. Mediating [KG20]. Mediator [LYH20, VKI21]. Medium [SLN21]. Medium-Fidelity [SLN21]. Meeter [HSG19]. Meetings [KDH^{+}21]. meets [RYCC21]. Meetup [CHC^{+}17]. Members [HMB^{+}21, PBC20]. Memory [JA21, KSX21, WGGS20, WGM21]. Men [RYD21]. Mental [LYH20]. Mentoring [NHH20]. Menus [BCCV17, FLC20]. Merlynne [CGS^{+}21]. Meso [CSJ^{+}18]. Message [KMHLF18, LGS19]. Messaging [ABQ17, CKP^{+}18, COKL20, HKF19, MFM^{+}20, RAZ^{+}20]. Meta [OR19]. Meta-Knowledge [OR19]. Metaphoric [LOSD20]. Method [CS18a, WGM21]. Methods [ABC21, BDL21, GB18, SSGB20]. metric [BLC^{+}21]. Mexico [GYD17]. Micro [CSJ^{+}18]. Microgestures [WGN^{+}21]. Microtask [QGB20]. Mid [BLC^{+}21, YMG^{+}20]. Mid-Air [YMG^{+}20, BLC^{+}21]. Midair [ZMZL21]. middle [LRSS20]. Migration [FD20, KWdCL^{+}21]. Migrations [RJZ^{+}21]. Mile [ETE^{+}19]. Military [DRA^{+}17]. Mind [AMR21]. Miniaturized [LZY^{+}21]. Mining [TD19]. Misaligned [DPA17, TBDB21]. Misfires [TBDB21]. Misinformation [HJM20, JZB^{+}21]. Misinterpretations [CKP^{+}18]. Missed [TBDB21]. mitigating [GN20]. mix [BKJ^{+}20]. Mixed [ABC21, CCH^{+}18, Pri19]. Mixed-Methods [ABC21]. ML [ADD21]. Mobile [AHCD17, ASP20, ABQ17, BB20, CEM17, COKL20, GRG20, KCC20, KFSC18, LZY^{+}21, SCY^{+}18, STM^{+}19, TCH20, WYY^{+}21, YRIV20, IRA^{+}19]. MoCaDiX [VN19]. modal [BLC^{+}21]. Model [BCCV17, CMN^{+}21, LAHW19, RH19, RS19, TBR20, YU20]. Model-Based [CMN^{+}21, BCCV17, TBR20]. Modeling [BDL21, TG18, UYM21, Wil20]. Modelling [BDW21, TVS17]. Models [ADD21, AFT^{+}18, AM21, BR17, BF17, CFG^{+}17, CMN^{+}21, CNH^{+}17, DS20, PCMP18, See20, SWT20]. Moderating [Gil20]. Moderation [FTP20, KJH19, KG20, RJZ^{+}21, See20, VXHK21]. Modern [MFM21]. Modular [PCL^{+}21, SDF^{+}21, SLT^{+}21]. Molecular [PCL^{+}21]. Momentary [BahRKM18]. Moneywork [PKC^{+}21]. Monitoring [WGM21]. MOOCs [KCC20]. Moods [MMH21]. Most [KPW19]. Motherhood [BBS19]. Mothers [BBS19]. Motion [GB18]. Motivating [AS17, BRM21, CGS^{+}21]. Motivations [GPH19, SGF20]. Motivators [PM21]. Motives [BZC^{+}17]. Mountain [BBW21]. Mourning [BHM19]. Movement [IOES20, JCD19, UYM21]. Movements [PM21]. Moving [FD20]. MTBalance [BBW21]. Multi [ABQ17, BLC^{+}21, GVNK20, HOZ^{+}21, LAHW19, LDP18, MDM^{+}19, OSM^{+}20, RLGG21, SCP^{+}21, SBPP20, SAE^{+}21, TD19, TVS17]. Multi-Application [TD19]. Multi-Channel [ABQ17, RLGG21]. Multi-Device [LAHW19]. Multi-layered [TVS17]. Multi-modal [BLC^{+}21]. Multi-Resident [SAE^{+}21]. Multi-Staged [LDP18]. Multi-Stakeholder [SCP^{+}21]. Multi-stroke [GVNK20, OSM^{+}20]. Multi-Surface [MDM^{+}19]. Multi-tasking [SBPP20]. Multi-Touch [HOZ^{+}21]. Multilingual [BKJ^{+}20, CYW18]. Multimodal [BCCV17, CCNY19, WO20, YOFC19]. Multiparty [CYW18]. Multiple [BLH^{+}17, LDP18, MON^{+}21]. Multitasking [WGN^{+}21]. Multitouch [RRF20]. Multiview [SEC^{+}20]. Music [CILM18, JCD19]. My [BVM21, CYC^{+}21, CLL^{+}21, GSC21, IRA^{+}19]. Mysteries [JA21].

n [LZY^{+}21]. Nairobi [AMJ^{+}19]. Named [CH20]. Narrative [BTM^{+}21, NM21, WB21]. Narratives [NM21]. Nature [MT21, MYS^{+}21, RVBC^{+}21]. Navigate [KG21, LKG21]. Navigation [AMR21, LZY^{+}21]. Near [CSMV21]. Near-Keyboard [CSMV21]. Need [CEM17]. Needs [BDW21]. Negatively [DRK18]. Negotiating [CS18b]. Neighborhood [HSVR21]. Net [TD19]. Network [ALG^{+}21, HSVR21, SS21, SPT^{+}20]. Networked [BHM19]. Networking [Ann20]. Networks [BB18, CHC^{+}17, MDM^{+}19, WYY^{+}21]. never [IRA^{+}19]. Newcomer [JVP^{+}19]. Newcomers [NHH20]. News [BZSM20, LHPH20, MAMP21, RAZ^{+}20]. Next [RS19]. NEXTGEN [SLD21]. Nicer [WB21]. Nobody [SS21]. Non [BSS18, FLC17, GPH19, RVBC^{+}21, SSGB20, SSGB21, UAAR20, ZSOJ20, Gar21]. Non- [Gar21]. Non-/ [Gar21]. Non-use [GPH19, SSGB21]. Non-Verbal [FLC17]. Non-Violent [RVBC^{+}21]. Non-Visual [UAAR20, ZSOJ20]. Non/use [BSS18, SSGB20]. Noncontact [MMM^{+}21]. Nondominant [YGHR21]. Norm [CSJ^{+}18, FB19, NSJ21]. Norms [RKNES20]. Notation [HG19]. Notations [MPB^{+}19]. Note [KFMH17]. Note/Chairs [KFMH17]. Notebooks [WMBO19]. Notifications [COKL20]. November [CM20, FJIH21]. Novice [BBW21, DDFK19, FGG17, WGM21]. Nudge [BKJ^{+}20]. Number [RA21a].

Object [MMM^{+}21, ZMZL21]. Objects [CS18b, WGG20, XHH17]. Obligations [BR17]. Occupations [MGG^{+}21]. Oct [KLS20]. October [BLS21]. Odin [HOSK21]. Off [TCH20, BKG^{+}21]. Office [LAR21]. Older [KSX21]. Omni [MNP^{+}20]. Omni-Channel [MNP^{+}20]. On-Hand [ZSOJ20]. On-Line [VZV19]. On-screen-start [YU20]. One [FC20, UYM21]. One-dimensional [UYM21]. One-Handed [FC20]. Online [ADAY21, AOT^{+}18, BTM^{+}21, BDSL17, BKKD21, BBNT21, BPO21, CILM18, CS18a, CCX^{+}20, CCWH21, CWB^{+}18, DHH18, DS20, FcLDM19, FD20, FKDG21, JØC^{+}20, JAB19, LX20, OM20, OR19, PM21, SDF^{+}21, SSH^{+}21, See20, WWT^{+}21, XZL^{+}20]. Open [BP17a, LPFD21, WWSG21]. Open-Ended [BP17a]. OpenUIDL [MNP^{+}20]. Operating [HBJR21]. Operations [ASS18, LZY^{+}21]. Operators [SDO21]. Opportunities [Ann20, BWDP20, KOOL19, MWH19, SHY^{+}18, SWF^{+}20, BL17]. Optimization [Sal17]. Optional [BRM21]. Order [EHH^{+}18]. Organizational [RYCC21]. Organizing [BSS18]. Oriented [MCC^{+}18]. Orienting [BHM19, KG21]. OT [SSNC20]. Outages [ALG^{+}21]. Outcomes [ALD^{+}20, ERB^{+}17]. Outdoor [HFBL19, LAW19]. Outer [KG21]. Output [SCMS18]. outreach [Gil20]. Overcome [ETE^{+}19]. Owned [LX20].

PACM [CM20]. PACMHCI [BKL^{+}21, BLS21, FJIH21, GMM21, KLS20, KLS21]. Page [NB20]. Pages [Sal17]. Painter [GGF^{+}20]. Pandemic [HSVR21, RIK21, YGHR21]. Paper [Hig20, HGG20]. Paper-cutting [Hig20]. Papers [CCH^{+}18]. Parallax [GB18]. Parallel [RRF20, SWF^{+}20]. Parent [ADAY21, MFM21]. Parent-Child [MFM21]. Parental [YGHR21]. Part [ASS18, MH19]. Part-Time [MH19]. Partial [CS18a]. Participation [AS17, CYW18, Hao21, KWdCL^{+}21, TAS^{+}19]. Participatory [BKM20, FKK19]. Partner [BTM^{+}21]. Partnerships [KTBL^{+}20]. Party [GN20, PKC^{+}21]. Passing [VKI21]. Passive [ZBL19]. Past [BWV20]. Paste [LZY^{+}21, LZY^{+}21]. Patent [SLD21]. Pathologists [GHHC21]. Patient [FSC21, HKM^{+}20]. Patients [BLH^{+}17, SWF^{+}20]. Patterns [AV17]. Payment [CEM17]. PDF [UAAR20]. Pedagogy [CLKEF21, MD20]. Peer [BZC^{+}17, CGS^{+}21, CKS18, DC17, DRK18, HKM^{+}20, Kee19, LM20]. Peer-Produced [Kee19]. Peer-to-Peer [CGS^{+}21]. People [AM21, CR20, KRK^{+}21, MRGN20, PSM21, SKV^{+}21, Tan21, WB21, BRN^{+}19]. Perceived [DRAQ18, GPH21, HRC^{+}21, KML^{+}21, KYB21, ZBL19]. Perception [ADAY21, ALD^{+}20, HWYW21]. Perceptions [CH20, vBGH^{+}19]. Perform [BRM21]. Performance [BRM21, FC20, LUFW20]. Peripheral [SVW19]. Perpetual [WMM^{+}19]. Personal [Ann20, BLL^{+}21, DWB18, EJB^{+}20, HFL19, LLDH21]. Personalised [HLV21]. Personalized [FCR^{+}17, LZS^{+}21]. Perspective [VKI21]. Perspectives [BHM19, RYCC21]. Persuasive [DDFK19]. Pervasive [NMBS21]. Petri [TD19]. Phone [AHCD17, IRA^{+}19]. Phones [YRIV20]. Phonocardiograms [CGM^{+}18]. Photography [WGG20]. Physical [GRG20, MLSH21, XHH17]. Physicalizations [HOZ^{+}21]. Pick [PKC^{+}21]. Pickers [RA21a]. Place [FKK19, SAE^{+}21]. Plan [ABC21, KYB21]. Planning [ABC21]. Platform [FD20, LHI20, LX20, MH19, MV19, RJZ^{+}21]. Platforms [BBNT21, HJM20, NKV21]. Play [AMJ^{+}19, GHHS21, LKG21, MFM21, APW^{+}21, GMM21]. Players [BRM21, MHG^{+}21, RGR^{+}21]. Playful [AKIS21]. Playing [AL21, BJ21]. Please [KWdCL^{+}21]. Pointing [DWMV21, YU20, ZMZL21]. Points [KIK20]. polarisation [GN20]. Policy [CCS^{+}18]. Policymaking [AS17]. Politeness [LHT^{+}20]. Political [Kee19, SCP^{+}21, GN20]. Politicians [VMP21]. Politics [GN20, SHD21, SDF^{+}21]. Polymodal [BCCV17]. Polyphony [RH19]. Pool [MMH21]. Porous [CS18b]. Portfolios [FKDG21]. Post [CCS^{+}18, GSC21]. Post-College [GSC21]. Posting [JZB^{+}21]. Posture [IOES20, WDP^{+}20]. Postures [CSMV21]. Potential [CLKEF21, GH21]. Power [LNM21, MSY20]. powerful [BRN^{+}19]. Practices [AL18, BL17, CFD^{+}21, FSC21, Gar21, HGG20, JA21, LUFW20, PBC20, RYD21, RYCC21, SAR21, WMM^{+}19]. Practitioner [RYCC21]. Predict [GHHS21]. Predicting [RGR^{+}21, YU20]. Prediction [MMH21]. Predictive [ASP20]. Predictors [SSGB21, vBGH^{+}19]. Prefer [BKJ^{+}20]. Preferences [dACZ19]. Presence [BPO21]. Present [BWV20, CWB^{+}18]. Presentation [LUFW20, RS19]. Presenting [DWB18]. Preserving [KSX21]. Presidential [Kee19]. Press [LZY^{+}21]. Press-n-Paste [LZY^{+}21]. Pressure [BKM21, LZY^{+}21]. Pressure-sensitive [LZY^{+}21]. Prevent [MBT^{+}20]. Prevention [RAZ^{+}20]. Primary [BBV19]. Principal [HLV21]. Principal-Agent [HLV21]. Principles [EJB^{+}20, GC19]. Printed [KIK20]. Prison [ETE^{+}19, TAS^{+}19]. Privacy [AFKL20, AHCD17, CHC^{+}17, KSX21, LLDH21, RKNES20, SB20b, BTM^{+}21]. Privacy-Preserving [KSX21]. PrivacyPrimer [KSX21]. Private [COKL20, DC17]. Proactive [CCS^{+}18]. Probability [GSK^{+}17]. Probe [HKM^{+}20].

Problem [BP17a, FCR^{+}17, MTA^{+}17, SCP^{+}21]. Problems [BBV19]. Procedures [MAN^{+}20]. Process [CMN^{+}21, CMP^{+}21, CNH^{+}17]. Produce [KDEA21]. Produced [Kee19]. Production [BZC^{+}17, CSMG17, Hig20]. Productivity [LAR21]. Professional [LYH20]. Programming [BahRKM18, BPO21, LZS^{+}21, MCC^{+}18, RH19, Sal17, TDC17]. Project [BL17, Gil20, AMJ^{+}19, SJ21]. Projection [GB18]. Projections [KFSC18]. Projects [LPFD21, WWSG21]. Promote [BGRK17, ETE^{+}19, LOSD20]. Promoting [dACZ19, LYH20]. Prompts [DAQ17]. Properties [CPB18]. Proprioceptive [BBW21]. Props [MLSH21]. Prospects [CCWH21]. Protection [BUSA^{+}21]. ProtoChat [CMP^{+}21]. Protocol [KE18]. Protocols [BahRKM18]. Prototyping [HDPL21, KKE^{+}21, KVDG20, SLN21]. Provider [KYB21]. Providers [BLH^{+}17, FSC21]. Providing [BPO21]. Proxy [DHH18]. Psychophysiological [BDL21]. Psychosocial [ETE^{+}19]. Public [AL18, AKIS21, Fie19, Gil20, MD20, STM^{+}19, TAS^{+}19]. Publications [KKS19]. Pull [LO17]. Puppetry [LO17]. Puts [SS21]. PWA [dACZ19]. PWA-EU [dACZ19].

Q&A [CXL17, CCS^{+}18, CT18]. Q&As [DMC^{+}20]. Qualitative [BWDP20, LE21]. Quality [ADD21, CCS^{+}18, WYY^{+}21]. Quantifying [TSS21]. Querying [HOZ^{+}21]. Question [MCYZ19, MWM^{+}19, MMH21]. Questions [DMC^{+}20]. Quilters [MWH19]. Quotidian [AL18].

r [RJZ^{+}21, Gil20]. r/AskHistorians [Gil20]. r/Incels [RJZ^{+}21]. r/The Donald [RJZ^{+}21]. Racial [MGG^{+}21]. Racism [TSHK20]. Rapid [KVDG20]. Rapidly [MTA^{+}17]. Raters [BZSM20]. Rates [UYM21]. Re [AWM21]. Re-Working [AWM21]. Reachability [YRIV20]. React [ZRL21]. Reading [JCD19]. Real [BBW21, DDZ18, FLC17, GHHS21, LYH20, MRRS19, PCMP18, SBPP20, SSNC20, WMBO19]. Real-Time [BBW21, DDZ18, MRRS19, PCMP18, SBPP20, WMBO19, FLC17]. Reality [ALG^{+}21, BLC^{+}21, BRM21, FSN^{+}20, HOSK21, KMM21, LZY^{+}21, MAN^{+}20, MLSH21, OG21, Pri19, RYCC21, SCMS18, SCY^{+}18, SHY^{+}18, SLN21, TRO^{+}21, TG18, WO20]. Receipts [CEM17]. Recidivism [vBGH^{+}19]. Recognition [BKG^{+}21, CS18a, DS20, DWZ^{+}20, GRG20, RA21b]. Recognition-Enabled [RA21b]. Recognizing [MH19, OSM^{+}20]. Recommendation [HvBKG20]. Recommendations [BBNT21, TD19]. Recommending [CRN21]. Reconsidering [HFBL19, See20]. Recording [AV19]. Recruitment [AOT^{+}18, KOOL19]. Reddit [CPS^{+}17, CSJ^{+}18, GKS21, Gil20, LLDH21, XZL^{+}20]. Redditor [SS21]. Reduce [JZB^{+}21, RVBC^{+}21]. Referencing [CMY^{+}17, CWB^{+}18]. Referential [MHN^{+}21]. Reflect [GBN21]. Reflection [BLL^{+}21]. Reflections [ERT21]. Reflects [JAB19]. Rehabilitation [HRC^{+}21]. Relating [SBAK21]. Relational [MT21, WGG20]. Relationship [SK20]. Relationships [ABQ17, HLV21, MFM21]. Relative [DWMV21]. Reliable [KMR20]. Remote [ALG^{+}21, MMM^{+}21, PBC20, RIK21]. Repair [SSH^{+}21]. Repairing [CLL^{+}21]. Replay [MON^{+}21]. Report [AL18]. Reporting [APW^{+}21, JKW^{+}21]. Reports [RAZ^{+}20]. Represent [NM21]. Representation [MON^{+}21, MGG^{+}21]. Representations [BLC^{+}21]. Reproducibility [FSN^{+}20]. Reproducing [BV21]. Requirements [SWT20]. Research [BUSA^{+}21, BL17, CLKEF21, CCH^{+}18, FSN^{+}20, Fie19, NHH20, See20]. Researchers [ZD19]. Resident [SAE^{+}21]. Resistance [KDEA21]. Resocialization [ETE^{+}19]. resolution [SJ21]. Resource [RIK21]. Resources [CH20]. Response [BL18, HKF19]. Responsible [MNS^{+}21, RYCC21]. Responsive [LZS^{+}21, XHH17]. Results [MGG^{+}21, VH21]. Retention [BGRK17]. Rethinking [BP17b, Hao21, YU20]. Returning [BSS18]. Retweet [VMP21]. Reveals [GSC21]. Reversion [GPH19]. Review [APW^{+}21, DWW^{+}17, DFJ^{+}21, LOSD20]. Reviews [DAQ17]. Revise [KYB21]. Revisiting [SB17]. RFID [DWZ^{+}20]. Ride [MH19]. Ride-Share [MH19]. Ride-Sharing [MH19]. Right [MWM^{+}19]. Rights [GYD17]. Rise [LX20]. Risky [BUSA^{+}21]. Role [AL21, CCX^{+}20, CR20, FSN^{+}20, IRA^{+}19, KYB21, OR19, PSM21, See20, VKI21]. Roles [LDP18, PBC20, PM21, ZMW20]. Romantic [ABQ17]. Routines [ZBL19]. Routing [MCYZ19]. Rules [CSJ^{+}18]. run [Gil20]. Runtime [MNP^{+}20]. Rural [SAR21, WGG20]. Rwandan [HMB^{+}21].

Safe [RAZ^{+}20]. Safety [AL18]. Sample [AOT^{+}18, DS20]. Sampling [CLP^{+}21]. Sanctioning [RKNES20]. SAPIENS [SVW19]. Sarcasm [OM20]. Satisfaction [YHR^{+}19]. Saudi [ADAY21, RYD21]. Savings [MAR^{+}19]. say [CNH^{+}17]. Scalable [ZBL19]. Scale [LKS21, POG^{+}21]. Scales [CSJ^{+}18]. Scandinavia [JAB19]. Scene [MON^{+}21]. Schizophrenia [ERB^{+}17]. Scholarship [Gil20]. Science [JØC^{+}20, MHG^{+}21, ZMW20]. Sciences [LKS21]. Scientific [BP17a, CRN21, FSN^{+}20, HRRS20, KKS19, MWM^{+}19, MRRS19, NHH20]. Scientists [KG21, MWM^{+}19, WMBO19]. Score [DRK18, JCD19]. Score-Group [DRK18].

Screen [BPO21, GB18, YU20]. Screens [BCCV17, DWMV21]. Scripting [HRRS20]. Sculptures [Hao21]. Search [ERLW17, HJM20, MGG^{+}21, VH21]. Seat [RI20]. See [BKG^{+}21]. Seeing [CSS^{+}17, SJ21, SAR21]. Seek [HWYW21]. Seeking [KBFK21, KYB21]. Selection [BDL21]. Selective [DRA^{+}17]. Self [BR20, FCR^{+}17, FCE^{+}18, KRMM21, LYH20, LUFW20, See20]. Self-Attach [BR20]. Self-Disclosure [LYH20]. Self-Moderation [See20]. Self-Presentation [LUFW20]. Self-Similar [KRMM21]. Self-Tracking [FCR^{+}17, FCE^{+}18]. SEME [ERLW17]. Sender [COKL20]. Sender-Controlled [COKL20]. Sending [AMF18]. Sense [BSS18]. Sensemaking [FGG17, MNS^{+}21]. Sensing [FLC17, VMD^{+}20, ZBL19]. Sensitive [HKF19, LZY^{+}21]. Sensitivity [BS21]. Sensor [AFKL20, MNS^{+}21]. September [GMM21]. Sequences [TBR20]. Serious [AL21, CGS^{+}21, GH21, KML^{+}21]. Services [BBV19, GN20, HLV21]. Session [SB17]. Sets [FLC20]. Setting [ALD^{+}20, KYB21]. Shape [PCL^{+}21]. Shape-changing [PCL^{+}21]. Share [COKL20, MH19]. Shared [AHCD17, FBG17, RCM21]. Sharing [BBNT21, BPO21, COKL20, DWW^{+}17, EJB^{+}20, HLV21, JZB^{+}21, MH19, NKV21, PBC20, VKI21]. Shift [RIK21]. Shifting [JØC^{+}20, RYCC21]. shops [IRA^{+}19]. Short [Pri19]. Short-term [Pri19]. Should [FLC17]. Side [FLC20, YRIV20, CLP^{+}21]. Side-Crossing [FLC20]. Signals [MFM^{+}20, ZWZ^{+}21]. Signature [ZWZ^{+}21]. Similar [KRMM21]. Similarities [HGG20]. Simplified [LAHW19]. simply [Pri19]. Simulate [HOSK21]. Simulated [AM21]. Site [DMC^{+}20, Gil20]. Sites [CXL17, CCS^{+}18, MD20]. Situ [KFSC18]. Situated [BHM19, KG20, MT21]. Situational [ZAS^{+}21]. Situations [HLV21].

Sizes [UYM21]. SketchADoodle [GVNK20]. Sketched [SB20a]. Skill [CKS18, HvBKG20, Hig20, MON^{+}21, MHG^{+}21]. Slidable [SVV21]. Sliders [GSK^{+}17]. Slime [SVV21]. Slogans [AOT^{+}18]. SMAC [LAHW19]. Smala [MCC^{+}18]. Small [BCCV17, FLC20, KBFK21]. Smart [Gar21, SVW19, TDC17, ZMZL21]. Smartest [GSC21]. Smartglasses [AV19]. Smartglasses-based [AV19]. Smartphone [IOES20, ZWZ^{+}21]. Smartphone-based [ZWZ^{+}21]. Smartwatches [RA21a]. So-called [BTM^{+}21]. Sochiatrist [MFM^{+}20]. Social [AMR21, AMF18, Ann20, AWM21, ALD^{+}20, AM21, BSS18, BWDP20, BBNT21, BPO21, CEM17, CSMG17, CH20, CCX^{+}20, CHC^{+}17, DWB18, DBS^{+}21, DRAQ18, EJB^{+}20, ERB^{+}17, ETE^{+}19, FB19, GYD17, GAW19, GSC21, GPH19, GPH21, GN20, HLV21, Hao21, HFL19, JZB^{+}21, KBFK21, LKS21, LNM21, MCYZ19, MAMP21, PSM21, RKNES20, RA21b, SKV^{+}21, SSH^{+}21, SPT^{+}20, TSHK20, TCH20, VKI21, WWT^{+}21, WYY^{+}21, ZD19]. Social-Support [WWT^{+}21]. Sociality [BRN^{+}19]. Socially [AMF18]. Societal [JAB19]. Society [MGG^{+}21]. Socio [KBKH21, SCP^{+}21]. Socio-Ecological [KBKH21]. Socio-Political [SCP^{+}21]. Sociocultural [OM20]. Sociotechnical [RYD21]. Software [CMN^{+}21, CLKEF21, LPFD21, MMdSP21, SVW19, TG18, WWSG21]. Solicitors [PM21]. Solidarity [NKV21]. Solutions [RS19]. SOLVENT [CCH^{+}18]. Solving [BP17a, MTA^{+}17]. Someone [PKC^{+}21]. Soul [BVM21]. Sound [CSS^{+}17]. Source [LPFD21, WWSG21]. Space [BAAP20, KDH^{+}21, LKG21, NSJ21, PCL^{+}21]. Spaces [AKIS21, LBM21, SEC^{+}20]. Spacetime [DDZ18]. Spark [SKV^{+}21]. Spatial [POG^{+}21, TRO^{+}21, TCH20]. Speakers [Gar21]. Specifying [BF17]. Spectator [WACdA^{+}21]. Spectators [WACdA^{+}21]. Speculative [Fie19]. Speech [CPS^{+}17, MIS^{+}20, WO20, ZMR21]. Speed [ASP20, RLGG21]. Speedrunning [SS19]. Spillover [NSJ21]. Spinal [BHNA18]. Spoil [CT18]. Spontaneous [LUFW20]. Spread [Wil20]. SQUAMATA [CMN^{+}21]. Stage [SLN21]. Staged [LDP18, CS18b]. Stairlifts [SAE^{+}21]. Stakeholder [SCP^{+}21]. start [YU20]. State [CLV^{+}19, LX20]. State-Owned [LX20]. Statement [CNC^{+}17]. States [CRN21]. Status [BL17, COKL20]. Stay [CPS^{+}17, GPH19]. Step [AMR21]. Stepper [RA21a]. Stereotype [BKG^{+}21]. Sticker [CKP^{+}18]. Sticky [BGRK17]. Stigmatized [AMF18]. Stories [JA21]. Storyboards [HWS21]. stranger [BLH^{+}17]. Strangers [SK20]. Strategic [BWDP20]. Strategies [HFL19, RKNES20]. Streamed [SS19]. streamer [CCWH21]. Streamers [JSP^{+}21]. Streaming [CCWH21, LHPH20]. Streams [CEM17]. Strenuous [BRM21]. Stress [GSC21, RVBC^{+}21, ZBL19]. StressMon [ZBL19]. Strip [HGG20]. Strips [HGG20]. Stroke [CFD^{+}21, CS18a, HRC^{+}21, GVNK20, OSM^{+}20]. Stronger [HSVR21]. Structural [MDM^{+}19]. Structuring [PCL^{+}21]. Struggles [FCE^{+}18]. Stuck [DFJ^{+}21, LRSS20]. Student [MMH21]. Students [LKS21]. Studies [MV19]. Studious [LUFW20]. Study [ABC21, ALG^{+}21, ADD21, AKIS21, BKM20, BKKD21, BS21, BHNA18, CSJ^{+}18, CCX^{+}20, CNH^{+}17, CHC^{+}17, FB19, Gil20, HKM^{+}20, HJM20, KFSC18, LLDH21, LAW19, MAN^{+}20, MIS^{+}20, NHH20, WGGS20, XZL^{+}20, vBGH^{+}19]. Studying [BB20, LHT^{+}20, SB20b]. Style [GAW19]. Styles [QGB20]. Subject [BL17, RA21b]. Subjective [CLP^{+}21, GBN21]. Subjectivity [MSY20]. Subreddit [SS21]. Subtletee [WDP^{+}20]. Success [GAW19, UYM21]. Successful [OR19]. Suicide [RAZ^{+}20]. Support [BMPP20, CPB18, FCR^{+}17, GKS21, GB18, HKM^{+}20, KSX21, KTBL^{+}20, LW19, LM20, LDP18, MMM^{+}21, MRRS19, NHH20, PCMP18, SVW19, SB17, WWT^{+}21]. Supported [MPB^{+}19, TG18]. Supporting [BLL^{+}21, BWDP20, BBS19, BHNA18, CFD^{+}21, CR20, CMP^{+}21, HRRS20, OTH20, See20, TDC17, XHH17]. Suppressing [ERLW17]. Surface [GRG20, LZY^{+}21, MDM^{+}19, GVNK20]. Surfaces [CYGW21, FBG17, FLC20, KE18, XHH17]. Surgery [MON^{+}21]. Surgical [MON^{+}21]. Surround [GB18]. Surround-Screen [GB18]. Surveillance [BTM^{+}21]. Survey [AM21, BDW21, SSGB21]. Survey-Based [SSGB21]. Sustainable [NPT19]. Swipe [RA21a]. Switching [CYW18, DWMV21]. Symbiosis [CR20]. Symbolic [PBC20]. Sympathizers [PM21]. Synchronized [MON^{+}21]. Synchronized-Video [MON^{+}21]. Synchronous [NSJ21, Pri19]. System [AV19, BDW21, BP17b, BGRK17, CRN21, CCH^{+}18, CYW18, HKM^{+}20, HvBKG20, LNM21, MCYZ19, MON^{+}21, dQPMdHRdAN21, RH19, SBPP20, SLD21, VN19, ZRL21, ZD19]. System-Facilitated [CYW18]. Systematic [APW^{+}21, DWW^{+}17, DAQ17, DFJ^{+}21]. Systems [BWV20, CPB18, CSN19, FLP^{+}21, HGG20, JVP^{+}19, LDP18, MRRS19, NPT19, TG18, TBR20].

Table [MMM^{+}21, RI20]. Tabletop [MTA^{+}17]. Tabletops [MRGN20]. Tabloids [CSMG17]. Tactile [CYGW21, GRG20, RLGG21]. Tag [AV19, CT18]. Tagging [BKG^{+}21]. Tags [AV19, AWM21]. Tailored [LDP18]. Take [ZRL21]. Talk [LLDH21]. Tangible [AFKL20, KE18, MDM^{+}19, MTA^{+}17]. Taps [YRIV20]. Targets [UYM21]. Task [BDW21, BF17, CFG^{+}17, DBS^{+}21, HBJR21, HvBKG20, MPB^{+}19, MHN^{+}21, PCMP18, SWT20, SGF20, WYY^{+}21]. tasking [SBPP20]. Tasks [BZSM20, WO20, YU20, YOFC19]. Taxing [MYS^{+}21]. Taxonomy [AM21]. Teachers [GLZ^{+}21]. Team [CYC^{+}21]. Teams [ADD21, BKM20, BS21, CYC^{+}21, KJH19]. Technical [VMD^{+}20]. Techniques [LHI20]. Technological [IRA^{+}19, KJH19]. Technologies [ADAY21, Ann20, HFBL19, TSHK20]. Technology [ADD21, BDL21, BS21, CFD^{+}21, HKM^{+}20, KOOL19, SWF^{+}20, YGHR21]. Technology-based [YGHR21]. Technology-Enabled [SWF^{+}20]. Tedious [MYS^{+}21]. Teens [BUSA^{+}21]. Telehealth [FLC17]. Telerehabilitation [CFD^{+}21]. Telework [Tan21]. TEMPEST [BahRKM18]. Temporal [MIS^{+}20]. Tennis [MMM^{+}21]. Tensions [KPW19, LZAD21]. Term [CWB^{+}18, HFBL19, Pri19]. Test [CFG^{+}17]. Tester [VZV19]. Testing [AMF18, AV17, CMN^{+}21, MP19, TBR20]. Tests [BR17]. Text [ASP20, BB20]. Texture [GRG20]. The Donald [RJZ^{+}21]. Their [HWYW21, CS18b]. Theories [KDEA21]. Therapeutic [ERB^{+}17]. Therapy [CGS^{+}21]. Things [LAR21]. think [BRN^{+}19]. Thinking [MCI20]. Thinking-Aloud [MCI20]. Third [PKC^{+}21]. thons [MD20]. Thought [RI20]. Threaded [BGRK17]. Threading [BGRK17]. Throughout [DWB18]. Thumb [YRIV20]. Ties [DRAQ18]. TikTok [KDEA21]. Tilt [RA21a]. Time [BBW21, DDZ18, HKF19, JZB^{+}21, KDH^{+}21, MH19, MRRS19, MFM21, PCMP18, SBPP20, WMBO19, ZRL21, BLH^{+}17, FLC17]. Time-Sensitive [HKF19]. Times [UYM21]. Today [FcLDM19]. Together [HSVR21, MWM^{+}19, SAE^{+}21]. Tomorrow [FcLDM19]. Too [CT18, DWB18]. Tool [CMN^{+}21, GHHC21, KTBL^{+}20, LE21, MPB^{+}19, MP19, NB20, WGGS20]. Tool-Supported [MPB^{+}19]. Toolbox [CPB18]. Tooling [WMM^{+}19]. Toolkit [FBG17, KKE^{+}21, TVS17]. Tools [WGM21, ZMW20]. Topic [ABQ17, KWdCL^{+}21]. Topic-Based [ABQ17]. Touch [CYGW21, DWMV21, GVNK20, HOZ^{+}21, KIK20, TRO^{+}21, YU20, YRIV20]. Touch-surface [GVNK20]. Touchable [UYM21]. Town [JKW^{+}21]. Toxic [CSA^{+}21]. Toxicity [CSA^{+}21, XZL^{+}20]. Toys [Gar21]. Traces [KDH^{+}21]. Tracing [CYGW21]. Tracking [FCR^{+}17, FCE^{+}18, KCC20, MNS^{+}21]. Trade [BKG^{+}21]. Trade-off [BKG^{+}21]. Traditional [SAR21]. Traffic [HGG20]. Trajectories [BKKD21, OSM^{+}20]. Transaction [LRSS20]. Transition [BBS19, GSC21]. Transitioning [DRA^{+}17]. Transitions [BSS18]. Transparency [APW^{+}21]. Treatment [BUSA^{+}21, TBDB21]. Trends [BDW21]. Triage [DAQ17]. Trust [BDL21, KPW19, SSH^{+}21]. Trustworthiness [CH20, GPH21]. Trustworthy [KPW19]. Truth [KMR20]. TUIO [KE18]. TUIOFX [FBG17]. Turn [TCH20]. Tutorial [KMM21]. Tutorial-Based [KMM21]. Tutoring [ZMR21]. Twitch [SK20]. Twitter [BL18, CSMG17, LHT^{+}20]. Two [YRIV20]. Two-Dimensional [YRIV20]. Typealike [CSMV21]. Typologies [SSGB20].

U.S. [Kee19, MM19]. UI [BR17, PKB^{+}19]. UiLab [BV21]. Ultrasound [MMM^{+}21]. Ultrasound-driven [MMM^{+}21]. UMI [CPB18]. Uncertain [GSK^{+}17, UYM21]. Unconstrained [WO20]. Understand [GKS21]. Understanding [Ann20, BSS18, CKP^{+}18, CSMG17, CCX^{+}20, CCWH21, CLP^{+}21, CWB^{+}18, EHH^{+}18, FKDG21, Gar21, KOOL19, LW19, RCM21, SGF20, Tan21, WO20]. Unemployment [GKS21]. Unified [TVS17]. Unifying [RDS18]. Unique [KBFK21]. Unity [CPB18]. Unknown [FSC21]. Untold [JA21]. Urban [SCP^{+}21]. Usability [MCI20]. Usage [LW19]. Use [AHCD17, BB20, BWDP20, GYD17, Gar21, GBN21, HFBL19, MH19, MHN^{+}21, MNS^{+}21, RVBC^{+}21, SSGB21, TAS^{+}19, WMBO19, BSS18, GPH19, SSGB20]. Used [AOT^{+}18, FC20]. User [AFKL20, ALG^{+}21, BR20, BahRKM18, BGRK17, dACZ19, CPB18, CLwH^{+}20, KCC20, KVDG20, KJH19, LLDH21, MAN^{+}20, MNP^{+}20, MLSH21, dQPMdHRdAN21, PCL^{+}21, RLGG21, Sal17, SWT20, SVV21, VZV19, VN19, WGGS20, YHR^{+}19, YMG^{+}20, ZMR21]. User-Centric [AFKL20]. User-Configurable [ZMR21]. User-Defined [MLSH21]. Users [BKJ^{+}20, KDEA21, WGM21, dQPMdHRdAN21]. Uses [LX20]. Using [BahRKM18, BRM21, CCWH21, FSC21, GKS21, GHHS21, JCD19, LOSD20, LKG21, MBT^{+}20, NMBS21, RGR^{+}21, STM^{+}19, WO20, ZBL19, KKS19, LHT^{+}20, MAN^{+}20, Sal17, SCY^{+}18, TRO^{+}21, TBR20, YRIV20, ZWZ^{+}21]. Utility [GPH21]. UX [GGF^{+}20]. UX-Painter [GGF^{+}20].

V4 [CM20, KLS20, KLS21]. V5 [BKL^{+}21, BLS21, FJIH21, GMM21]. Validation [BMPP20]. Value [KML^{+}21]. Values [BLH^{+}17, BLL^{+}21, CS18b, GN20, NPT19, SHD21]. Variables [OM20]. Venmo [CEM17]. Verbal [FLC17]. Verification [ZWZ^{+}21]. Versus [MYS^{+}21]. Vertical [RRF20]. via [MMM^{+}21]. Viability [CYC^{+}21]. VibHand [ZSOJ20]. Vibrotactile [ZSOJ20]. Vicariously [LAW19]. Video [BKM21, DHH18, FLC17, HDPL21, HJM20, HWYW21, MON^{+}21, SCY^{+}18].

Video-Based [DHH18]. Videogames [BVM21]. Videos [DDFK19]. VidLyz [DDFK19]. View [WACdA^{+}21]. Violations [CSJ^{+}18, GYD17, NSJ21]. Violent [RVBC^{+}21]. Virtual [BLC^{+}21, BRM21, GB18, HSVR21, HOSK21, HWYW21, KMM21, MAN^{+}20, MLSH21, OG21, SCMS18, SK20, SLN21, XHH17, YOFC19]. Vision [LAR21, MSY20, SHD21]. Visual [BV21, CMY^{+}17, JSP^{+}21, SB17, SAR21, TRO^{+}21, UAAR20, ZSOJ20]. Visualization [LDP18, NB20]. Visualizations [CSS^{+}17, LHI20]. Visualizing [HWS21]. Visually [HRRS20]. VizSciFlow [HRRS20]. Voice [CLL^{+}21, KRMM21, PKB^{+}19]. Volitionality [SSGB21]. Volunteer [JØC^{+}20]. VR [CCNY19, KTF^{+}21]. vs [RVBC^{+}21].

Walkthrough [WGGS20]. Warning [SLD21]. Was [FKK19, Rib19, BZC^{+}17, CCWH21, Pri19]. watched [Pri19]. Waters [AMF18]. Wave [BMdS^{+}21]. Wearers [WACdA^{+}21]. Weather [BJ21]. Web [AV17, BahRKM18, BMPP20, LZS^{+}21, NB20, OTH20, dQPMdHRdAN21, SCMS18]. Web-Based [BahRKM18]. Web5VR [SCMS18]. website [Gil20]. Websites [MMH21]. Weibo [LHT^{+}20]. Welcome [EVC^{+}19, KFMH17, SN18]. Welfare [HFL19]. Wellbeing [RA21b, Vya19]. WevQuery [AV17]. Whale [RAZ^{+}20]. Where [RYCC21, Pri19]. which [MH19]. Who [HRC^{+}21, PKC^{+}21, SKV^{+}21, SSGB21]. Widgets [CLV^{+}19]. Wikipedia [MD20, VH21]. Wild [BB20]. Will [CYC^{+}21, CT18]. Willed [HRC^{+}21]. Willingness [GPH21]. WIMP [CMN^{+}21]. Wings [SBAK21]. Winning [MMdSP21]. Winter [BJ21]. Within [ASS18]. Without [HBJR21, LAW19, TRO^{+}21]. Women [ERT21, IRA^{+}19, RYD21]. Work [ABC21, BKKD21, BBV19, FSN^{+}20, HGG20, MWM^{+}19, MYS^{+}21, MDZ19, MNS^{+}21, TSS21, Vya19, WMM^{+}19, ZBL19]. Workbench [BV21]. Worker [CKS18]. Workers [HMB^{+}21, HBJR21, ZMW20]. Workflow [CRN21]. Workflows [FKDG21, ZMW20]. Working [AWM21]. Workplace [MNS^{+}21]. Workspace [MRGN20]. World [GHHS21, Gil20]. Worn [HOSK21]. WProfX [NB20]. Wrappers [BBS19]. Write [FM21, LM20].

XAL [GLZ^{+}21]. XD [SHY^{+}18]. XD-AR [SHY^{+}18]. xPress [BP17b].

Young [YGHR21]. you’re [BLH^{+}17]. Yours [SLN21]. Youth [BWDP20, SSH^{+}21]. YouTube [BBNT21, HJM20].

References

Ahmetoglu:2021:PPM

[ABC21] Yoana Ahmetoglu, Duncan P. Brumby, and Anna L. Cox. To plan or not to plan? a mixedmethods diary study examining when, how and why knowledge work planning is inaccurate. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW3):222:1–222:20, January 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                                         https://dl.acm.org/
                                                         doi/10.1145/3432921.

Andalibi:2017:MCT

[ABQ17] Nazanin Andalibi, Frank Bentley, and Katie Quehl.

Multi-channel topic-based mobile messaging in romantic relationships. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 1 (CSCW):20:1–20:18, December 2017. CODEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3134655.

Al-Dawood:2021:SAP

[ADAY21] Adel Al-Dawood, Serene Alhajhussein, and Svetlana Yarosh. Saudi Arabian parents’ perception of online marital matchmaking technologies. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4(CSCW3): 211:1–211:32, January 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3432910.

Almahmoud:2021:HTC

[ADD21] Jumana Almahmoud, Robert DeLine, and Steven M. Drucker. How teams communicate about the quality of ML models: a case study at an international technology company. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (GROUP):222:1–222:24, July 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3463934.

Ahmad:2020:TPT

[AFKL20] Imtiaz Ahmad, Rosta Farzan, Apu Kapadia, and Adam J. Lee. Tangible privacy: Towards user-centric sensor designs for bystander privacy. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW2):116:1–116:28, October 2020. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3415187.

Aromaa:2018:DHM

[AFT^{+}18] Susanna Aromaa, Nikos Frangakis, Domenico Tedone, Juhani Viitaniemi, and Iina Aaltonen. Digital human models in human factors and ergonomics evaluation of gesture interfaces. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 2(EICS):6:1–6:14, June 2018. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3229088.

Asthana:2018:FEA

[AH18] Sumit Asthana and Aaron Halfaker. With few eyes, all hoaxes are deep. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 2(CSCW): 21:1–21:18, November 2018. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3274290.

Ahmed:2017:DPC

[AHCD17] Syed Ishtiaque Ahmed, Md. Romael Haque, Jay Chen, and Nicola Dell. Digital privacy challenges with shared mobile phone use in Bangladesh. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 1 (CSCW):17:1–17:20, December 2017. CODEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3134652.

Araake:2021:PEP

[AKIS21] Koichi Araake, Michinari Kono, Eiji Iwata, and Norio Sasaki. Playful engagement for public spaces: a case study on a mall escalator. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(ISS): 498:1–498:19, November 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3488543.

AlvaradoGarcia:2018:QRG

[AL18] Adriana Alvarado Garcia and Christopher A. Le Dantec. Quotidian report: Grassroots data practices to address public safety. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 2(CSCW): 17:1–17:18, November 2018. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3274286.

Ahmad:2021:EGC

[AL21] Azhan Ahmad and Effie Lai-Chong Law. Educators as gamemasters: Creating serious role playing game with. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CHI PLAY):230:1– 230:29, September 2021. CODEN ???? ISSN 2573-0142 (electronic).

                                        URL https://dl.acm.
                                        org/doi/10.1145/3474657.

Ashktorab:2020:HAC

[ALD^{+}20] Zahra Ashktorab, Q. Vera Liao, Casey Dugan, James Johnson, Qian Pan, Wei Zhang, Sadhana Kumaravel, and Murray Campbell. Human-AI collaboration in a cooperative game setting: Measuring social perception and outcomes. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4(CSCW2):96:1–96:20, October 2020. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3415167.

Ahsen:2021:ENO

[ALG^{+}21] Tooba Ahsen, Zi Yi Lim, Aaron L. Gardony, Holly A. Taylor, Jan P de Ruiter, and Fahad Dogar. The effects of network outages on user experience in augmented reality based remote collaboration — an empirical study. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (CSCW2):313:1–313:27, October 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3476054.

Azad:2021:LCP

[AM21] Sasha Azad and Chris Martens. Little computer people: a survey and taxonomy of simulated models of social interaction. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CHI PLAY):245:1– 245:30, September 2021. CODEN ???? ISSN 2573-0142 (electronic).

                                        URL https://dl.acm.
                                        org/doi/10.1145/3474672.

Ames:2018:HCC

[Ame18] Morgan G. Ames. Hackers, computers, and cooperation: a critical history of logo and constructionist learning. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 2(CSCW): 18:1–18:19, November 2018. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3274287.

Andalibi:2018:TWS

[AMF18] Nazanin Andalibi, Margaret E. Morris, and Andrea Forte. Testing waters, sending clues: Indirect disclosures of socially stigmatized experiences on social media. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 2(CSCW): 19:1–19:23, November 2018. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3274288.

Arawjo:2019:CEI

[AMJ^{+}19] Ian Arawjo, Ariam Mogos, Steven J. Jackson, Tapan Parikh, and Kentaro Toyama. Computing education for intercultural learning: Lessons from the Nairobi Play Project. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3(CSCW): 52:1–52:24, November 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3359154.

Albarrak:2021:DMS

[AMR21] Luluah Albarrak, Oussama Metatla, and Anne Roudaut. (don’t) mind the step: Investigating the effect of digital social cues on navigation decisions. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(ISS): 492:1–492:18, November 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3488537.

Annett:2020:UHO

[Ann20] Michelle Annett. Understanding the homepreneurship opportunities afforded by social networking and personal fabrication technologies. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4(CSCW2):99:1–99:48, October 2020. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3415170.

August:2018:FEC

[AOT^{+}18] Tal August, Nigini Oliveira, Chenhao Tan, Noah Smith, and Katharina Reinecke. Framing effects: Choice of slogans used to advertise online experiments can boost recruitment and lead to sample biases. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 2(CSCW): 22:1–22:19, November 2018. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3274291.

Aeschbach:2021:TMR

[APW^{+}21] Lena Fanya Aeschbach, Sebastian A. C. Perrig, Lorena Weder, Klaus Opwis, and Florian Brühlmann. Transparency in measurement reporting: a systematic literature review of CHI PLAY. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CHI PLAY):233:1– 233:21, September 2021. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3474660.

Aitamurto:2017:MPC

[AS17] Tanja Aitamurto and Jorge Saldivar. Motivating participation in crowdsourced policymaking: The interplay of epistemic and interactive aspects. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 1 (CSCW):18:1–18:22, December 2017. CODEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3134653.

Alharbi:2020:EPF

[ASP20] Ohoud Alharbi, Wolfgang Stuerzlinger, and Felix Putze. The effects of predictive features of mobile keyboards on text entry speed and errors. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4(ISS): 183:1–183:16, November 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3427311.

Arif:2018:APE

[ASS18] Ahmer Arif, Leo Graiden Stewart, and Kate Starbird. Acting the part: Examining information operations within #Black-LivesMatter discourse. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 2 (CSCW):20:1–20:27, November 2018. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3274289.

Apaolaza:2017:WTH

[AV17] Aitor Apaolaza and Markel Vigo. Wev-Query: Testing hypotheses about Web interaction patterns. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 1(EICS):4:1–4:17, June 2017. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3095806.

Aiordachioae:2019:LTS

[AV19] Adrian Aiordachioae and Radu-Daniel Vatavu. Life-tags: a smartglassesbased system for recording and abstracting life with tag clouds. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3 (EICS):15:1–15:22, June 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3331157.

Antoniak:2021:TBC

[AWM21] Maria Antoniak, Melanie Walsh, and David Mimno. Tags, borders, and catalogs: Social re-working of genre on LibraryThing. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CSCW1):29:1–29:29, April 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3449103.

Bhatia:2020:EDS

[BAAP20] Arpit Bhatia, Yajur Ahuja, Suyash Agarwal, and Aman Parnami. Exploring the design space of badge based input. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4(ISS): 185:1–185:20, November 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3427313.

Batalas:2018:UTE

[BahRKM18] Nikolaos Batalas, Marije aan het Rot, Vassilis Javed Khan, and Panos Markopoulos. Using TEMPEST: Enduser programming of Web-based ecological momentary assessment protocols. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 2 (EICS):3:1–3:24, June 2018. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3179428.

Berenberg:2018:ECE

[BB18] Daniel Berenberg and James P. Bagrow. Efficient crowd exploration of large networks: The case of causal attribution. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 2(CSCW): 24:1–24:25, November 2018. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3274293.

Bemmann:2020:LMK

[BB20] Florian Bemmann and Daniel Buschek. LanguageLogger: a mobile keyboard application for studying language use in everyday text communication in the wild. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (EICS):84:1–84:24, June 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3397872.

Buntain:2021:YRE

[BBNT21] Cody Buntain, Richard Bonneau, Jonathan Nagler, and Joshua A. Tucker. YouTube recommendations and effects on sharing across online social platforms. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CSCW1):11:1–11:26, April 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3449085.

Britton:2019:MCW

[BBS19] Lauren Britton, Louise Barkhuus, and Bryan Semaan. “Mothers as Candy Wrappers”: Critical infrastructure supporting the transition into motherhood. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3 (GROUP):232:1–232:21, December 2019. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3361113.

Bopp:2019:CPK

[BBV19] Chris Bopp, Lehn M. Benjamin, and Amy Voida. The coerciveness of the primary key: Infrastructure problems in human services work. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3(CSCW): 51:1–51:26, November 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3359153.

Berentsen:2021:MAN

[BBW21] Mark J. Berentsen, Marit Bentvelzen, and Paweł W. Woźniak. MTBalance: Assisting novice mountain bikers with realtime proprioceptive feedback. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(ISS): 506:1–506:25, November 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3488551.

Bouzit:2017:PMM

[BCCV17] Sara Bouzit, Gaëlle Calvary, Denis Chêne, and Jean Vanderdonckt. Polymodal menus: a model-based approach for designing multimodal adaptive menus for small screens. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 1 (EICS):15:1–15:19, June 2017. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3099585.

BenAjenaghughrure:2021:PMT

[BDL21] Ighoyota Ben Ajenaghughrure, Sonia Cláudia Da Costa Sousa, and David Lamas. Psychophysiological modeling of trust in technology: Influence of feature selection methods. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(EICS): 203:1–203:25, June 2021. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3459745.

Blackwell:2017:CCO

[BDSL17] Lindsay Blackwell, Jill Dimond, Sarita Schoenebeck, and Cliff Lampe. Classification and its consequences for online harassment: Design insights from HeartMob. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 1 (CSCW):24:1–24:19, December 2017. CODEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3134659.

Bowen:2021:TMI

[BDW21] Judy Bowen, Anke Dittmar, and Benjamin Weyers. Task modelling for interactive system design: a survey of historical trends, gaps and future needs. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(EICS): 214:1–214:22, June 2021. CODEN ???? ISSN 2573-0142 (electronic).

                                        URL https://dl.acm.
                                        org/doi/10.1145/3461736.

Buchholz:2017:EFT

[BF17] Gregor Buchholz and Peter Forbrig. Extended features of task models for specifying cooperative activities. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 1(EICS):7:1–7:21, June 2017. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3095809.

Budak:2017:TSH

[BGRK17] Ceren Budak, R. Kelly Garrett, Paul Resnick, and Julia Kamin. Threading is sticky: How threaded conversations promote comment system user retention. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 1 (CSCW):27:1–27:20, December 2017. CODEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3134662.

Brubaker:2019:ONG

[BHM19] Jed R. Brubaker, Gillian R.

Hayes, and Melissa Mazmanian. Orienting to networked grief: Situated perspectives of communal mourning on Facebook. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3(CSCW): 27:1–27:19, November 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3359129.

Buyuktur:2018:SCC

[BHNA18] Ayse G. Buyuktur, Pei-Yao Hung, Mark W. Newman, and Mark S. Ackerman. Supporting collaboratively constructed independence: a study of spinal cord injury. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 2(CSCW): 26:1–26:25, November 2018. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3274295.

Back:2021:PCW

[BJ21] Jon Back and Karin B. Johansson. Playing cool — winter weather’s influence on locationbased gaming. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CHI PLAY):242:1– 242:16, September 2021. CODEN ???? ISSN 2573-0142 (electronic).

                                        URL https://dl.acm.
                                        org/doi/10.1145/3474669.

Barlas:2021:SSI

[BKG^{+}21] Pinar Barlas, Kyriakos Kyriakou, Olivia Guest, Styliani Kleanthous, and Jahna Otterbacher. To “see” is to stereotype: Image tagging algorithms, gender recognition, and the accuracy– fairness trade-off. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW3):232:1–232:31, January 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3432931.

Bawa:2020:DMU

[BKJ^{+}20] Anshul Bawa, Pranav Khadpe, Pratik Joshi, Kalika Bali, and Monojit Choudhury. Do multilingual users prefer chat-bots that code-mix? Let’s nudge and find out! Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW1):041:1–041:23, May 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3392846.

Blaising:2021:MIW

[BKKD21] Allie Blaising, Yasmine Kotturi, Chinmay Kulkarni, and Laura Dabbish. Making it work, or not: a longitudinal study of career trajectories among online freelancers. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW3):226:1–226:29, January 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3432925.

Bardzell:2021:PVCa

[BKL^{+}21] Shaowen Bardzell, Juho Kim, Siân Lindley, Aleksandra Sarcevic, and Sarita Schoenebeck. PACMHCI V5 CSCW1 April 2021 editorial. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CSCW1):1:1, April 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3449254.

Benke:2020:CBE

[BKM20] Ivo Benke, Michael Thomas Knierim, and Alexander Maedche. Chatbot-based emotion management for distributed teams: a participatory design study. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW2):118:1–118:30, October 2020. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3415189.

Beres:2021:UPE

[BKM21] Nicole A. Beres, Madison Klarkowski, and Regan L. Mandryk. Under pressure: Exploring choke and clutch in competitive video games. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CHI PLAY):239:1– 239:22, September 2021. CODEN ???? ISSN 2573-0142 (electronic).

                                        URL https://dl.acm.
                                        org/doi/10.1145/3474666.

Bossen:2017:PMP

[BL17] Claus Bossen and Timo Leimbach. Project management practices as a subject of research for CSCW: Status and future opportunities. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 1 (CSCW):25:1–25:25, December 2017. CODEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3134660.

Buntain:2018:PCR

[BL18] Cody L. Buntain and Jung Kyu Rhys Lim.

#pray4victims: Con-

sistencies in response to disaster on Twitter. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 2(CSCW): 25:1–25:18, November 2018. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3274294.

Bermejo:2021:EBD

[BLC^{+}21] Carlos Bermejo, Lik Hang Lee, Paul Chojecki, David Przewozny, and Pan Hui. Exploring button designs for mid-air interaction in virtual reality: a hexa-metric evaluation of key representations and multi-modal cues. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(EICS): 194:1–194:26, June 2021. CODEN ???? ISSN 2573-0142 (electronic).

                                        URL https://dl.acm.
                                        org/doi/10.1145/3457141.

Berry:2017:GKY

[BLH^{+}17] Andrew BL Berry, Catherine Y. Lim, Andrea L. Hartzler, Tad Hirsch, Evette Ludman, Edward H. Wagner, and James D. Ralston. “It’s good to know you’re not a stranger every time”: Communication about values between patients with multiple chronic conditions and healthcare providers. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 1 (CSCW):23:1–23:20, December 2017. CODEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3134658.

Berry:2021:SCR

[BLL^{+}21] Andrew B. L. Berry, Catherine Y. Lim, Calvin A. Liang, Andrea L. Hartzler, Tad Hirsch, Dawn M. Ferguson, Zoë A. Bermet, and James D. Ralston. Supporting collaborative reflection on personal values and health. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (CSCW2):299:1–299:39, October 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3476040.

Bardzell:2021:PVCb

[BLS21] Shaowen Bardzell, Siân Lindley, and Aleksandra Sarcevic. PACMHCI V5 CSCW2 October 2021 editorial. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (CSCW2):291:1, October 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3476022.

Baumer:2021:EIG

[BMdS^{+}21] Eric P. S. Baumer, Naja L. Holten Møller, Cleidson R. B. de Souza, Casey Fiesler, Aparecido Fabiano Pinatti de Carvalho, Juliana Baptista dos Santos França, and Maria Menendez-Blanco. Editorial introduction: GROUP 2022 first wave. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(GROUP):215:1–215:2, July 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3463927.

Broccia:2020:FAS

[BMPP20] Giovanna Broccia, Marco Manca, Fabio Paternò, and Francesca Pulina. Flexible automatic support for Web accessibility validation. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (EICS):83:1–83:24, June 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3397871.

Bauer:2017:CPS

[BP17a] Aaron Bauer and Zoran Popović. Collaborative problem solving in an open-ended scientific discovery game. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 1 (CSCW):22:1–22:21, December 2017. CODEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3134657.

Brewer:2017:XRD

[BP17b] Robin N. Brewer and Anne Marie Piper. xPress: Rethinking design for aging and accessibility through an IVR blogging system. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 1 (CSCW):26:1–26:17, December 2017. CODEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3139354.

Byun:2021:CPS

[BPO21] Jeongmin Byun, Jungkook Park, and Alice Oh. Cocode: Providing social presence with co-learner screen sharing in online programming classes. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (CSCW2):300:1–300:28, October 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3476041.

Bowen:2017:GOA

[BR17] Judy Bowen and Steve Reeves. Generating obligations, assertions and tests from UI models. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 1(EICS):5:1–5:18, June 2017. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3095807.

Barnaby:2020:AEF

[BR20] Gareth Barnaby and Anne Roudaut. Autogrip: Enabling force feedback devices to selfattach to end-users’ fingers. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4(ISS): 184:1–184:14, November 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3427312.

Born:2021:MPP

[BRM21] Felix Born, Adrian Rygula, and Maic Masuch. Motivating players to perform an optional strenuous activity in a virtual reality exergame using virtual performance augmentation. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CHI PLAY):225:1– 225:21, September 2021. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3474652.

Burgess:2019:TPP

[BRN^{+}19] Eleanor R. Burgess, Kathryn E. Ringland, Jennifer Nicholas, Ashley A. Knapp, Jordan Eschler, David C. Mohr, and Madhu C. Reddy. “I think people are powerful”: The sociality of individuals managing depression. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3(CSCW): 41:1–41:29, November 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3359143.

Boyd:2021:AES

[BS21] Karen L. Boyd and Katie Shilton. Adapting ethical sensitivity as a construct to study technology design teams. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (GROUP):217:1–217:29, July 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3463929.

Baumer:2018:DRS

[BSS18] Eric P. S. Baumer, Rui Sun, and Peter Schaedler. Departing and returning: Sense of agency as an organizing concept for understanding social media non/use transitions. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 2(CSCW): 23:1–23:19, November 2018. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3274292.

Bellini:2021:CPB

[BTM^{+}21] Rosanna Bellini, Emily Tseng, Nora McDonald, Rachel Greenstadt, Damon McCoy, Thomas Ristenpart, and Nicola Dell. “So-called privacy breeds evil”: Narrative justifications for intimate partner surveillance in online forums. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW3):210:1–210:27, January 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3432909.

Badillo-Urquiola:2021:CRR

[BUSA^{+}21] Karla Badillo-Urquiola, Zachary Shea, Zainab Agha, Irina Lediaeva, and Pamela Wisniewski. Conducting risky research with teens: Codesigning for the ethical treatment and protection of adolescents. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW3):231:1–231:46, January 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3432930.

Burny:2021:UWC

[BV21] Nicolas Burny and Jean Vanderdonckt. UiLab, a workbench for conducting and reproducing experiments in GUI visual design. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(EICS): 196:1–196:31, June 2021. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3457143.

Bopp:2021:MSG

[BVM21] Julia A. Bopp, Jan B.

Vornhagen, and Elisa D. Mekler. “My Soul Got a Little Bit Cleaner”: Art experience in videogames. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CHI PLAY):237:1– 237:19, September 2021. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3474664.

Bisafar:2020:SYA

[BWDP20] Farnaz Irannejad Bisafar, Brooke Foucault Welles, Catherine D’Ignazio, and Andrea G. Parker. Supporting youth activists? strategic use of social media: a qualitative investigation of design opportunities. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW2):109:1–109:25, October 2020. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3415180.

Bowen:2020:GPP

[BWV20] Judy Bowen, Marco Winckler, and Jean Vanderdonckt. A glimpse into the past, present, and future of engineering interactive computing systems. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (EICS):71:1–71:32, June 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3394973.

Balestra:2017:IWF

[BZC^{+}17] Martina Balestra, Lior Zalmanson, Coye Cheshire, Ofer Arazy, and Oded Nov. It was fun, but did it last?: The dynamic interplay between fun motives and contributors’ activity in peer production. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 1 (CSCW):21:1–21:13, December 2017. CODEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3134656.

Bhuiyan:2020:IDC

[BZSM20] Md Momen Bhuiyan, Amy X. Zhang, Connie Moon Sehat, and Tanushree Mitra. Investigating differences in crowdsourced news credibility assessment: Raters, tasks, and expert criteria. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4(CSCW2):93:1–93:26, October 2020. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3415164.

Chan:2018:SMI

[CCH^{+}18] Joel Chan, Joseph Chee Chang, Tom Hope, Dafna Shahaf, and Aniket Kittur. SOLVENT: a mixed initiative system for finding analogies between research papers. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 2(CSCW): 31:1–31:21, November 2018. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3274300.

Chow:2019:CDC

[CCNY19] Kevin Chow, Caitlin Coyiuto, Cuong Nguyen, and Dongwook Yoon.

Challenges and design considerations for multimodal asynchronous collaboration in VR. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3 (CSCW):40:1–40:24, November 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3359142.

Chen:2018:DDP

[CCS^{+}18] Chunyang Chen, Xi Chen, Jiamou Sun, Zhenchang Xing, and Guoqiang Li. Data-driven proactive policy assurance of post quality in community Q&A sites. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 2(CSCW): 33:1–33:22, November 2018. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3274302.

Chen:2021:WAN

[CCWH21] Xinyue Chen, Si Chen, Xu Wang, and Yun Huang. “I was afraid, but now I enjoy being a streamer!”: Understanding the challenges and prospects of using live streaming for online education. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW3):237:1–237:32, January 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3432936.

Chen:2020:URI

[CCX^{+}20] Zhilong Chen, Hancheng Cao, Fengli Xu, Mengjie Cheng, Tao Wang, and Yong Li. Understanding the role of intermediaries in online social e-commerce: an exploratory study of Beidian. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW2):114:1–114:24, October 2020. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3415185.

Caraway:2017:FDN

[CEM17] Monica Caraway, Daniel A.

Epstein, and Sean A. Munson. Friends don’t need receipts: The curious case of social awareness streams in the mobile payment app venmo. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 1 (CSCW):28:1–28:17, December 2017. CODEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3134663.

Caldeira:2021:TSD

[CFD^{+}21] Clara Caldeira, Mayara Costa Figueiredo, Lucy Dodakian, Cleidson R. B. de Souza, Steven C. Cramer, and Yunan Chen. Towards supporting data-driven practices in stroke telerehabilitation technology. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CSCW1):25:1–25:33, April 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3449099.

Campos:2017:MIT

[CFG^{+}17] José Creissac Campos, Camille Fayollas, Marcelo Gonçalves, Célia Martinie, David Navarre, Philippe Palanque, and Miguel Pinto. A more intelligent test case generation approach through task models manipulation. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 1(EICS): 9:1–9:20, June 2017. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3095811.

Callaghan:2018:MHM

[CGM^{+}18] William Callaghan, Joslin Goh, Michael Mohareb, Andrew Lim, and Edith Law. MechanicalHeart: a human-machine framework for the classification of phonocardiograms. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 2(CSCW): 28:1–28:17, November 2018. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3274297.

Chan:2021:MMP

[CGS^{+}21] Tina Chan, Robert P.

Gauthier, Ally Suarez, Nicholas F. Sia, and James R. Wallace. Merlynne: Motivating peerto-peer cognitive behavioral therapy with a serious game. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CHI PLAY):250:1– 250:23, September 2021. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3474677.

Chauhan:2020:TPS

[CH20] Apoorva Chauhan and Amanda Lee Hughes. Trustworthiness perceptions of social media resources named after a crisis event. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW1):044:1–044:23, May 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3392849.

Chung:2017:PLE

[CHC^{+}17] Taejoong Chung, Jinyoung Han, Daejin Choi, Ted Taekyoung Kwon, Jong-Youn Rha, and Hyunchul Kim. Privacy leakage in eventbased social networks: a Meetup case study. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 1 (CSCW):35:1–35:22, December 2017. CODEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3134670.

Calefato:2018:ICC

[CILM18] Fabio Calefato, Giuseppe Iaffaldano, Filippo Lanubile, and Federico Maiorano. Investigating crowd creativity in online music communities. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 2(CSCW): 27:1–27:21, November 2018. CODEN ????

ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3274296.

Cha:2018:CAU

[CKP^{+}18] Yoonjeong Cha, Jongwon Kim, Sangkeun Park, Mun Yong Yi, and Uichin Lee. Complex and ambiguous: Understanding sticker misinterpretations in instant messaging. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 2(CSCW): 30:1–30:22, November 2018. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3274299.

Chiang:2018:CCP

[CKS18] Chun-Wei Chiang, Anna Kasunic, and Saiph Savage. Crowd coach: Peer coaching for crowd workers’ skill growth. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 2 (CSCW):37:1–37:17, November 2018. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3274306.

Cardoso-Llach:2021:AIE

[CLKEF21] Daniel Cardoso-Llach, Eric Kaltman, Emek Erdolu, and Zachary Furste. An archive of interfaces: Exploring the potential of emulation for software research, pedagogy, and design. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (CSCW2):294:1–294:22, October 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3476035.

Cuadra:2021:MBR

[CLL^{+}21] Andrea Cuadra, Shuran Li, Hansol Lee, Jason Cho, and Wendy Ju. My bad! repairing intelligent voice assistant errors improves interaction. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CSCW1):27:1–27:24, April 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3449101.

Chung:2021:UHS

[CLP^{+}21] Chaeyeon Chung, Jungsoo Lee, Kyungmin Park, Junsoo Lee, Minjae Kim, Mookyung Song, Yeonwoo Kim, Jaegul Choo, and Sungsoo Ray Hong. Understanding humanside impact of sampling image batches in subjective attribute labeling. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (CSCW2):296:1–296:26, October 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3476037.

Coppers:2019:FFA

[CLV^{+}19] Sven Coppers, Kris Luyten, Davy Vanacken, David Navarre, Philippe Palanque, and Christine Gris. Fortunettes: Feedforward about the future state of GUI widgets. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3 (EICS):20:1–20:20, June 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3331162.

Chen:2020:EUD

[CLwH^{+}20] Yu-Chun Chen, Chia-Ying Liao, Shuo wen Hsu, Da-Yuan Huang, and Bing-Yu Chen. Exploring user defined gestures for ear-based interactions. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4(ISS): 186:1–186:20, November 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3427314.

Chevalier:2020:PHV

[CM20] Fanny Chevalier and Nicolai Marquardt. PACM HCI V4 ISS, November 2020 — editorial. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (ISS):182:1, November 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3427310.

Canny:2021:EMB

[CMN^{+}21] Alexandre Canny, Célia Martinie, David Navarre, Philippe Palanque, Eric Barboni, and Christine Gris. Engineering modelbased software testing of WIMP interactive applications: a process based on formal models and the SQUAMATA tool. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(EICS): 207:1–207:30, June 2021. CODEN ???? ISSN 2573-0142 (electronic).

                                        URL https://dl.acm.
                                        org/doi/10.1145/3461729.

Choi:2021:PSC

[CMP^{+}21] Yoonseo Choi, Toni-Jan Keith Palma Monserrat, Jeongeon Park, Hyungyu Shin, Nyoungwoo Lee, and Juho Kim. ProtoChat: Supporting the conversation design process with crowd feedback. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW3):225:1–225:27, January 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3432924.

Chua:2017:KFC

[CMY^{+}17] Soon Hau Chua, Toni-Jan Keith Palma Monserrat, Dongwook Yoon, Juho Kim, and Shengdong Zhao. Korero: Facilitating complex referencing of visual materials in asynchronous discussion interface. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 1 (CSCW):34:1–34:19, December 2017. CODEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3134669.

Calvary:2017:ES

[CNC^{+}17] Gaelle Calvary, Jeffrey Nichols, José Creissac Campos, Nuno Nunes, and Pedro Campos. Editorial statement. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 1(EICS):1:1–1:2, June 2017. CODEN ????

ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3106386.

Chounta:2017:WSE

[CNH^{+}17] Irene-Angelica Chounta, Alexander Nolte, Tobias Hecking, Rosta Farzan, and Thomas Herrmann. When to say “Enough is Enough!”: a study on the evolution of collaboratively created process models. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 1 (CSCW):33:1–33:21, December 2017. CODEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3134668.

Cho:2020:SYC

[COKL20] Hyunsung Cho, Jinyoung Oh, Juho Kim, and Sung-Ju Lee. I share, you care: Private status sharing and sender-controlled notifications in mobile instant messaging. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW1):034:1–034:25, May 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3392839.

Casarin:2018:UUT

[CPB18] Julien Casarin, Nicolas Pacqueriaud, and Dominique Bechmann. UMI3D: a Unity3D toolbox to support CSCW systems properties in generic 3D user interfaces. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 2(CSCW): 29:1–29:20, November 2018. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3274298.

Chandrasekharan:2017:YCS

[CPS^{+}17] Eshwar Chandrasekharan, Umashanthi Pavalanathan, Anirudh Srinivasan, Adam Glynn, Jacob Eisenstein, and Eric Gilbert. You can’t stay here: The efficacy of Reddit’s 2015 ban examined through hate speech. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 1 (CSCW):31:1–31:22, December 2017. CODEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3134666.

Cho:2020:RCG

[CR20] Janghee Cho and Emilee Rader. The role of conversational grounding in supporting symbiosis between people and digital assistants. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW1):033:1–033:28, May 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3392838.

Chakroborti:2021:DRI

[CRN21] Debasish Chakroborti, Banani Roy, and Sristy Sumana Nath. Designing for recommending intermediate states in a scientific workflow management system. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(EICS): 198:1–198:29, June 2021. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3457145.

Carcangiu:2018:GGG

[CS18a] Alessandro Carcangiu and Lucio Davide Spano. G-Gene: a gene alignment method for online partial stroke gestures recognition. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 2 (EICS):13:1–13:17, June 2018. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3229095.

Cheon:2018:SLN

[CS18b] EunJeong Cheon and Norman Makoto Su. ’Staged for Living’: Negotiating objects and their values over a porous boundary. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 2(CSCW): 36:1–36:24, November 2018. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3274305.

Canossa:2021:HTD

[CSA^{+}21] Alessandro Canossa, Dmitry Salimov, Ahmad Azadvar, Casper Harteveld, and Georgios Yannakakis. For honor, for toxicity: Detecting toxic behavior through gameplay. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CHI PLAY):253:1– 253:29, September 2021. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3474680.

Chandrasekharan:2018:IHR

[CSJ^{+}18] Eshwar Chandrasekharan, Mattia Samory, Shagun Jhaver, Hunter Charvat, Amy Bruckman, Cliff Lampe, Jacob Eisenstein, and Eric Gilbert. The Internet’s hidden rules: an empirical study of Reddit norm violations at micro, meso, and macro scales. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 2(CSCW): 32:1–32:25, November 2018. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3274301.

Chakraborty:2017:TES

[CSMG17] Abhijnan Chakraborty, Rajdeep Sarkar, Ayushi Mrigen, and Niloy Ganguly. Tabloids in the era of social media?: Understanding the production and consumption of clickbaits in Twitter. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 1 (CSCW):30:1–30:21, December 2017. CODEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3134665.

Chhibber:2021:TNK

[CSMV21] Nalin Chhibber, Hemant Bhaskar Surale, Fabrice Matulic, and Daniel Vogel. Typealike: Near-keyboard hand postures for expanded laptop interaction. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(ISS): 486:1–486:20, November 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3486952.

Cho:2019:IEB

[CSN19] Bryden Cho, Chengzheng Sun, and Agustina Ng. Issues and experiences in building heterogeneous co-editing systems. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3 (GROUP):245:1–245:28, December 2019. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3361126.

Cartwright:2017:SSI

[CSS^{+}17] Mark Cartwright, Ayanna Seals, Justin Salamon, Alex Williams, Stefanie Mikloska, Duncan Mac-Connell, Edith Law, Juan P. Bello, and Oded Nov. Seeing sound: Investigating the effects of visualizations and complexity on crowdsourced audio annotations. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 1 (CSCW):29:1–29:21, December 2017. CODEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3134664.

Choi:2018:WTM

[CT18] Joohee Choi and Yla Tausczik. Will too many editors spoil the tag?: Conflicts and alignment in Q&A categorization. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 2(CSCW): 38:1–38:19, November 2018. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3274307.

Compton:2018:LPU

[CWB^{+}18] Ryan J. Compton, Jeff Warshaw, Hernan Badenes, Barton Smith, and Steve Whittaker. Living in the present: Understanding long-term content referencing in enterprise online communities. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 2(CSCW): 39:1–39:21, November 2018. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3274308.

Chen:2017:CCD

[CXL17] Chunyang Chen, Zhenchang Xing, and Yang Liu. By the community & for the community: a deep learning approach to assist collaborative editing in Q&A sites. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 1 (CSCW):32:1–32:21, December 2017. CODEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3134667.

Cao:2021:MTW

[CYC^{+}21] Hancheng Cao, Vivian Yang, Victor Chen, Yu Jin Lee, Lydia Stone, N’godjigui Junior Diarrassouba, Mark E. Whiting, and Michael S. Bernstein. My team will go on: Differentiating high and low viability teams through team interaction. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW3):230:1–230:27, January 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3432929.

Chau:2021:CLD

[CYGW21] Edwin Chau, Jiakun Yu, Cagatay Goncu, and Anusha Withana. Composite line designs and accuracy measurements for tactile line tracing on touch surfaces. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(ISS): 491:1–491:17, November 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3488536.

Chen:2018:BLF

[CYW18] Mei-Ling Chen, Naomi Yamashita, and Hao-Chuan Wang. Beyond lingua franca: Systemfacilitated language switching diversifies participation in multiparty multilingual communication. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 2(CSCW): 34:1–34:22, November 2018. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3274303.

Cardieri:2019:PEE

[dACZ19] Giulia de Andrade Cardieri and Luciana A. M. Zaina. PWA-EU: Extending PWA approach for promoting customization based on user preferences. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3 (EICS):10:1–10:28, June 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3331152.

Dittus:2017:CET

[DAQ17] Martin Dittus, Luca Maria Aiello, and Daniele Quercia. Community engagement triage: Lightweight prompts for systematic reviews. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 1(CSCW): 39:1–39:22, December 2017. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3134674.

Dechant:2021:HAC

[DBS^{+}21] Martin Johannes Dechant, Max V. Birk, Youssef Shiban, Knut Schnell, and Regan L. Mandryk. How avatar customization affects fear in a game-based digital exposure task for social anxiety. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CHI PLAY):248:1– 248:27, September 2021. CODEN ???? ISSN 2573-0142 (electronic).

                                        URL https://dl.acm.
                                        org/doi/10.1145/3474675.

Dittus:2017:PPF

[DC17] Martin Dittus and Licia Capra. Private peer feedback as engagement driver in humanitarian mapping. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 1(CSCW): 40:1–40:18, December 2017. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3134675.

Dey:2019:VIA

[DDFK19] Sanorita Dey, Brittany R. L. Duff, Wai-Tat Fu, and Karrie Karahalios. VidLyz: an interactive approach to assist novice entrepreneurs in making persuasive campaign videos. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3(CSCW): 43:1–43:26, November 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3359145.

DAngelo:2018:SCR

[DDZ18] Gabriele D’Angelo, Angelo Di Iorio, and Stefano Zacchiroli. Spacetime characterization of real-time collaborative editing. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 2(CSCW): 41:1–41:19, November 2018. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3274310.

Drey:2021:SIC

[DFJ^{+}21] Tobias Drey, Fabian Fischbach, Pascal Jansen, Julian Frommel, Michael Rietzler, and Enrico Rukzio. To be or not to be stuck, or is it a continuum?: a systematic literature review on the concept of being stuck in games. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CHI PLAY):229:1– 229:35, September 2021. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3474656.

Davis:2018:ECP

[DHH18] Dan Davis, Claudia Hauff, and Geert-Jan Houben. Evaluating crowdworkers as a proxy for online learners in video-based learning contexts. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 2(CSCW): 42:1–42:16, November 2018. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3274311.

Dubois:2020:GDG

[DMC^{+}20] Patrick Marcel Joseph Dubois, Mahya Maftouni, Parmit K. Chilana, Joanna McGrenere, and Andrea Bunt. Gender differences in graphic design Q&As: How community and site characteristics contribute to gender gaps in answering questions. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW2):113:1–113:26, October 2020. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3415184.

Datta:2017:IMI

[DPA17] Srayan Datta, Chanda Phelan, and Eytan Adar. Identifying misaligned inter-group links and communities. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 1 (CSCW):37:1–37:23, December 2017. CODEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3134672.

Proenca:2021:CDS

[dQPMdHRdAN21] Mailson de Queiroz Proença, Vivian Genaro Motti, Kamila Rios da Hora Rodrigues, and Vânia Paula de Almeida Neris. Coping with diversity — a system for end-users to customize Web user interfaces. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(EICS): 201:1–201:27, June 2021. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3457151.

Dosono:2017:CTC

[DRA^{+}17] Bryan Dosono, Yasmeen Rashidi, Taslima Akter, Bryan Semaan, and Apu Kapadia. Challenges in transitioning from civil to military culture: Hyper-selective disclosure through ICTs. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 1 (CSCW):41:1–41:23, December 2017. CODEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3134676.

Deri:2018:CLC

[DRAQ18] Sebastian Deri, Jeremie Rappaz, Luca Maria Aiello, and Daniele Quercia. Coloring in the links: Capturing social ties as they are perceived. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 2(CSCW): 43:1–43:18, November 2018. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3274312.

Durkin:2018:SGF

[DRK18] Celia Durkin, Federico Rossano, and Scott Klemmer. Score-group framing negatively impacts peer evaluations. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 2 (CSCW):45:1–45:19, November 2018. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3274314.

Dessi:2020:DEG

[DS20] Stefano Dessì and Lucio Davide Spano. DG3: Exploiting gesture declarative models for sample generation and online recognition. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (EICS):82:1–82:21, June 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3397870.

DeVito:2018:TGF

[DWB18] Michael A. DeVito, Ashley Marie Walker, and Jeremy Birnholtz. ’Too Gay for Facebook’: Presenting LGBTQ+ identity throughout the personal social media ecosystem. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 2(CSCW): 44:1–44:23, November 2018. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3274313.

Dickson:2021:HTS

[DWMV21] Terence Dickson, Rina R.

Wehbe, Fabrice Matulic, and Daniel Vogel. HybridPointing for touch: Switching between absolute and relative pointing on large touch screens. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(ISS): 495:1–495:22, November 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3488540.

Dillahunt:2017:SEC

[DWW^{+}17] Tawanna R. Dillahunt, Xinyi Wang, Earnest Wheeler, Hao Fei Cheng, Brent Hecht, and Haiyi Zhu. The sharing economy in computing: a systematic literature review. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 1 (CSCW):38:1–38:26, December 2017. CODEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3134673.

Dian:2020:TDI

[DWZ^{+}20] Cao Dian, Dong Wang, Qian Zhang, Run Zhao, and Yinggang Yu. Towards domain-independent complex and fine-grained gesture recognition with RFID. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4(ISS): 187:1–187:22, November 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3427315.

Early:2018:UGE

[EHH^{+}18] Kirstin Early, Jessica Hammer, Megan Kelly Hofmann, Jennifer A. Rode, Anna Wong, and Jennifer Mankoff. Understanding gender equity in author order assignment. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 2(CSCW): 46:1–46:21, November 2018. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3274315.

Epstein:2020:EDP

[EJB^{+}20] Daniel A. Epstein, Siyun Ji, Danny Beltran, Griffin D’Haenens, Zhaomin Li, and Tan Zhou. Exploring design principles for sharing of personal informatics data on ephemeral social media. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4(CSCW2):95:1–95:24, October 2020. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3415166.

Ernala:2017:LMI

[ERB^{+}17] Sindhu Kiranmai Ernala, Asra F. Rizvi, Michael L. Birnbaum, John M. Kane, and Munmun De Choudhury. Linguistic markers indicating therapeutic outcomes of social media disclosures of schizophrenia. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 1 (CSCW):43:1–43:27, December 2017. CODEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3134678.

Epstein:2017:SSE

[ERLW17] Robert Epstein, Ronald E. Robertson, David Lazer, and Christo Wilson. Suppressing the search engine manipulation effect (SEME). Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 1 (CSCW):42:1–42:22, December 2017. CODEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3134677.

Erete:2021:CBR

[ERT21] Sheena Erete, Yolanda A. Rankin, and Jakita O. Thomas. I can’t breathe: Reflections from black women in CSCW and HCI. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4(CSCW3): 234:1–234:23, January 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3432933.

Ertl:2019:SMH

[ETE^{+}19] Tanja Ertl, Sebastian Taugerbeck, Margarita Esau, Konstantin Aal, Peter Tolmie, and Volker Wulf. The social mile — how (psychosocial)

ICT can help to promote resocialization and to overcome prison. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3 (GROUP):248:1–248:31, December 2019. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3370270.

Erickson:2019:EW

[EVC^{+}19] Ingrid Erickson, Adriana S. Vivacqua, Lars Rune Christensen, Naja L. Holten Moller, Eric Baumer, Yvette Wohn, Louise Barkhuus, and Lionel P. Robert. Editors’ welcome. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3(GROUP): 229:1–229:2, December 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3361110.

Fiesler:2019:CCC

[FB19] Casey Fiesler and Amy S.

Bruckman. Creativity, copyright, and close-Knit communities: a case study of social norm formation and enforcement. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3 (GROUP):241:1–241:24, December 2019. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3361122.

Fetter:2017:TJT

[FBG17] Mirko Fetter, David Bimamisa, and Tom Gross. TUIOFX: a JavaFX toolkit for shared interactive surfaces. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 1 (EICS):10:1–10:18, June 2017. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3095812.

Fan:2020:IHU

[FC20] Zhuzhi Fan and Céline Coutrix. Impact of hand used on one-handed back-of-device performance. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4(ISS): 188:1–188:19, November 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3427316.

Figueiredo:2018:EHD

[FCE^{+}18] Mayara Costa Figueiredo, Clara Caldeira, Elizabeth Victoria Eikey, Melissa Mazmanian, and Yunan Chen. Engaging with health data:

The interplay between self-tracking activities and emotions in fertility struggles. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 2(CSCW): 40:1–40:20, November 2018. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3274309.

Faas:2019:JTJ

[FcLDM19] Travis Faas, I ching Liu, Lynn Dombrowski, and Andrew D. Miller. Jam today, jam tomorrow: Learning in online game jams. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3 (GROUP):240:1–240:27, December 2019. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3361121.

Figueiredo:2017:STF

[FCR^{+}17] Mayara Costa Figueiredo, Clara Caldeira, Tera L. Reynolds, Sean Victory, Kai Zheng, and Yunan Chen. Self-tracking for fertility care: Collaborative support for a highly personalized problem. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 1 (CSCW):36:1–36:21, December 2017. CODEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3134671.

Fiesler:2020:MAL

[FD20] Casey Fiesler and Brianna Dym. Moving across lands: Online platform migration in fandom communities. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW1):042:1–042:25, May 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3392847.

Foong:2017:NES

[FGG17] Eureka Foong, Darren Gergle, and Elizabeth M. Gerber. Novice and expert sensemaking of crowdsourced design feedback. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 1 (CSCW):45:1–45:18, December 2017. CODEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3134680.

Fiesler:2019:ECR

[Fie19]  Casey Fiesler.  Ethi-

cal considerations for research involving (speculative) public data. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3 (GROUP):249:1–249:13, December 2019. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3370271.

Fjeld:2021:PVI

[FJIH21] Morten Fjeld, Hans-Christian Jetter, Petra Isenberg, and Mark Hancock. PACMHCI V5, ISS, November 2021 editorial. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (ISS):483:1, November 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3486933.

Foong:2021:CUH

[FKDG21] Eureka Foong, Joy O.

Kim, Mira Dontcheva, and Elizabeth M. Gerber. CrowdFolio: Understanding how holistic and decomposed workflows influence feedback on online portfolios. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CSCW1):22:1–22:31, April 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3449096.

Frey:2019:PDW

[FKK19] Seth Frey, P. M. Krafft, and Brian C. Keegan. “This Place Does What It Was Built For”: Designing digital institutions for participatory change. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3(CSCW): 32:1–32:31, November 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3359134.

Faucett:2017:SLM

[FLC17] Heather A. Faucett, Matthew L. Lee, and Scott Carter. I should listen more: Real-time sensing and feedback of non-verbal communication in video telehealth. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 1 (CSCW):44:1–44:19, December 2017. CODEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3134679.

Fruchard:2020:SCM

[FLC20] Bruno Fruchard, Eric Lecolinet, and Olivier Chapuis. Side-crossing menus: Enabling large sets of gestures for small surfaces. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4(ISS): 189:1–189:19, November 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3427317.

Funk:2021:DEI

[FLP^{+}21] Mathias Funk, Rong-Hao Liang, Philippe Palanque, Jun Hu, and Panos Markopoulos. Designing and engineering interactive computing systems. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(EICS): 193:1–193:4, June 2021. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3457140.

Feldman:2021:HWW

[FM21] Molly Q. Feldman and Brian James McInnis. How we write with crowds. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4(CSCW3): 229:1–229:31, January 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3432928.

Figueiredo:2021:UDA

[FSC21] Mayara Costa Figueiredo, H. Irene Su, and Yunan Chen. Using data to approach the unknown: Patients’ and healthcare providers? Data practices in fertility challenges. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW3):227:1–227:35, January 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3432926.

Feinberg:2020:NRR

[FSN^{+}20] Melanie Feinberg, Will Sutherland, Sarah Beth Nelson, Mohammad Hossein Jarrahi, and Arcot Rajasekar. The new reality of reproducibility: The role of data work in scientific research. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW1):035:1–035:22, May 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3392840.

Feuston:2020:CED

[FTP20] Jessica L. Feuston, Alex S. Taylor, and Anne Marie Piper. Conformity of eating disorders through content moderation. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW1):040:1–040:28, May 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3392845.

Garg:2021:UFN

[Gar21] Radhika Garg. Understanding families’ Non-/ use practices and choices: The case of smart speakers and smart interactive toys. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (CSCW2):295:1–295:26, October 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3476036.

Gligoric:2019:CEB

[GAW19] Kristina Gligorić, Ashton Anderson, and Robert West. Causal effects of brevity on style and success in social media. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3(CSCW): 45:1–45:23, November 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3359147.

Goncalves:2018:KBK

[GB18] Afonso Gonçalves and Sergi Bermúdez. KAVE: Building Kinect based CAVE automatic virtual environments, methods for surround-screen projection management, motion parallax and fullbody interaction support. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 2(EICS): 10:1–10:15, June 2018. CODEN ???? ISSN 2573-0142 (electronic).

                                        URL https://dl.acm.
                                        org/doi/10.1145/3229092.

Greipl:2021:FBE

[GBN21] Simon Greipl, Katharina Bernecker, and Manuel Ninaus. Facial and bodily expressions of emotional engagement: How dynamic measures reflect the use of game elements and subjective experience of emotions and effort. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CHI PLAY):240:1– 240:25, September 2021. CODEN ???? ISSN 2573-0142 (electronic).

                                        URL https://dl.acm.
                                        org/doi/10.1145/3474667.

Green:2019:PLA

[GC19] Ben Green and Yiling Chen. The principles and limits of algorithmin-the-loop decision making. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3(CSCW): 50:1–50:24, November 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3359152.

Gardey:2020:UPA

[GGF^{+}20] Juan Cruz Gardey, Alejandra Garrido, Sergio Firmenich, Julián Grigera, and Gustavo Rossi. UX-Painter: an approach to explore interaction fixes in the browser. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (EICS):89:1–89:21, June 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3397877.

Galeote:2021:GBC

[GH21] Daniel Fernández Galeote and Juho Hamari. Game-based climate change engagement: Analyzing the potential of entertainment and serious games. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CHI PLAY):226:1– 226:21, September 2021. CODEN ???? ISSN 2573-0142 (electronic).

                                        URL https://dl.acm.
                                        org/doi/10.1145/3474653.

Gu:2021:LLD

[GHHC21] Hongyan Gu, Jingbin Huang, Lauren Hung, and Xiang ’Anthony’ Chen. Lessons learned from designing an AI-Enabled diagnosis tool for pathologists. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CSCW1):10:1–10:25, April 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3449084.

Guttman:2021:PRI

[GHHS21] Rotem D. Guttman, Jessica Hammer, Erik Harpstead, and Carol J. Smith. Play for real(ism)

— using games to predict Human-AI interactions in the real world. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CHI PLAY):228:1– 228:17, September 2021. CODEN ???? ISSN 2573-0142 (electronic).

                                        URL https://dl.acm.
                                        org/doi/10.1145/3474655.

Gilbert:2020:RWL

[Gil20] Sarah A. Gilbert. “I run the world’s largest historical outreach project and it’s on a cesspool of a website.” moderating a public scholarship site on Reddit: a case study of r/AskHistorians. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW1):019:1–019:27, May 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3392822.

Garg:2021:ULE

[GKS21] Radhika Garg, Yash Kapadia, and Subhasree Sengupta. Using the lenses of emotion and support to understand unemployment discourse on Reddit. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CSCW1):14:1–14:24, April 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3449088.

Ghai:2021:EAL

[GLZ^{+}21] Bhavya Ghai, Q. Vera Liao, Yunfeng Zhang, Rachel Bellamy, and Klaus Mueller. Explainable active learning (XAL): Toward AI explanations as interfaces for machine teachers. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW3):235:1–235:28, January 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3432934.

Griggio:2019:CEB

[GMM19] Carla F. Griggio, Joanna McGrenere, and Wendy E. Mackay. Customizations and expression breakdowns in ecosystems of communication apps. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3(CSCW): 26:1–26:26, November 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3359128.

Gerling:2021:PVC

[GMM21] Kathrin Gerling, Elisa Mekler, and Regan L. Mandryk. PACMHCI V5, CHI PLAY, September 2021 editorial. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CHI PLAY):224:1– 224:2, September 2021. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3474651.

Gron:2020:PPV

[GN20] Kirsikka Grön and Matti Nelimarkka. Party politics, values and the design of social media services: Implications of political elites’ values and ideologies to mitigating of political polarisation through design. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW2):104:1–104:29, October 2020. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3415175.

Grandhi:2019:DSD

[GPH19] Sukeshini A. Grandhi, Linda Plotnick, and Starr Roxanne Hiltz. Do I stay or do I go?: Motivations and decision making in social media non-use and reversion. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3 (GROUP):235:1–235:27, December 2019. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3361116.

Grandhi:2021:CCP

[GPH21] Sukeshini Grandhi, Linda Plotnick, and Starr Roxanne Hiltz. By the crowd and for the crowd: Perceived utility and willingness to contribute to trustworthiness indicators on social media. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (GROUP):218:1–218:24, July 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3463930.

Guettaf:2020:EPC

[GRG20] Adnane Guettaf, Yosra Rekik, and Laurent Grisoni. Effect of physical challenging activity on tactile texture recognition for mobile surface. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4(ISS): 190:1–190:12, November 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3427318.

Gong:2021:SDM

[GSC21] Crystal Gong, Koustuv Saha, and Stevie Chancellor. “The Smartest Decision for My Future”: Social media reveals challenges and stress during post-college life transition. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (CSCW2):298:1–298:29, October 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3476039.

Greis:2017:ICE

[GSK^{+}17] Miriam Greis, Hendrik Schuff, Marius Kleiner, Niels Henze, and Albrecht Schmidt. Input controls for entering uncertain data: Probability distribution sliders. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 1(EICS):3:1–3:17, June 2017. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3095805.

Grolaux:2020:STS

[GVNK20] Donatien Grolaux, Jean Vanderdonckt, Thanh-Diane Nguyen, and Iyad Khaddam. SketchADoodle: Touch-surface multistroke gesture handling by Bézier curves. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (EICS):87:1–87:30, June 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3397875.

Garcia:2017:MDA

[GYD17] Adriana Alvarado Garcia, Alyson L. Young, and Lynn Dombrowski. On making data actionable: How activists use imperfect data to foster social change for human rights violations in Mexico. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 1(CSCW): 19:1–19:19, December 2017. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3134654.

Hao:2021:CGS

[Hao21]  Yu Hao.  Computer

games as social sculptures: Rethinking the discourse of participation and its implications for digital game design. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CHI PLAY):241:1– 241:15, September 2021. CODEN ???? ISSN 2573-0142 (electronic).

                                        URL https://dl.acm.
                                        org/doi/10.1145/3474668.

Heinz:2021:DTA

[HBJR21] Mario Heinz, Sebastian Büttner, Sascha Jenderny, and Carsten Röcker. Dynamic task allocation based on individual abilities — experiences from developing and operating an inclusive assembly line for workers with and without disabilities. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(EICS): 206:1–206:19, June 2021. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3461728.

Hirskyj-Douglas:2021:FDI

[HDPL21] Ilyena Hirskyj-Douglas, Roosa Piitulainen, and Andrés Lucero. Forming the dog Internet: Prototyping a dog-to-human video call device. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(ISS): 494:1–494:20, November 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3488539.

Helms:2019:DCR

[HFBL19] Karey Helms, Pedro Ferreira, Barry Brown, and Airi Lampinen. Away and (dis)connection: Reconsidering the use of digital technologies in light of long-term outdoor activities. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3 (GROUP):230:1–230:20, December 2019. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3361111.

HoltenMoller:2019:ACC

[HFL19] Naja L. Holten Møller, Geraldine Fitzpatrick, and Christopher A. Le Dantec. Assembling the case: Citizens’ strategies for exercising authority and personal autonomy in social welfare. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3 (GROUP):244:1–244:21, December 2019. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3361125.

Hesenius:2019:GHG

[HG19] Marc Hesenius and Volker Gruhn. GestureCards: a hybrid gesture notation. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3 (EICS):22:1–22:35, June 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3331164.

Huber:2020:PFS

[HGG20] Stephan Huber, Johanna Gramlich, and Tobias Grundgeiger. From paper flight strips to digital strip systems: Changes and similarities in air traffic control work practices. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW1):028:1–028:21, May 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3392833.

Higashi:2020:ESI

[Hig20] Takafumi Higashi. Evaluation of skill improvement by combining skill and difficulty levels during paper-cutting production. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4(ISS): 191:1–191:17, November 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3427319.

Hussein:2020:MMV

[HJM20] Eslam Hussein, Prerna Juneja, and Tanushree Mitra. Measuring misinformation in video search platforms: an audit study on YouTube. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW1):048:1–048:27, May 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3392854.

Hwang:2019:WDH

[HKF19] Sun Young Hwang, Negar Khojasteh, and Susan R. Fussell. When delayed in a hurry: Interpretations of response delays in time-sensitive instant messaging. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3 (GROUP):234:1–234:20, December 2019. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3361115.

Haldar:2020:PAS

[HKM^{+}20] Shefali Haldar, Yoojung Kim, Sonali R. Mishra, Andrea L. Hartzler, Ari H. Pollack, and Wanda Pratt. The patient advice system: a technology probe study to enable peer support in the hospital. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW2):112:1–112:23, October 2020. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3415183.

Haapoja:2021:PSS

[HLV21] Jesse Haapoja, Airi Lampinen, and Kari Mikko Vesala. Personalised services in social situations: Principal-agent relationships in account sharing. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW3):219:1–219:21, January 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3432918.

Hanrahan:2021:AEE

[HMB^{+}21] Benjamin V. Hanrahan, Carleen Maitland, Timothy Brown, Anita Chen, Fraterne Kagame, and Beatrice Birir. Agency and extraction in emerging industrial drone applications: Imaginaries of rwandan farm workers and community members. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW3):233:1–233:21, January 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3432932.

Hoppe:2021:OHH

[HOSK21] Matthias Hoppe, Daria Oskina, Albrecht Schmidt, and Thomas Kosch.

Odin’s helmet: a headworn haptic feedback device to simulate Gforces on the human body in virtual reality. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(EICS): 212:1–212:15, June 2021. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3461734.

Herman:2021:MTQ

[HOZ^{+}21] Bridger Herman, Maxwell Omdal, Stephanie Zeller, Clara A. Richter, Francesca Samsel, Greg Abram, and Daniel F. Keefe. Multi-touch querying on data physicalizations in immersive AR. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (ISS):497:1–497:20, November 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3488542.

Hougaard:2021:WWI

[HRC^{+}21] Bastian Ilsø Hougaard, Ingeborg Goll Rossau, Jedrzej Jacek Czapla, Mózes Adorján Mikó, Rasmus Bugge Skammelsen, Hendrik Knoche, and Mads Jochumsen. Who willed it?: Decreasing frustration by manipulating perceived control through fabricated input for stroke rehabilitation BCI games. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CHI PLAY):235:1– 235:19, September 2021. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3474662.

Hossain:2020:VVG

[HRRS20] Muhammad M. Hossain, Banani Roy, Chanchal K. Roy, and Kevin A. Schneider. VizSciFlow: a visually guided scripting framework for supporting complex scientific data analysis. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (EICS):74:1–74:37, June 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3394976.

Huber:2019:AAB

[HSG19] Bernd Huber, Stuart Shieber, and Krzysztof Z. Gajos. Automatically analyzing brainstorming language behavior with meeter. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3(CSCW):

30:1–30:17, November 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3359132.

Haesler:2021:STH

[HSVR21] Steffen Haesler, Stefka Schmid, Annemike Sophia Vierneisel, and Christian Reuter. Stronger together: How neighborhood groups build up a virtual network during the COVID-19 pandemic. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (CSCW2):304:1–304:31, October 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3476045.

Hettiachchi:2020:CCS

[HvBKG20] Danula Hettiachchi, Niels van Berkel, Vassilis Kostakos, and Jorge Goncalves. Crowd-Cog: a cognitive skill based system for heterogeneous task assignment and recommendation in crowdsourcing. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW2):110:1–110:22, October 2020. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3415181.

Homaeian:2021:JAS

[HWS21] Leila Homaeian, James R.

Wallace, and Stacey D. Scott. Joint action storyboards: a framework for visualizing communication grounding costs. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CSCW1):28:1–28:27, April 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3449102.

Hwang:2021:HSC

[HWYW21] Angel Hsing-Chi Hwang, Cheng Yao Wang, Yao-Yuan Yang, and Andrea Stevenson Won. Hide and seek: Choices of virtual backgrounds in video chats and their effects on perception. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (CSCW2):303:1–303:30, October 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3476044.

Ikematsu:2020:IHS

[IOES20] Kaori Ikematsu, Haruna Oshima, Rachel Eardley, and Itiro Siio. Investigating how smartphone movement is affected by lying down body posture. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4(ISS): 192:1–192:17, November 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3427320.

Ibtasam:2019:MCB

[IRA^{+}19] Samia Ibtasam, Lubna Razaq, Maryam Ayub, Jennifer R. Webster, Syed Ishtiaque Ahmed, and Richard Anderson. “My cousin bought the phone for me. I never go to mobile shops.”: The role of family in women’s technological inclusion in Islamic culture. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3(CSCW): 46:1–46:33, November 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3359148.

Jones:2021:LFM

[JA21] Jasmine Jones and Mark S. Ackerman. Learning from family mysteries: Accounting for untold stories in family memory practices. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW3):221:1–221:22, January 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3432920.

Jensen:2019:SWO

[JAB19] Jonas Kjeldmand Jensen, Tawfiq Ammari, and Pernille Bjørn. Into Scandinavia: When online fatherhood reflects societal infrastructures. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3 (GROUP):231:1–231:21, December 2019. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3361112.

Ju:2019:AMS

[JCD19] Qinjie Ju, René Chalon, and Stéphane Derrode. Assisted music score reading using fixed-gaze head movement: Empirical experiment and design implications. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3(EICS):3:1–3:29, June 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3300962.

Jasim:2021:CCR

[JKW^{+}21] Mahmood Jasim, Pooya Khaloo, Somin Wadhwa, Amy X. Zhang, Ali Sarvghad, and Narges Mahyar. CommunityClick: Capturing and reporting community feedback from town halls to improve inclusivity. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW3):213:1–213:32, January 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3432912.

Jackson:2020:SFE

[JØC^{+}20] Corey Brian Jackson, Carsten Østerlund, Kevin Crowston, Mahboobeh Harandi, and Laura Trouille. Shifting forms of engagement: Volunteer learning in online citizen science. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW1):036:1–036:19, May 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3392841.

Jun:2021:EES

[JSP^{+}21] Joonyoung Jun, Woosuk Seo, Jihyeon Park, Subin Park, and Hyunggu Jung. Exploring the experiences of streamers with visual impairments. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (CSCW2):297:1–297:23, October 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3476038.

Jaquero:2019:NGE

[JVP^{+}19] Víctor Manuel López Jaquero, Radu-Daniel Vatavu, Jose Ignacio Panach, Oscar Pastor, and Jean Vanderdonckt. A Newcomer’s guide to EICS, the engineering interactive computing systems community. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3(EICS):1:1–1:9, June 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3300960.

Jahanbakhsh:2021:ELI

[JZB^{+}21] Farnaz Jahanbakhsh, Amy X. Zhang, Adam J. Berinsky, Gordon Pennycook, David G. Rand, and David R. Karger. Exploring lightweight interventions at posting time to reduce the sharing of misinformation on social media. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CSCW1):18:1–18:42, April 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3449092.

Kotturi:2021:UCC

[KBFK21] Yasmine Kotturi, Allie Blaising, Sarah E. Fox, and Chinmay Kulkarni. The unique challenges for creative small businesses seeking feedback on social media. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CSCW1):15:1–15:27, April 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3449089.

Kapuscinska:2021:SEA

[KBKH21] Adela Kapuscinska, Payal M. Bhujwala, Melissa Kalarchian, and Jessica Hammer. A socio-ecological approach to activity games for girls. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CHI PLAY):246:1– 246:28, September 2021. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3474673.

Kar:2020:GEU

[KCC20] Pragma Kar, Samiran Chattopadhyay, and Sandip Chakraborty. Gestatten: Estimation of user’s attention in mobile MOOCs from eye gaze and gaze gesture tracking. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (EICS):72:1–72:32, June 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3394974.

Karizat:2021:AFT

[KDEA21] Nadia Karizat, Dan Delmonaco, Motahhare Eslami, and Nazanin Andalibi. Algorithmic folk theories and identity: How TikTok users coproduce knowledge of identity and engage in algorithmic resistance. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (CSCW2):305:1–305:44, October 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3476046.

Kirshenbaum:2021:TTT

[KDH^{+}21] Nurit Kirshenbaum, Kylie Davidson, Jesse Harden, Chris North, Dylan Kobayashi, Ryan Theriot, Roderick S. Tabalba, Michael L. Rogers, Mahdi Belcaid, Andrew T. Burks, Krishna N. Bharadwaj, Luc Renambot, Andrew E. Johnson, Lance Long, and Jason Leigh. Traces of time through space: Advantages of creating complex canvases in collaborative meetings. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(ISS): 502:1–502:20, November 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3488552.

Kaltenbrunner:2018:TPA

[KE18] Martin Kaltenbrunner and Florian Echtler. The TUIO 2.0 protocol: an abstraction framework for tangible interactive surfaces. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 2(EICS):8:1–8:35, June 2018. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3229090.

Keegan:2019:DPP

[Kee19] Brian C. Keegan. The dynamics of peer-produced political information during the 2016 U.S. Presidential campaign. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3 (CSCW):33:1–33:20, November 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3359135.

Karahalios:2017:ENC

[KFMH17] Karrie Karahalios, Geraldine Fitzpatrick, and Andrés Monroy-Hernández. Editor’s note/Chairs’ welcome. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 1 (CSCW):16:1, December 2017. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3134651.

Kosch:2018:ICA

[KFSC18] Thomas Kosch, Markus Funk, Albrecht Schmidt, and Lewis L. Chuang. Identifying cognitive assistance with mobile electroencephalography: a case study with in-situ projections for manual assembly. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 2 (EICS):11:1–11:20, June 2018. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3229093.

Kou:2020:MCA

[KG20] Yubo Kou and Xinning Gui. Mediating community-AI interaction through situated explanation: The case of AI-led moderation. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW2):102:1–102:27, October 2020. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3415173.

Kross:2021:OFB

[KG21] Sean Kross and Philip Guo. Orienting, framing, bridging, magic, and counseling: How data scientists navigate the outer loop of client collaborations in industry and academia. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (CSCW2):311:1–311:28, October 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3476052.

Kato:2020:CPI

[KIK20] Kunihiro Kato, Kaori Ikematsu, and Yoshihiro Kawahara. CA-Path: 3D-printed interfaces with conductive points in grid layout to extend capacitive touch inputs. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4(ISS): 193:1–193:17, November 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3427321.

Kiene:2019:TFU

[KJH19] Charles Kiene, Jialun Aaron Jiang, and Benjamin Mako Hill. Technological frames and user innovation: Exploring technological change in community moderation teams. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3(CSCW): 44:1–44:23, November 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3359146.

Karolus:2021:EDC

[KKE^{+}21] Jakob Karolus, Francisco Kiss, Caroline Eckerth, Nicolas Viot, Felix Bachmann, Albrecht Schmidt, and Pawel W. Wozniak. EMBody: a datacentric toolkit for EMG-Based interface prototyping and experimentation. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(EICS): 195:1–195:29, June 2021. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3457142.

Konkol:2019:CIS

[KKS19] Markus Konkol, Christian Kray, and Jan Suleiman. Creating interactive scientific publications using bindings. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3 (EICS):16:1–16:18, June 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3331158.

Kim:2020:PVC

[KLS20] Juho Kim, Siân Lindley, and Sarita Schoenebeck. PACMHCI V4 CSCW2 Oct 2020 editorial. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW2):90:1, October 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3415161.

Kim:2021:PVC

[KLS21] Juho Kim, Sian Lindley, and Sarita Schoenebeck. PACMHCI V4 CSCW3 December 2020 continued editorial. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW3):208:1, January 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3432907.

Karahalios:2018:EM

[KMHLF18] Karrie Karahalios, Andrés Monroy-Hernández, Airi Lampinen, and Geraldine Fitzpatrick. Editors’ message. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 2 (CSCW):16:1, November 2018. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3274285.

Kniestedt:2021:DDE

[KML^{+}21] Isabelle Kniestedt, Marcello A. Gómez Maureira, Iulia Lefter, Stephan Lukosch, and Frances M. Brazier. Dive deeper: Empirical analysis of game mechanics and perceived value in serious games. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CHI PLAY):236:1– 236:25, September 2021. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3474663.

Kao:2021:ETB

[KMM21] Dominic Kao, Alejandra J. Magana, and Christos Mousas. Evaluating tutorial-based instructions for controllers in virtual reality games. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CHI PLAY):234:1– 234:28, September 2021. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3474661.

Kundinger:2020:TRG

[KMR20] Thomas Kundinger, Celena Mayr, and Andreas Riener. Towards a reliable ground truth for drowsiness: a complexity analysis on the example of driver fatigue. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (EICS):78:1–78:18, June 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3394980.

Koivunen:2019:UDM

[KOOL19] Sami Koivunen, Thomas Olsson, Ekaterina Olshannikova, and Aki Lindberg. Understanding decision-making in recruitment: Opportunities and challenges for information technology. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3 (GROUP):242:1–242:22, December 2019. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3361123.

Knittel:2019:MTC

[KPW19] Megan Knittel, Shelby Pitts, and Rick Wash. “The Most Trustworthy Coin”: How ideological tensions drive trust in Bitcoin. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3(CSCW): 36:1–36:23, November 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3359138.

Karaosmanoglu:2021:LLH

[KRK^{+}21] Sukran Karaosmanoglu, Sebastian Rings, Lucie Kruse, Christian Stein, and Frank Steinicke.

Lessons learned from a human-centered design of an immersive exergame for people with dementia. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CHI PLAY):252:1– 252:27, September 2021. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3474679.

Kao:2021:ESS

[KRMM21] Dominic Kao, Rabindra Ratan, Christos Mousas, and Alejandra J. Magana. The effects of a self-similar avatar voice in educational games. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CHI PLAY):238:1– 238:28, September 2021. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3474665.

Kandappu:2021:PTP

[KSX21] Thivya Kandappu, Vigneshwaran Subbaraju, and Qianli Xu. PrivacyPrimer: Towards privacy-preserving episodic memory support for older adults. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (CSCW2):306:1–306:32, October 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3476047.

Koch:2020:IIC

[KTBL^{+}20] Janin Koch, Nicolas Taffin, Michel Beaudouin-Lafon, Markku Laine, Andrés Lucero, and Wendy E. Mackay. ImageSense: an intelligent collaborative ideation tool to support diverse human-computer partnerships. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW1):045:1–045:27, May 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3392850.

Kudo:2021:TBV

[KTF^{+}21] Yoshiki Kudo, Anthony Tang, Kazuyuki Fujita, Isamu Endo, Kazuki Takashima, and Yoshifumi Kitamura. Towards balancing VR immersion and bystander awareness. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(ISS): 484:1–484:22, November 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3486950.

Khaddam:2020:TRP

[KVDG20] Iyad Khaddam, Jean Vanderdonckt, Salah Dowaji, and Donatien Grolaux. Towards rapid prototyping of foldable graphical user interfaces with Flecto. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4(ISS): 194:1–194:33, November 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3427322.

Kruger:2021:WTG

[KWdCL^{+}21] Max Krüger, Anne Weibert, Debora de Castro Leal, Dave Randall, and Volker Wulf. “What is the Topic of the Group, Please?” on migration, care and the challenges of participation in design. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (CSCW2):309:1–309:29, October 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3476050.

Kumaran:2021:PER

[KYB21] Sneha R. Krishna Kumaran, Yue Yin, and Brian P. Bailey. Plan early, revise more: Effects of goal setting and perceived role of the feedback provider on feedback seeking behavior. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CSCW1):24:1–24:22, April 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3449098.

Li:2019:SSM

[LAHW19] Zhen Li, Michelle Annett, Ken Hinckley, and Daniel Wigdor. SMAC: a simplified model of attention and capture in multi-device desk-centric environments. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3 (EICS):2:1–2:47, June 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3300961.

Lee:2021:BTC

[LAR21] Hae-Na Lee, Vikas Ashok, and IV Ramakrishnan. Bringing things closer: Enhancing low-vision interaction experience with Office productivity applications. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(EICS):

197:1–197:18, June 2021. CODEN ???? ISSN 2573-0142 (electronic).

                                        URL https://dl.acm.
                                        org/doi/10.1145/3457144.

Lu:2019:VEI

[LAW19] Zhicong Lu, Michelle Annett, and Daniel Wigdor. Vicariously experiencing it all without going outside: a study of outdoor livestreaming in China. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3(CSCW): 25:1–25:28, November 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3359127.

Long:2021:CDA

[LBM21] Duri Long, Takeria Blunt, and Brian Magerko. Co-designing AI literacy exhibits for informal learning spaces. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (CSCW2):293:1–293:35, October 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3476034.

Logre:2018:MSV

[LDP18] Ivan Logre and Anne-Marie Déry-Pinna. MDE in support of visualization systems design: a multi-staged approach tailored for multiple roles. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 2 (EICS):14:1–14:17, June 2018. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3229096.

Liu:2021:ACD

[LE21] Jiali Liu and James Eagan. ADQDA: a crossdevice affinity diagramming tool for fluid and holistic qualitative data analysis. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(ISS): 489:1–489:19, November 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3488534.

Lampinen:2019:EM

[LGS19] Airi Lampinen, Darren Gergle, and David A. Shamma. Editors’ message. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3 (CSCW):23:1, November 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3359125.

Lobo:2020:FPE

[LHI20] María Jesús Lobo, Christophe Hurter, and Pourang Irani. Flex-ER: a platform to evaluate interaction techniques for immersive visualizations. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4(ISS): 195:1–195:20, November 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3427323.

Luo:2020:EAD

[LHPH20] Mufan Luo, Tiffany W. Hsu, Joon Sung Park, and Jeffrey T. Hancock. Emotional amplification during livestreaming: Evidence from comments during and after news events. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW1):047:1–047:19, May 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3392853.

Li:2020:SPA

[LHT^{+}20] Mingyang Li, Louis Hickman, Louis Tay, Lyle Ungar, and Sharath Chandra Guntuku. Studying politeness across cultures using English Twitter and Mandarin Weibo. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW2):119:1–119:15, October 2020. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3415190.

Lomas:2021:DSC

[LKG21] James Derek Lomas, Mihovil Karac, and Mathieu Gielen. Design space cards: Using a card deck to navigate the design space of interactive play. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CHI PLAY):227:1– 227:21, September 2021. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3474654.

Li:2021:AAC

[LKS21] Tiffany Wenting Li, Karrie Karahalios, and Hari Sundaram. “It’s all about conversation”: Challenges and concerns of faculty and students in the arts, humanities, and the social sciences about education at scale. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW3):216:1–216:37, January 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3432915.

Li:2021:HDT

[LLDH21] Tianshi Li, Elizabeth Louie, Laura Dabbish, and Jason I. Hong. How developers talk about personal data and what it means for user privacy: a case study of a developer forum on Reddit. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW3):220:1–220:28, January 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3432919.

Liaqat:2020:LPS

[LM20] Amna Liaqat and Cosmin Munteanu. Leveraging peer support for mature immigrants learning to write in informal contexts. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW2):106:1–106:24, October 2020. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3415177.

Loria:2021:SCG

[LNM21] Enrica Loria, Lennart E.

Nacke, and Annapaola Marconi. On social contagion in gamification: The power of influencers in a locationbased gameful system. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CHI PLAY):243:1– 243:20, September 2021. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3474670.

LEite:2017:MPA

[LO17] Luis LEite and Veronica Orvalho. Mani-pull-action: Hand-based digital puppetry. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 1 (EICS):2:1–2:16, June 2017. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3095804.

Lee:2020:LRU

[LOSD20] Seyeon Lee, Hyunyoung Oh, Chung-Kon Shi, and Young Yim Doh. Life review using a life metaphoric game to promote intergenerational communication. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4(CSCW2):98:1–98:21, October 2020. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3415169.

Li:2021:CCC

[LPFD21] Renee Li, Pavitthra Pandurangan, Hana Frluckaj, and Laura Dabbish. Code of conduct conversations in open source software projects on Github. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CSCW1):19:1–19:31, April 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3449093.

Lustig:2020:SMY

[LRSS20] Caitlin Lustig, Sean Rintel, Liane Scult, and Siddharth Suri. Stuck in the middle with you: The transaction costs of corporate employees hiring freelancers. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW1):037:1–037:28, May 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3392842.

Li:2020:SYS

[LUFW20] Lingyuan Li, Jirassaya Uttarapong, Guo Freeman, and Donghee Yvette Wohn. Spontaneous, yet studious: Esports commentators’ live performance and self-presentation practices. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW2):103:1–103:25, October 2020. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3415174.

Liao:2019:GIC

[LW19] Jingxian Liao and Hao-Chuan Wang. Gestures as intrinsic creativity support: Understanding the usage and function of hand gestures in computermediated group brainstorming. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3 (GROUP):243:1–243:16, December 2019. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3361124.

Lu:2020:LRC

[LX20] Alex Jiahong Lu and Xuecong Xu. “Learning for the Rise of China”:

Exploring uses and gratifications of state-owned online platform. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW1):030:1–030:25, May 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3392835.

Lee:2020:DCM

[LYH20] Yi-Chieh Lee, Naomi Yamashita, and Yun Huang. Designing a chatbot as a mediator for promoting deep selfdisclosure to a real mental health professional. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW1):031:1–031:27, May 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3392836.

Logas:2021:TBA

[LZAD21] Jacob Logas, Ruican Zhong, Stephanie Almeida, and Sauvik Das. Tensions between access and control in Makerspaces. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW3):215:1–215:33, January 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3432914.

Laine:2021:RPW

[LZS^{+}21] Markku Laine, Yu Zhang, Simo Santala, Jussi P. P. Jokinen, and Antti Oulasvirta. Responsive and personalized Web layouts with integer programming. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(EICS): 213:1–213:23, June 2021. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3461735.

Lee:2021:PPC

[LZY^{+}21] Lik Hang Lee, Yiming Zhu, Yui-Pan Yau, Pan Hui, and Susanna Pirttikangas. Press-n-Paste: Copy-and-paste operations with pressuresensitive caret navigation for miniaturized surface in mobile augmented reality. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(EICS): 199:1–199:29, June 2021. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3457146.

Micallef:2021:FGI

[MAMP21] Nicholas Micallef, Mihai Avram, Filippo Menczer, and Sameer Patil. Fakey: a game intervention to improve news literacy on social media. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (CSCW1):6:1–6:27, April 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3449080.

Marques:2020:IVC

[MAN^{+}20] Bernardo Marques, João Alves, Miguel Neves, Inês Justo, André Santos, Raquel Rainho, Rafael Maio, Dany Costa, Carlos Ferreira, Paulo Dias, and Beatriz Sousa Santos. Interaction with virtual content using augmented reality: a user study in assembly procedures. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4(ISS): 196:1–196:17, November 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3427324.

Mehmood:2019:TDC

[MAR^{+}19] Hamid Mehmood, Tallal Ahmad, Lubna Razaq, Shrirang Mare, Maryem Zafar Usmani, Richard Anderson, and Agha Ali Raza. Towards digitization of collaborative savings among low-income groups. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3(CSCW): 35:1–35:30, November 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3274304.

Mandel:2020:UCP

[MBT^{+}20] Travis Mandel, Jahnu Best, Randall H. Tanaka, Hiram Temple, Chansen Haili, Sebastian J. Carter, Kayla Schlechtinger, and Roy Szeto. Using the crowd to prevent harmful AI behavior. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4(CSCW2):97:1–97:25, October 2020. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3415168.

Magnaudet:2018:DSC

[MCC^{+}18]  Mathieu  Magnaudet,

Stéphane Chatty, Stéphane Conversy, Sébastien Leriche, Celia Picard, and Daniel Prun. Djnn/Smala: a conceptual framework and a language for interaction-oriented programming. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 2 (EICS):12:1–12:27, June 2018. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3229094.

McDonald:2020:ITA

[MCI20] Sharon McDonald, Gilbert Cockton, and Alastair Irons. The impact of thinking-aloud on usability inspection. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (EICS):88:1–88:22, June 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3397876.

Ma:2019:EAE

[MCYZ19] Haiwei Ma, Hao-Fei Cheng, Bowen Yu, and Haiyi Zhu. Effects of anonymity, ephemerality, and system routing on cost in social question asking. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3 (GROUP):238:1–238:21, December 2019. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3361119.

March:2020:WET

[MD20] Laura March and Sayamindu Dasgupta. Wikipedia edit-a-thons as sites of public pedagogy. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW2):100:1–100:26, October 2020. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3415171.

Manshaei:2019:TBM

[MDM^{+}19] Roozbeh Manshaei, Sean DeLong, Uzair Mayat, Dhrumil Patal, Matthew Kyan, and Ali Mazalek. Tangible BioNets: Multisurface and tangible interactions for exploring structural features of biological networks. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3 (EICS):14:1–14:22, June 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3331156.

Meng:2019:CDW

[MDZ19] Amanda Meng, Carl DiSalvo, and Ellen Zegura. Collaborative data work towards a caring democracy. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3(CSCW): 42:1–42:23, November 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3359144.

Massachi:2020:SSA

[MFM^{+}20] Talie Massachi, Grant Fong, Varun Mathur, Sachin R. Pendse, Gabriela Hoefer, Jessica J. Fu, Chong Wang, Nikita Ramoji, Nicole R. Nugent, Megan L. Ranney, Daniel P. Dickstein, Michael F. Armey, Ellie Pavlick, and Jeff Huang. Sochiatrist: Signals of affect in messaging data. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW2):111:1–111:25, October 2020. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3415182.

Musick:2021:GFT

[MFM21] Geoff Musick, Guo Freeman, and Nathan J. Mc-Neese. Gaming as family time: Digital game coplay in modern parentchild relationships. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CHI PLAY):251:1– 251:25, September 2021. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3474678.

Metaxa:2021:ISG

[MGG^{+}21] Danaë Metaxa, Michelle A.

Gan, Su Goh, Jeff Hancock, and James A. Landay. An image of society: Gender and racial representation and impact in image search results for occupations. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CSCW1):26:1–26:23, April 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3449100.

Ma:2019:PTR

[MH19] Ning F. Ma and Benjamin V. Hanrahan. Part-time ride-sharing: Recognizing the context in which drivers rideshare and its impact on platform use. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3 (GROUP):247:1–247:17, December 2019. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3361128.

Miller:2021:HDP

[MHG^{+}21] Josh Aaron Miller, Britton Horn, Matthew Guthrie, Jonathan Romano, Guy Geva, Celia David, Amy Robinson Sterling, and Seth Cooper. How do players and developers of citizen science games conceptualize skill chains? Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CHI PLAY):244:1– 244:29, September 2021. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3474671.

Mieczkowski:2021:AMC

[MHN^{+}21] Hannah Mieczkowski, Jeffrey T. Hancock, Mor Naaman, Malte Jung, and Jess Hohenstein. AI-mediated communication: Language use and interpersonal effects in a referential communication task. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CSCW1):17:1–17:14, April 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3449091.

Mathew:2020:HBH

[MIS^{+}20] Binny Mathew, Anurag Illendula, Punyajoy Saha, Soumya Sarkar, Pawan Goyal, and Animesh Mukherjee. Hate begets hate: a temporal study of hate speech. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4(CSCW2):92:1–92:24, October 2020. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3415163.

Moran-Ledesma:2021:UDG

[MLSH21] Marco Moran-Ledesma, Oliver Schneider, and Mark Hancock. Userdefined gestures with physical props in virtual reality. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(ISS): 488:1–488:23, November 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3486954.

McDonald:2019:IMC

[MM19]  Samantha  McDonald

and Melissa Mazmanian. Information materialities of citizen communication in the U.S. Congress. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3(CSCW): 47:1–47:20, November 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3359149.

Machado:2021:CBW

[MMdSP21] Leticia S. Machado, Ricardo Rodrigo M. Melo, Cleidson R. B. de Souza, and Rafael Prikladnicki. Collaborative behavior and winning challenges in competitive software crowdsourcing. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (GROUP):220:1–220:25, July 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3463932.

Mogavi:2021:CSE

[MMH21] Reza Hadi Mogavi, Xiaojuan Ma, and Pan Hui. Characterizing student engagement moods for dropout prediction in question pool Websites. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CSCW1):12:1–12:22, April 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3449086.

Morisaki:2021:UDC

[MMM^{+}21] Tao Morisaki, Ryoma Mori, Ryosuke Mori, Kohki Serizawa, Yasutoshi Makino, Yuta Itoh, Yuji Yamakawa, and Hiroyuki Shinoda. Ultrasound-driven curveball in table tennis: Human activity support via noncontact remote object manipulation. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(ISS): 503:1–503:20, November 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3488548.

Moldovan:2020:OUI

[MNP^{+}20] Alex Moldovan, Vlad Nicula, Ionut Pasca, Mihai Popa, Jaya Krishna Namburu, Anamaria Oros, and Paul Brie. OpenUIDL, a user interface description language for runtime omnichannel user interfaces. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (EICS):86:1–86:52, June 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3397874.

Moller:2021:CWT

[MNS^{+}21] Naja Holten Møller, Gina Neff, Jakob Grue Simonsen, Jonas Christoffer Villumsen, and Pernille Bjørn. Can workplace tracking ever empower? Collective sensemaking for the responsible use of sensor data at work. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (GROUP):219:1–219:21, July 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3463931.

Matsuda:2021:SSR

[MON^{+}21] Akira Matsuda, Toru Okuzono, Hiromi Nakamura, Hideaki Kuzuoka, and Jun Rekimoto. A surgical scene replay system for learning gastroenterological endoscopic surgery skill by multiple synchronizedvideo and gaze representation. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(EICS): 204:1–204:22, June 2021. CODEN ???? ISSN 2573-0142 (electronic).

                                        URL https://dl.acm.
                                        org/doi/10.1145/3461726.

Morgado:2019:ITA

[MP19] Inês Coimbra Morgado and Ana C. R. Paiva. The iMPAcT tool for Android testing. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3(EICS):4:1–4:23, June 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3300963.

Martinie:2019:ADT

[MPB^{+}19] Célia Martinie, Philippe Palanque, Elodie Bouzekri, Andy Cockburn, Alexandre Canny, and Eric Barboni. Analysing and demonstrating toolsupported customizable task notations. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3 (EICS):12:1–12:26, June 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3331154.

Mendes:2020:CTB

[MRGN20] Daniel Mendes, Sofia Reis, João Guerreiro, and Hugo Nicolau. Collaborative tabletops for blind people: The effect of auditory design on workspace awareness. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4(ISS): 197:1–197:19, November 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3427325.

Mostaeen:2019:DRT

[MRRS19] Golam Mostaeen, Banani Roy, Chanchal Roy, and Kevin Schneider. Designing for real-time groupware systems to support complex scientific data analysis. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3 (EICS):9:1–9:28, June 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3331151.

Miceli:2020:BSI

[MSY20] Milagros Miceli, Martin Schuessler, and Tianling Yang. Between subjectivity and imposition: Power dynamics in data annotation for computer vision. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW2):115:1–115:25, October 2020. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3415186.

Marathe:2021:SRE

[MT21] Megh Marathe and Kentaro Toyama. The situated, relational, and evolving nature of epilepsy diagnosis. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW3):217:1–217:18, January 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3432916.

Maquil:2017:CRI

[MTA^{+}17] Valerie Maquil, Eric Tobias, Dimitra Anastasiou, Hélène Mayer, and Thibaud Latour. COPSE: Rapidly instantiating problem solving activities based on tangible tabletop interfaces. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 1(EICS):6:1–6:16, June 2017. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3095808.

Magrofuoco:2019:GCP

[MV19] Nathan Magrofuoco and Jean Vanderdonckt. Gelicit: a cloud platform for distributed gesture elicitation studies. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3 (EICS):6:1–6:41, June 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3331148.

MacArthur:2019:MQI

[MWH19] Cayley MacArthur, Caroline Wong, and Mark Hancock. Makers and quilters: Investigating opportunities for improving gender-imbalanced maker groups. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3 (CSCW):29:1–29:24, November 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3359131.

Mao:2019:HDS

[MWM^{+}19] Yaoli Mao, Dakuo Wang, Michael Muller, Kush R. Varshney, Ioana Baldini, Casey Dugan, and Aleksandra Mojsilović. How data scientists work together with domain experts in scientific collaborations: To find the right answer or to ask the right question? Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3 (GROUP):237:1–237:23, December 2019. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3361118.

Marathe:2021:TVT

[MYS^{+}21] Megh Marathe, Yoonseon Yi, Chia-Hsuan Su, Ting-Wei Chang, and Gabriela Marcu. Tedious versus taxing: The nature of work in a behavioral health context. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (CSCW2):302:1–302:24, October 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3476043.

Nejati:2020:WFG

[NB20] Javad Nejati and Aruna Balasubramanian. WProfX: a fine-grained visualization tool for Web page loads. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (EICS):73:1–73:22, June 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3394975.

Nolte:2020:HSN

[NHH20] Alexander Nolte, Linda Bailey Hayden, and James D. Herbsleb. How to support newcomers in scientific hackathons — an action research study on expert mentoring. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW1):025:1–025:23, May 2020. CODEN ????

ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3392830.

Ntouros:2021:DSE

[NKV21] Vasilis Ntouros, Hara Kouki, and Vasilis Vlachokyriakos. Designing sharing economy platforms through a ‘solidarity HCI’ lens. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CSCW1):23:1–23:25, April 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3449097.

Norambuena:2021:NMA

[NM21] Brian Felipe Keith Norambuena and Tanushree Mitra. Narrative maps: an algorithmic approach to represent and extract information narratives. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW3):228:1–228:33, January 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3432927.

Niyazov:2021:DDP

[NMBS21] Aziz Niyazov, Nicolas Mellado, Loic Barthe, and Marcos Serrano. Dynamic decals: Pervasive freeform interfaces using constrained deformable graphical elements. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(ISS): 493:1–493:27, November 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3488538.

Norton:2019:IGS

[NPT19] Juliet Norton, Birgit Penzenstadler, and Bill Tomlinson. Implications of grassroots sustainable agriculture community values on the design of information systems. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3(CSCW): 34:1–34:22, November 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3359136.

Nelimarkka:2021:ENV

[NSJ21] Matti Nelimarkka, Antti Salovaara, and Giulio Jacucci. Exploring norm violations and norm management in collocated synchronous communication: Dualchannel communication as hybrid space with spillover and asymmetries. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(GROUP): 221:1–221:24, July 2021. CODEN ???? ISSN 2573-0142 (electronic).

                                        URL https://dl.acm.
                                        org/doi/10.1145/3463933.

Ouverson:2021:CFC

[OG21] Kaitlyn M. Ouverson and Stephen B. Gilbert. A composite framework of co-located asymmetric virtual reality. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (CSCW1):5:1–5:20, April 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3449079.

Oprea:2020:ESV

[OM20] Silviu Vlad Oprea and Walid Magdy. The effect of sociocultural variables on sarcasm communication online. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW1):029:1–029:22, May 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3392834.

Osinski:2019:TSK

[OR19] Meike Osinski and Nikol Rummel. Towards successful knowledge integration in online collaboration: an experiment on the role of metaknowledge. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3(CSCW): 31:1–31:17, November 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3359133.

Ousmer:2020:RTM

[OSM^{+}20] Mehdi Ousmer, Arthur Sluÿters, Nathan Magrofuoco, Paolo Roselli, and Jean Vanderdonckt. Recognizing 3D trajectories as 2D multi-stroke gestures. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4(ISS): 198:1–198:21, November 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3427326.

Oppenlaender:2020:CSW

[OTH20]  Jonas  Oppenlaender,

Thanassis Tiropanis, and Simo Hosio. CrowdUI: Supporting Web design with the crowd. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (EICS):76:1–76:28, June 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3394978.

Panicker:2020:CRC

[PBC20] Aswati Panicker, Kavya Basu, and Chia-Fang Chung. Changing roles and contexts: Symbolic interactionism in the sharing of food and eating practices between remote, intergenerational family members. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW1):043:1–043:19, May 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3392848.

Pruszko:2021:MHS

[PCL^{+}21] Laura Pruszko, Céline Coutrix, Yann Laurillau, Benoît Piranda, and Julien Bourgeois. Molecular HCI: Structuring the cross-disciplinary space of modular shapechanging user interfaces. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(EICS): 211:1–211:33, June 2021. CODEN ???? ISSN 2573-0142 (electronic).

                                        URL https://dl.acm.
                                        org/doi/10.1145/3461733.

Parvin:2018:RTA

[PCMP18] Parvaneh Parvin, Stefano Chessa, Marco Manca, and Fabio Paterno’. Real-time anomaly detection in elderly behavior with the support of task models. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 2 (EICS):15:1–15:18, June 2018. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3229097.

Park:2019:GMU

[PKB^{+}19] Chang Min Park, Taeyeon Ki, Ali J. Ben Ali, Nikhil Sunil Pawar, Karthik Dantu, Steven Y. Ko, and Lukasz Ziarek. Gesto: Mapping UI events to gestures and voice commands. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3(EICS):5:1–5:22, June 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3300964.

Pena:2021:PSW

[PKC^{+}21] Belén Barros Pena, Bailey Kursar, Rachel E. Clarke, Katie Alpin, Merlyn Holkar, and John Vines. “Pick Someone Who Can Kick Your Ass”

— moneywork in financial third party access. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW3):218:1–218:28, January 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3432917.

Phadke:2021:ESF

[PM21] Shruti Phadke and Tanushree Mitra. Educators, solicitors, flamers, motivators, sympathizers: Characterizing roles in online extremist movements. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (CSCW2):310:1–310:35, October 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3476051.

Paananen:2021:IHS

[POG^{+}21] Ville Paananen, Jonas Oppenlaender, Jorge Goncalves, Danula Hettiachchi, and Simo Hosio. Investigating human scale spatial experience. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(ISS): 496:1–496:18, November 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3488541.

Prilla:2019:SWW

[Pri19] Michael Prilla. “I simply watched where she was looking at”: Coordination in short-term synchronous cooperative mixed reality. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3 (GROUP):246:1–246:21, December 2019. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3361127.

Phadke:2021:WMP

[PSM21] Shruti Phadke, Mattia Samory, and Tanushree Mitra. What makes people join conspiracy communities?: Role of social factors in conspiracy engagement. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW3):223:1–223:30, January 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3432922.

Qiu:2020:ECS

[QGB20] Sihang Qiu, Ujwal Gadiraju, and Alessandro Bozzon. Estimating conversational styles in conversational microtask crowdsourcing. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW1):032:1–032:23, May 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3392837.

Ren:2021:SST

[RA21a] Yuan Ren and Ahmed Sabbir Arif. Stepper, swipe, tilt, force: Comparative evaluation of four number pickers for smartwatches. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(ISS): 500:1–500:21, November 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3488545.

Roemmich:2021:DSC

[RA21b] Kat Roemmich and Nazanin Andalibi. Data subjects’ conceptualizations of and attitudes toward automatic emotion recognition-enabled wellbeing interventions on social media. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (CSCW2):308:1–308:34, October 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3476049.

Roth:2020:ENM

[RAZ^{+}20] Rebecca Roth, Jaclyn Abraham, Heidi Zinzow, Pamela Wisniewski, Amro Khasawneh, and Kapil Chalil Madathil. Evaluating news media reports on the ’blue whale challenge’ for adherence to suicide prevention safe messaging guidelines. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW1):026:1–026:27, May 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3392831.

Richards:2021:SUC

[RCM21] Olivia K. Richards, Adrian Choi, and Gabriela Marcu. Shared understanding in care coordination for children’s behavioral health. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CSCW1):21:1–21:25, April 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3449095.

Roels:2018:IUF

[RDS18] Reinout Roels, Arno De Witte, and Beat Signer. INFEX: a unifying framework for crossdevice information exploration and exchange. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 2(EICS):2:1–2:26, June 2018. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3179427.

Roohi:2021:PGD

[RGR^{+}21] Shaghayegh Roohi, Christian Guckelsberger, Asko Relas, Henri Heiskanen, Jari Takatalo, and Perttu Hämäläinen. Predicting game difficulty and engagement using AI players. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CHI PLAY):231:1– 231:17, September 2021. CODEN ???? ISSN 2573-0142 (electronic).

                                        URL https://dl.acm.
                                        org/doi/10.1145/3474658.

Raffaillac:2019:PPI

[RH19] Thibault Raffaillac and Stéphane Huot. Polyphony: Programming interfaces and interactions with the entity-component-system model. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3(EICS):8:1–8:22, June 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3331150.

Rankin:2020:STB

[RI20] Yolanda A. Rankin and India Irish. A seat at the table: Black feminist thought as a critical framework for inclusive game design. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW2):117:1–117:26, October 2020. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3415188.

Ribes:2019:HLW

[Rib19] David Ribes. How I learned what a domain was. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3(CSCW): 38:1–38:12, November 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3359140.

Ravi:2021:PSR

[RIK21] Prerna Ravi, Azra Ismail, and Neha Kumar.

The pandemic shift to remote learning under resource constraints. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (CSCW2):314:1–314:28, October 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3476055.

Ribeiro:2021:DPM

[RJZ^{+}21] Manoel Horta Ribeiro, Shagun Jhaver, Savvas Zannettou, Jeremy Blackburn, Gianluca Stringhini, Emiliano De Cristofaro, and Robert West. Do platform migrations compromise content moderation? Evidence from r/The Donald and r/Incels. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (CSCW2):316:1–316:24, October 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3476057.

Rashidi:2020:ETC

[RKNES20] Yasmeen Rashidi, Apu Kapadia, Christena Nippert-Eng, and Norman Makoto Su. “It’s easier than causing confrontation”: Sanctioning strategies to maintain social norms and privacy on social media. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW1):023:1–023:25, May 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3392827.

Rekik:2021:MCT

[RLGG21] Yosra Rekik, Edward Lank, Adnane Guettaf, and Laurent Grisoni. Multi-channel tactile feedback based on user finger speed. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(ISS): 504:1–504:17, November 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3488549.

Reibert:2020:MIP

[RRF20] Joshua Reibert, Patrick Riehmann, and Bernd Froehlich. Multitouch interaction with parallel coordinates on large vertical displays. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (ISS):199:1–199:22, November 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3427327.

Roels:2019:CFC

[RS19] Reinout Roels and Beat Signer. A conceptual framework and content model for next generation presentation solutions. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3(EICS): 7:1–7:22, June 2019. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3331149.

Reetz:2021:NVS

[RVBC^{+}21] Adrian Reetz, Deltcho Valtchanov, Michael Barnett-Cowan, Mark Hancock, and James R. Wallace. Nature vs. stress: Investigating the use of biophilia in non-violent exploration games to reduce stress. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CHI PLAY):247:1– 247:13, September 2021. CODEN ???? ISSN 2573-0142 (electronic).

                                        URL https://dl.acm.
                                        org/doi/10.1145/3474674.

Rakova:2021:WRA

[RYCC21] Bogdana Rakova, Jingying Yang, Henriette Cramer, and Rumman Chowdhury. Where responsible AI meets reality: Practitioner perspectives on enablers for shifting organizational practices. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (CSCW1):7:1–7:23, April 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3449081.

Rabaan:2021:DMS

[RYD21] Hawra Rabaan, Alyson L.

Young, and Lynn Dombrowski. Daughters of men: Saudi Women’s sociotechnical agency practices in addressing domestic abuse. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW3):224:1–224:31, January 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3432923.

Soubutts:2021:APT

[SAE^{+}21] Ewan Soubutts, Amid Ayobi, Rachel Eardley, Kirsten Cater, and Aisling Ann O’Kane. Aging in place together: The journey towards adoption and acceptance of stairlifts in multiresident homes. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (CSCW2):320:1–320:26, October 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3476061.

Salem:2017:UIO

[Sal17] Paulo Salem. User interface optimization using genetic programming with an application to landing pages. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 1 (EICS):13:1–13:17, June 2017. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3099583.

Sultana:2021:SCT

[SAR21] Sharifa Sultana, Syed Ishtiaque Ahmed, and Jeffrey M. Rzeszotarski. Seeing in context: Traditional visual communication practices in rural Bangladesh. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW3):214:1–214:31, January 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3432913.

Segura:2017:HIV

[SB17] Vinícius Segura and Simone D. J. Barbosa. HistoryViewer: Instrumenting a visual analytics application to support revisiting a session of interactive data analysis. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 1 (EICS):11:1–11:18, June 2017. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3095813.

Samuelsson:2020:ESE

[SB20a] Sigurdur Gauti Samuelsson and Matthias Book. Eliciting sketched expressions of command intentions in an IDE. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (ISS):200:1–200:25, November 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3427328.

Skeba:2020:IFL

[SB20b] Patrick Skeba and Eric P. S. Baumer. Informational friction as a lens for studying algorithmic aspects of privacy. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW2):101:1–101:22, October 2020. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3415172.

Sandbhor:2021:LWR

[SBAK21] Prasad Sandbhor, Priti Bangal, Deepti Aggarwal, and Rohit Ashok Khot. Life on wings: Relating to a bird’s life in a city through a board game. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CHI PLAY):232:1– 232:28, September 2021. CODEN ???? ISSN 2573-0142 (electronic).

                                        URL https://dl.acm.
                                        org/doi/10.1145/3474659.

Shivakumar:2020:RTI

[SBPP20] Ashutosh Shivakumar, Aishwarya Bositty, Nia S. Peters, and Yong Pei. Real-time interruption management system for efficient distributed collaboration in multitasking environments. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW1):039:1–039:23, May 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3392844.

Serpi:2018:WFF

[SCMS18] Matteo Serpi, Alessandro Carcangiu, Alessio Murru, and Lucio Davide Spano. Web5VR: a flexible framework for integrating virtual reality input and output devices on the Web. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 2(EICS):4:1–4:19, June 2018. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3179429.

Saha:2021:UAS

[SCP^{+}21] Manaswi Saha, Devanshi Chauhan, Siddhant Patil, Rachel Kangas, Jeffrey Heer, and Jon E. Froehlich. Urban accessibility as a sociopolitical problem: a multi-stakeholder analysis. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4(CSCW3): 209:1–209:26, January 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3432908.

Speicher:2018:MAH

[SCY^{+}18]  Maximilian  Speicher,

Jingchen Cao, Ao Yu, Haihua Zhang, and Michael Nebeling. 360Anywhere: Mobile ad-hoc collaboration in any environment using 360 video and augmented reality. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 2(EICS):9:1–9:20, June 2018. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3229091.

Schneider:2021:MPT

[SDF^{+}21] Nathan Schneider, Primavera De Filippi, Seth Frey, Joshua Z. Tan, and Amy X. Zhang. Modular politics: Toward a governance layer for online communities. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CSCW1):16:1–16:26, April 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3449090.

Shiripour:2021:GBG

[SDO21] Morteza Shiripour, Niraj Ramesh Dayama, and Antti Oulasvirta. Gridbased genetic operators for graphical layout generation. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(EICS): 208:1–208:30, June 2021. CODEN ???? ISSN 2573-0142 (electronic).

                                        URL https://dl.acm.
                                        org/doi/10.1145/3461730.

Satriadi:2020:MAM

[SEC^{+}20] Kadek Ananta Satriadi, Barrett Ens, Maxime Cordeil, Tobias Czauderna, and Bernhard Jenny. Maps around me: 3D multiview layouts in immersive spaces. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4(ISS): 201:1–201:20, November 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3427329.

Seering:2020:RSM

[See20] Joseph Seering. Reconsidering self-moderation: the role of research in supporting communitybased models for online content moderation. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW2):107:1–107:28, October 2020. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3415178.

Simons:2020:HHU

[SGF20] Rachel N. Simons, Danna Gurari, and Kenneth R. Fleischmann. “I Hope This Is Helpful”: Understanding crowdworkers’ challenges and motivations for an image description task. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW2):105:1–105:26, October 2020. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3415176.

Scheuerman:2021:DDP

[SHD21] Morgan Klaus Scheuerman, Alex Hanna, and Emily Denton. Do datasets have politics? disciplinary values in computer vision dataset development. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (CSCW2):317:1–317:37, October 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3476058.

Speicher:2018:XAC

[SHY^{+}18]  Maximilian  Speicher,

Brian D. Hall, Ao Yu, Bowen Zhang, Haihua Zhang, Janet Nebeling, and Michael Nebeling. XD-AR: Challenges and opportunities in crossdevice augmented reality application development. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 2(EICS):7:1–7:24, June 2018. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3229089.

Singh:2021:SLI

[SJ21] Ranjit Singh and Steven Jackson. Seeing like an infrastructure: Lowresolution citizens and the Aadhaar Identification Project. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (CSCW2):315:1–315:26, October 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3476056.

Sheng:2020:VSI

[SK20] Jeff T. Sheng and Sanjay R. Kairam. From virtual strangers to IRL friends: Relationship development in livestreaming communities on twitch. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4(CSCW2): 94:1–94:34, October 2020. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3415165.

Saveski:2021:SCC

[SKV^{+}21] Martin Saveski, Farshad Kooti, Sylvia Morelli Vitousek, Carlos Diuk, Bryce Bartlett, and Lada A. Adamic. Social catalysts: Characterizing people who spark conversations among others. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (CSCW2):292:1–292:20, October 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3476023.

Soden:2021:PAN

[SLD21] Robert Soden, Nicolas James LaLone, and Dharma Dailey. A patent application for NEXTGEN flood early warning system. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (GROUP):216:1–216:16, July 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3463928.

Speicher:2021:DSY

[SLN21]  Maximilian  Speicher,

Katy Lewis, and Michael Nebeling. Designers, the stage is yours! mediumfidelity prototyping of augmented & virtual reality interfaces with 360theater. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(EICS): 205:1–205:25, June 2021. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3461727.

Smiley:2021:MAM

[SLT^{+}21] Jim Smiley, Benjamin Lee, Siddhant Tandon, Maxime Cordeil, Lonni Besançon, Jarrod Knibbe, Bernhard Jenny, and Tim Dwyer. The MADE-Axis: a modular actuated device to embody the axis of a data dimension. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(ISS): 501:1–501:23, November 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3488546.

Stumpf:2018:WL

[SN18] Simone Stumpf and Jeffrey Nichols. Welcome letter. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 2(EICS): 1:1, June 2018. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3179426.

Swofford:2020:ISA

[SPT^{+}20] Mason Swofford, John Peruzzi, Nathan Tsoi, Sydney Thompson, Roberto Martín-Martín, Silvio Savarese, and Marynel Vázquez. Improving social awareness through DANTE: Deep affinity network for clustering conversational interactants. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW1):020:1–020:23, May 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3392824.

Sher:2019:SCH

[SS19] Stephen Tsung-Han Sher and Norman Makoto Su. Speedrunning for charity: How donations gather around a live streamed couch. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3 (CSCW):48:1–48:26, November 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3359150.

Stewart:2021:NPR

[SS21] Leo G. Stewart and Emma S. Spiro. Nobody puts Redditor in a binary: Digital demography, collective identities, and gender in a Subreddit network. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (CSCW1):8:1–8:31, April 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3449082.

Saxena:2020:MGT

[SSGB20] Devansh Saxena, Patrick Skeba, Shion Guha, and Eric P. S. Baumer. Methods for generating typologies of non/use. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW1):027:1–027:26, May 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3392832.

Skeba:2021:WCS

[SSGB21] Patrick Skeba, Devansh Saxena, Shion Guha, and Eric P. S. Baumer. Who has a choice?: Surveybased predictors of volitionality in Facebook use and non-use. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (GROUP):223:1–223:25, July 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3463935.

Schoenebeck:2021:YTS

[SSH^{+}21] Sarita Schoenebeck, Carol F. Scott, Emma Grace Hurley, Tammy Chang, and Ellen Selkie. Youth trust in social media companies and expectations of justice: Accountability and repair after online harassment. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (CSCW1):2:1–2:18, April 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3449076.

Sun:2020:RDB

[SSNC20] David Sun, Chengzheng Sun, Agustina Ng, and Weiwei Cai. Real differences between OT and CRDT in correctness and complexity for consistency maintenance in co-editors. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW1):021:1–021:30, May 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3392825.

Swaminathan:2019:LUM

[STM^{+}19] Saiganesh Swaminathan, Indrani Medhi Thies, Devansh Mehta, Edward Cutrell, Amit Sharma, and William Thies. Learn2Earn: Using mobile airtime incentives to bolster public awareness campaigns. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3(CSCW): 49:1–49:20, November 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3359151.

Sluyters:2021:ESG

[SVV21] Arthur Sluÿters, Jean Vanderdonckt, and Radu-Daniel Vatavu. Engineering slidable graphical user interfaces with slime. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(EICS): 200:1–200:29, June 2021. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3457147.

Schipor:2019:STS

[SVW19] Ovidiu-Andrei Schipor, Radu-Daniel Vatavu, and Wenjun Wu. SAPI-ENS: Towards software architecture to support peripheral interaction in smart environments. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3 (EICS):11:1–11:24, June 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3331153.

Suh:2020:PJP

[SWF^{+}20] Jina Suh, Spencer Williams, Jesse R. Fann, James Fogarty, Amy M. Bauer, and Gary Hsieh. Parallel journeys of patients with cancer and depression: Challenges and opportunities for technology-enabled collaborative care. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW1):038:1–038:36, May 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3392843.

Silva:2020:ECB

[SWT20] Thiago Rocha Silva, Marco Winckler, and Hallvard Trætteberg. Ensuring the consistency between user requirements and task models: a behavior-based automated approach. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (EICS):77:1–77:32, June 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3394979.

Tang:2021:UTE

[Tan21] John Tang. Understanding the telework experience of people with disabilities. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CSCW1):30:1–30:27, April 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3449104.

Taugerbeck:2019:DPP

[TAS^{+}19] Sebastian Taugerbeck, Michael Ahmadi, Marén Schorch, David Unbehaun, Konstantin Aal, and Volker Wulf. Digital participation in prison — a public discourse analysis of the use of ICT by inmates. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3 (GROUP):233:1–233:26, December 2019. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3361114.

Taylor:2021:MMD

[TBDB21] Lauren C. Taylor, Kelsie Belan, Munmun De Choudhury, and Eric P. S. Baumer. Misfires, missed data, misaligned treatment: Disconnects in collaborative treatment of eating disorders. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CSCW1):31:1–31:28, April 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3449105.

Turner:2020:MBT

[TBR20] Jessica Turner, Judy Bowen, and Steve Reeves. Model-based testing of interactive systems using interaction sequences. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (EICS):85:1–85:37, June 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3397873.

Toch:2020:CYT

[TCH20] Eran Toch, Hadas Chassidim, and Tali Hatuka. Can you turn it off?: The spatial and social context of mobile disturbance. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4(CSCW2):91:1–91:18, October 2020. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3415162.

Theis:2019:BPN

[TD19] Julian Theis and Houshang Darabi. Behavioral Petri net mining and automated analysis for human-computer interaction recommendations in multi-application environments. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3 (EICS):13:1–13:16, June 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3331155.

Terrier:2017:CLB

[TDC17] Lénaïc Terrier, Alexandre Demeure, and Sybille Caffiau. CCBL: a language for better supporting context centered programming in the smart home. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 1 (EICS):14:1–14:18, June 2017. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3099584.

Tomlein:2018:ARS

[TG18] Matús Tomlein and Kaj Grønbæk. Augmented reality supported modeling of industrial systems to infer software configuration. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 2 (EICS):5:1–5:17, June 2018. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3229087.

Thevin:2021:IAE

[TRO^{+}21] Lauren Thevin, Nicolas Rodier, Bernard Oriola, Martin Hachet, Christophe Jouffrais, and Anke M. Brock. Inclusive adaptation of existing board games for gamers with and without visual impairments using a spatial augmented reality framework for touch detection and audio feedback. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(ISS): 505:1–505:33, November 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3488550.

To:2020:TJD

[TSHK20] Alexandra To, Wenxia Sweeney, Jessica Hammer, and Geoff Kaufman. “They Just Don’t Get It”: Towards social technologies for coping with interpersonal racism. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW1):024:1–024:29, May 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3392828.

Toxtli:2021:QIL

[TSS21] Carlos Toxtli, Siddharth Suri, and Saiph Savage. Quantifying the invisible labor in crowd work. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (CSCW2):319:1–319:26, October 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3476060.

Trullemans:2017:CMT

[TVS17] Sandra Trullemans, Lars Van Holsbeeke, and Beat Signer. The context modelling toolkit: a unified multi-layered context modelling approach. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 1(EICS):8:1–8:16, June 2017. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3095810.

Uckun:2020:BAB

[UAAR20] Utku Uckun, Ali Selman Aydin, Vikas Ashok, and IV Ramakrishnan. Breaking the accessibility barrier in nonvisual interaction with PDF forms. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (EICS):80:1–80:16, June 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3397868.

Usuba:2021:MMT

[UYM21] Hiroki Usuba, Shota Yamanaka, and Homei Miyashita. Modeling movement times and success rates for acquisition of one-dimensional targets with uncertain touchable sizes. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (ISS):487:1–487:15, November 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3486953.

vanBerkel:2019:CPF

[vBGH^{+}19] Niels van Berkel, Jorge Goncalves, Danula Hettiachchi, Senuri Wijenayake, Ryan M. Kelly, and Vassilis Kostakos. Crowdsourcing perceptions of fair predictors for machine learning: a recidivism case study. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3(CSCW): 28:1–28:21, November 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3359130.

Vincent:2021:DII

[VH21] Nicholas Vincent and Brent Hecht. A deeper investigation of the importance of Wikipedia links to search engine results. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (CSCW1):4:1–4:15, April 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3449078.

Vayanou:2021:PSC

[VKI21] Maria Vayanou, Akrivi Katifori, and Yannis Ioannidis. Perspective sharing in culture group games: Passing around the social mediator role. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CHI PLAY):249:1– 249:24, September 2021. CODEN ???? ISSN 2573-0142 (electronic).

                                        URL https://dl.acm.
                                        org/doi/10.1145/3474676.

Vallett:2020:TAS

[VMD^{+}20] Richard Vallett, Denisa Qori McDonald, Genevieve Dion, Youngmoo Kim, and Ali Shokoufandeh. Toward accurate sensing with knitted fabric: Applications and technical considerations. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (EICS):79:1–79:26, June 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3394981.

Vaghela:2021:BCH

[VMP21] Palashi Vaghela, Ramaravind K. Mothilal, and Joyojeet Pal. Birds of a caste — how caste hierarchies manifest in retweet behavior of Indian politicians. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW3):212:1–212:24, January 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3432911.

Vanderdonckt:2019:MDC

[VN19] Jean Vanderdonckt and Thanh-Diane Nguyen.

MoCaDiX: Designing cross-device user interfaces of an information system based on its class diagram. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3 (EICS):17:1–17:40, June 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3331159.

Vaccaro:2021:CCM

[VXHK21] Kristen Vaccaro, Ziang Xiao, Kevin Hamilton, and Karrie Karahalios. Contestability for content moderation. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (CSCW2):318:1–318:28, October 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3476059.

Vyas:2019:AWC

[Vya19] Dhaval Vyas. Altruism and wellbeing as care work in a craft-based maker culture. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3 (GROUP):239:1–239:12, December 2019. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3361120.

Vanderdonckt:2019:ALB

[VZV19] Jean Vanderdonckt, Mathieu Zen, and Radu-Daniel Vatavu. AB4Web: an on-line A/B tester for comparing user interface design alternatives. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3 (EICS):18:1–18:28, June 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3331160.

Welsford-Ackroyd:2021:SVE

[WACdA^{+}21] Finn Welsford-Ackroyd, Andrew Chalmers, Rafael Kuffner dos Anjos, Daniel Medeiros, Hyejin Kim, and Taehyun Rhee. Spectator View: Enabling asymmetric interaction between HMD wearers and spectators with a large display. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(ISS): 485:1–485:17, November 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3486951.

Wu:2021:BFN

[WB21] Y. Wayne Wu and Brian P. Bailey. Better feedback from nicer people: Narrative empathy and ingroup framing improve feedback exchange. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW3):236:1–236:20, January 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3432935.

Wozniak:2020:SAP

[WDP^{+}20] Mikołaj P. Wożniak, Julia Dominiak, Michał Pieprzowski, Piotr Ładoński, Krzysztof Grudzień, Lars Lischke, Andrzej Romanowski, and Paweł W. Wożniak. Subtletee: Augmenting posture awareness for beginner golfers. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (ISS):204:1–204:24, November 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3427332.

Wyche:2020:EPR

[WGG20] Susan Wyche, April Greenwood, and Brian Samuel Geyer. Exploring photography in rural Kenyan households: Considering “relational objects” in CSCW and HCI. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW1):046:1–046:26, May 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3392852.

Weninger:2020:EIM

[WGGS20] Markus Weninger, Paul Grünbacher, Elias Gander, and Andreas Schörgenhumer. Evaluating an interactive memory analysis tool: Findings from a cognitive walkthrough and a user study. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (EICS):75:1–75:37, June 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3394977.

Weninger:2021:GEM

[WGM21] Markus Weninger, Elias Gander, and Hanspeter Mössenböck. Guided exploration: a method for guiding novice users in interactive memory monitoring tools. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(EICS): 209:1–209:34, June 2021. CODEN ???? ISSN 2573-0142 (electronic).

                                        URL https://dl.acm.
                                        org/doi/10.1145/3461731.

Wambecke:2021:MEG

[WGN^{+}21] Jérémy Wambecke, Alix Goguey, Laurence Nigay, Lauren Dargent, Daniel Hauret, Stéphanie Lafon, and Jean-Samuel Louis de Visme. M[eye]cro: Eye-gaze+Microgestures for multitasking and interruptions. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(EICS): 210:1–210:22, June 2021. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3461732.

Wilson:2020:CSC

[Wil20] Andrew D. Wilson. Combating the spread of coronavirus by modeling fomites with depth cameras. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4(ISS): 203:1–203:13, November 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3427331.

Wang:2019:HDS

[WMBO19] April Yi Wang, Anant Mittal, Christopher Brooks, and Steve Oney. How data scientists use computational notebooks for real-time collaboration. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3(CSCW): 39:1–39:30, November 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3359141.

Williams:2019:PWL

[WMM^{+}19] Alex C. Williams, Gloria Mark, Kristy Milland, Edward Lank, and Edith Law. The perpetual work life of crowdworkers: How tooling practices increase fragmentation in crowdwork. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3(CSCW): 24:1–24:28, November 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3359126.

Williams:2020:UGS

[WO20] Adam S. Williams and Francisco R. Ortega. Understanding gesture and speech multimodal interactions for manipulation tasks in augmented reality using unconstrained elicitation. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4(ISS): 202:1–202:21, November 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3427330.

Wessel:2021:DDM

[WWSG21] Mairieli Wessel, Igor Wiese, Igor Steinmacher, and Marco Aurelio Gerosa. Don’t disturb me: Challenges of interacting with software bots on open source software projects. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (CSCW2):301:1–301:21, October 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3476042.

Wang:2021:CTB

[WWT^{+}21] Liuping Wang, Dakuo Wang, Feng Tian, Zhenhui Peng, Xiangmin Fan, Zhan Zhang, Mo Yu, Xiaojuan Ma, and Hongan Wang. CASS: Towards building a socialsupport chatbot for online health community. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (CSCW1):9:1–9:31, April 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3449083.

Wang:2021:TEQ

[WYY^{+}21] Liang Wang, Zhiwen Yu, Dingqi Yang, Tian Wang, En Wang, Bin Guo, and Daqing Zhang. Task execution quality maximization for mobile crowdsourcing in geosocial networks. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (CSCW2):312:1–312:29, October 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3476053.

Xiao:2017:SRC

[XHH17] Robert Xiao, Scott Hudson, and Chris Harrison. Supporting responsive cohabitation between virtual interfaces and physical objects on everyday surfaces. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 1 (EICS):12:1–12:17, June 2017. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3095814.

Xia:2020:EAC

[XZL^{+}20] Yan Xia, Haiyi Zhu, Tun Lu, Peng Zhang, and Ning Gu. Exploring antecedents and consequences of toxicity in online discussions: a case study on Reddit. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW2):108:1–108:23, October 2020. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3415179.

Yu:2021:PFY

[YGHR21] Junnan Yu, Julisa Granados, Ronni Hayden, and Ricarose Roque. Parental facilitation of young children’s technologybased learning experiences from nondominant groups during the COVID-19 pandemic. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (CSCW2):307:1–307:27, October 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3476048.

Yigitbas:2019:CDD

[YHR^{+}19] Enes Yigitbas, André Hottung, Sebastian Mansfield Rojas, Anthony Anjorin, Stefan Sauer, and Gregor Engels. Contextand data-driven satisfaction analysis of user interface adaptations based on instant user feedback. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3 (EICS):19:1–19:20, June 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3331161.

Young:2020:DMA

[YMG^{+}20] Gareth Young, Hamish Milne, Daniel Griffiths, Elliot Padfield, Robert Blenkinsopp, and Orestis Georgiou. Designing mid-air haptic gesture controlled user interfaces for cars. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (EICS):81:1–81:23, June 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3397869.

Yin:2019:EMF

[YOFC19] Guofan Yin, Martin J.- D. Otis, Pascal E. Fortin, and Jeremy R. Cooperstock. Evaluating multimodal feedback for assembly tasks in a virtual environment. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3 (EICS):21:1–21:11, June 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3331163.

Yeh:2020:EST

[YRIV20] Yen-Ting Yeh, Quentin Roy, Antony Albert Raj Irudayaraj, and Daniel Vogel. Expanding side touch input on mobile phones: Finger reachability and twodimensional taps and flicks using the index and thumb. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4(ISS): 206:1–206:20, November 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3427334.

Yamanaka:2020:RDG

[YU20] Shota Yamanaka and Hiroki Usuba. Rethinking the dual Gaussian distribution model for predicting touch accuracy in on-screen-start pointing tasks. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4(ISS): 205:1–205:20, November 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3427333.

Zhou:2021:LSE

[ZAS^{+}21] Ke Zhou, Luca Maria Aiello, Sanja Scepanovic, Daniele Quercia, and Sara Konrath. The language of situational empathy. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CSCW1):13:1–13:19, April 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3449087.

Zakaria:2019:SSD

[ZBL19] Camellia Zakaria, Rajesh Balan, and Youngki Lee. StressMon: Scalable detection of perceived stress and depression using passive sensing of changes in work routines and group interactions. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3(CSCW): 37:1–37:29, November 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3359139.

Zytko:2019:DSM

[ZD19] Douglas Zytko and Leanne DeVreugd. Designing a social matching system to connect academic researchers with local community collaborators. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3 (GROUP):236:1–236:15, December 2019. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3361117.

Zhang:2021:JAU

[ZMR21] Xinlei Zhang, Takashi Miyaki, and Jun Rekimoto. JustSpeak: Automated, user-configurable, interactive agents for speech tutoring. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (EICS):202:1–202:24, June 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3459744.

Zhang:2020:HDD

[ZMW20] Amy X. Zhang, Michael Muller, and Dakuo Wang. How do data science workers collaborate? roles, workflows, and tools. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW1):022:1–022:23, May 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3392826.

Zhang:2021:AMO

[ZMZL21] Futian Zhang, Sachi Mizobuchi, Wei Zhou, and Edward Lank. Analyzing midair object pointing mappings for smart display input. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(ISS): 490:1–490:18, November 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3488535.

Zhou:2021:TTB

[ZRL21] Xingchen Zhou, Pei-Luen Patrick Rau, and Xueqian Liu. “Time to Take a Break”: How heavy adult gamers react to a built-in gaming gradual intervention system. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (CSCW1):3:1–3:30, April 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3449077.

Zhao:2020:VHV

[ZSOJ20] Kaixing Zhao, Marcos Serrano, Bernard Oriola, and Christophe Jouffrais. VibHand: Onhand vibrotactile interface enhancing nonvisual exploration of digital graphics. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4(ISS): 207:1–207:19, November 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3427335.

Zhao:2021:SBH

[ZWZ^{+}21] Run Zhao, Dong Wang, Qian Zhang, Xueyi Jin, and Ke Liu. Smartphonebased handwritten signature verification using acoustic signals. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (ISS):499:1–499:26, November 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3488544.